A Bibliography of Publications in Communications of the ACM : 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 April 2025 Version 2.63

Title word cross-reference

#9627 [GH79, Lai79].

0 [Fia73]. 1 [Fia73]. 2^{n} [IGK71]. AX + XB = C [BS72b]. ⋃[a_{i}, b_{i}] [FW78a]. cos(x)/x [GK70b, Gau66]. e^{x}/x [GK70b, Gau66]. E_{i}(x) [Pac70]. Ei(x) [Ng70, Red70]. f(x) = A cos(Bx + C) [Spä70b]. g [ES74]. i [Bro76, FR75a]. ∫_{0}^{∞}[exp(−ct)dt/(t)^{1/2}(1 + t^{2})] [Act74]. k [Fox75, Law77]. K_{0}(z) [Bur74]. K_{1}(z) [Bur74]. L_{1} [BR74, FH75, RR73a]. m [LT73, Cha70a]. n [Bra70a, Bra70b, Bro76, FR75a, LT73, RG73, Wag74, Cha70a]. O(n) [Chi78]. p^{n} [Ack74]. ± [Sal73c]. ±f(±f(±f(⋯ ± f(x) ⋯))) [Sal73c]. R [Ehr74]. RC [Bou76, Mar72a]. R × C [Bou74, Han75a]. sin(x)/x [GK70b, Gau66]. t [Hil70a, Hil70b, Hil73a, Hil81a, Hil81b, eL79b]. we^{w} = x [FSC73]. w exp(w) = x [Ein74]. X + Y [HPSS75].

-Distribution [Hil70a, Hil81a, eL79b]. -f [Sal72c]. -Quantiles [Hil70b, Hil81b]. -Splines [ES74]. -word [IGK71].

1 [Rus78]. 1.5 [Bak78b]. 10 [BKHH78, BBMT72]. 149 [Mer62]. 176 [Art63, Sch72c]. 179 [BB74, Lud63]. 191 [Kop74, Rel63]. 195 [Sch72d, Thu63]. 1966 [Ame66].

2048-Word [Wil70a]. 219 [Kas63]. 222 [Gau64a]. 236 [Gau64b]. 245 [Flo64, Red71]. 246 [Boo64]. 259 [Gau65]. 266 [PH65, Sul72]. 282 [Gau66]. 284 [Fle66].

299 [HP67]. 2nd [Mor71a].

3 [Cla76b]. 3-D [Cla76b]. 300 [Gun67, Vos73]. 316 [Bro67, Rad71]. 323 [OS68, Roy73c]. 324 [Bay68, Bay73b]. 332 [Sko75b, Wit68]. 333 [HOK73, Obr73, SS68]. 343 [GB68, Nie72]. 347 [GR70, Pet70, Sin69a]. 352 [Cle69]. 357 [Sin69b]. 363 [Gau69, Köl72]. 368 [Ste70a]. 369 [Sch70a]. 370 [But70, CP78a, Pro72]. 371 [McK70b]. 372 [Dun70]. 373 [Whi70a]. 374 [Whi70b]. 375 [Spä70a]. 376 [Spä70b]. 377 [LS70]. 378 [Pan70]. 379 [HM72, Lyn70]. 380 [LB70a, Mac72]. 381 [Kno70, Sch72a]. 382 [Cha70a]. 383 [Cha70b]. 384 [Ste70b, Ste70d]. 385 [Pac70]. 386 [Bra70a, RG73]. 387 [Fie70]. 388 [HKLS70a]. 389 [HKLS70b]. 390 [HKLS70c]. 391 [McK70a, Sch72b]. 392 [Fri72, SM70b]. 393 [Abd70, Sal72a]. 394 [Dia70, Mar72b]. 395 [Hil70a]. 396 [Hil70b]. 397 [CG70, JK72]. 398 [Rob72, Sto70]. 399 [Sep70]. 3rd [DR72].

400 [Pie74b, Wal70]. 401 [Bra70c]. 402 [Whe73, van70a]. 403 [CT71]. 404 [AV73, LT71]. 405 [DWT71, Hei72]. 406 [How71, How73b]. 407 [Gea71a, Nik73]. 408 [Law73, McN71, Sal73b]. 409 [Sch71]. 410 [Cha71a]. 411 [MW71b]. 412 [Čer71, Wat73]. 413 [LS71]. 414 [GS71]. 415 [BL71b]. 416 [Gus71a]. 417 [Gus71b]. 418 [Ein72a, Ein72b, Pie74c]. 419 [JT72, Wit74]. 420 [Ell74a, Ell74b, Gai74b, MC73, Wil72, Wil73c]. 421 [Kuk72a]. 422 [Whi72]. 423 [Mol72a]. 424 [Gen72a, Goo73b]. 425 [HK72, Pag74b]. 426 [Bro72a, Bro74]. 427 [Lin72]. 428 [Byr73b, Yoh72]. 429 [Squ72, Wil73a]. 430 [PM72]. 431 [Pro74, Rav72, Rav74]. 432 [BS72b]. 433 [Aki72]. 434 [Bou74, Han75a, Mar72a]. 435 [Ful72]. 436 [Bol72a]. 437 [Bol72b]. 438 [Bol72c, HMR72]. 439 [Bol72d]. 440 [Gal73].

441 [Kno73]. 442 [HD73]. 443

[Ein74, FSC73]. 444 [Wag73a]. 445 [Cha73]. 446 [Bro73, PM75]. 447 [HT73]. 448 [BS73]. 449 [Fia73]. 450 [Bul74, KK74, MM73]. 451 [Gol73, Kni75]. 452 [LT73]. 453 [Pie73]. 454 [RK73, She74]. 455 [HW73]. 456 [Fen73, Tes74]. 457 [BK73a]. 458 [RR73a]. 459 [Sys73, Sys75]. 460 [SS73]. 461 [BH73]. 462 [Don73]. 463 [Lew73]. 464 [Rei73b]. 465 [Hil73a]. 466 [Ehr73a]. 467 [Bre73a]. 468 [Pat73]. 469 [LM73]. 470 [Kub73]. 471 [Gau73]. 472 [HR73]. 473 [Pie74a]. 474 [Aki74a]. 475 [Dor75, Fre75, Mas75, Wri74]. 476 [Cli74a]. 477 [Ehr74]. 478 [BR74, FH75]. 479 [Mag75, Pag74a]. 480 [LS74]. 481 [Cra74]. 482 [MR74]. 483 [Wat74]. 484 [Bur74]. 485 [ES74]. 486 [Vei74]. 487 [Pom74]. 488 [Bre74]. 489 [FR75a]. 49 [Her61]. 490 [GZ75]. 491 [Gib75a]. 492 [Gib75b].

60 [Atk73, BB77, Led72]. ’68 [ABE77, BLSW71, van72].

77 [Bra78]. ’78 [ABB^{+}79].

= [Kow79].

A1 [De 73, BS73, CT71, LM73]. Abstract [GW75, Gut77, GHM78, Lum73]. Abstraction [LT77, LSAS77, SWL77]. Abstractions [HRM78, SS77]. Academic [MM76, BT79]. accelerating [TT82]. Access [BCLS74, Cla77, JL78, Mul72, WS75, Wil70a, Bay73c]. Accesses [Wil76, Yao77]. accounting [Ham73]. Accumulators [KBS75]. Accuracy [CL79, Fis77, KC73b, Lin75, YS77]. Accurate [Lin70, Mal71]. Achieving [Fis77]. ACM [GH79, Lai79, Abd70, Aki72, Aki74a, Art63, Ash72, ABB^{+}79, BR74, BS72b, Bay68, BS73, Bol72a, Bol72b, Bol72c, Bol72d, Boo64, BL71b, Bra70a, Bra70c, Bre73a, Bre74, Bro72a, BK73a, Bro73, Bro67, BH73, Bur74, But70, Čer71, Cha71a, CG70, Cha73, Cha70a, Cha70b, Cle69, Cli74a, CT71, Cra74, DWT71, DR72, Dia70, Don73, Dun70, Ehr73a, Ehr74, ES74, Ein72a, Fen73, Fia73, Fie70, Fle66, Flo64, FR75a, FSC73, Ful72, Gal73, Gau64b, Gau64a, Gau65, Gau69, Gau73, Gea71a, Gen72a, Gib75a, Gib75b, GZ75, Gol73, GS71, GB68, Gun67, Gus71a, Gus71b, Her61, HR73, HP67, Hil70a, Hil70b, HD73, Hil73a, HT73, Hor77, HMR72, How71, HKLS70b, HKLS70a, HKLS70c, HW73, HK72, JT72, Kas63, Kno70, Kno73, Kub73, Kuk72a, LB70a, LM73, Lei69, LS70]. ACM [Lew73, Lin72, LT73, LT71, Lud63, LS74, Lyn70, LS71, MM73, Mar72a, McC75, McC74, MDB79, McK70b, McK70a, MR74, McN71, MW71b, Mer62, Mol72a, OS68, Pac70, Pag74a, Pan70, Pat73, Pie73, Pie74a, PH65, Pom74, Pro72, PM72, Ral72, Rav72, Rei73b, Rel63, Rev72, RK73, RR73a, SS68, SS73, Sch70a, Sch71, Sep70, Sin69a, Sin69b, SM70b, Spä70a, Spä70b, Squ72, Ste70a, Ste70b, Sto70, Sys73, Thu63, Vei74, Wag73a, Wal70, Wat74, Whi70a, Whi70b, Whi72, Wil72, Wit68, Wri74, Yoh72, van70a]. activities [Lee77]. acyclic [Han79]. Adapting [Sto73]. Adaptive [Agn76, Gal73, Hil70c, JP73, Sch70b]. Adaptively [Lyn70]. Adaptively-Noise [Lyn70]. Addendum [MDB79, MN73, MB73, Sim74]. Addition [Sou75, War77]. Additional [LY72]. Address [Jon70, LYD71, Lum73, LY72]. Addresses [FS76, Ric71]. Addressing [Fab74, FS76, SD76]. Adjacent [Hin75, Roy73b, Roy73a]. Administration [MN73]. Advanced [Aus77, Bob72]. advantage [Fra77]. Aesthetics [Ers72b, Ers72a]. Affix [Cro72]. agenda [KC79b]. Aggregation [SS77]. Aid [AC75a, Gel78, GB77, Lin73, SF79]. Aided [BS70, CP78b, NKHS76]. Algebra [Bac78, CP78b, SC75]. Algebraic [LS70, Mos71a, BF71a, Jef71, Mor71a]. Algol [Atk73, Bal74, BB77, BLSW71, Bro77, Kni70, Led72, van72]. Algol-based [Bro77]. Algorithm [Abd70, Aki72, Aki74a, Art63, BR74, BS72b, Bar78, Bay68, BS73, Bol72a, Bol72b, Bol72c, Bol72d, Boo64, BL71b, BM77, Bra70a, Bra70b, Bra70c, Bre73a, Bre74, Bre77, Bro72a, BK73a, Bro74, Bro73, Bro67, Bro76, BH73, Bur74, But70, CT77, Čer71, Cha71a, CG70, Cha73, Cha70a, Cha70b, Che70, Chi78, CO76a, Cla76a, Cla78, Cle69, Cli74a, CK77, CT71, Cra74, DWT71, Dia70, Don73, DL78, Dun70, Ear70, Ehr73a, Ehr74, ES74, Ein72a, Fen73, Fen71a, Fia73, Fie70, Fle66, Flo64, FR75a, FSC73, Ful72, Gal79, Gal73, Gau64b, Gau64a, Gau65, Gau69, Gau73, Gea71a, Gen72a, Gib75a, Gib75b, GZ75, Gol73, GS71, GB68, GW75, GW76, GM78, GR70, Gun67, Gus71a, Gus71b, Her61, HR73, HP67, Hil70a, Hil70b, HD73, Hil73a, Hin75, Hir75, HOK73]. Algorithm [HT73, Hor75, HMR72, How71, HKLS70b, HKLS70a, HKLS70c, HS77, HW73, HK72, JT72, JB73, Kas63, Kno70, Kno73, KD71, Kow79, Kro74b, Kro74a, Kub73, Kuk72a, LB70a, LM73, LCW74, Lei69, Lev76, LS70, Lew73, Lin72, LT73, Loe70, LPW79, LT71, Lud63, LSWL75, LS74, Lyn70, LS71, MM73, Mar72a, MN72, McK70b, McK70a, MR74, McN71, MW71b, Mer62, MB73, Mol72a, Mor78a, Obr73, OS68, Pac70, Pag74a, Pan70, Pat72, Pat73, PR79, Pet70, Pie73, Pie74a, PH65, Pom74, Pre79, PF76, Pro72, PM72, RZ79, Rav72, Rei73a, Rei73b, Rel63, RK73, RR73a, Rob77, SS68, SO74, SS73, Sch70a, Sch71, Sep70, Sho78, Shw71, Sin69a, Sin69b, SM70b, Spä70a, Spä70b, Squ72, Ste70a, Ste70b, Ste70c, Sto70, Sun90, Sys73, Thu63, Tie70, Vei74, Wad76]. Algorithm [Wag73a, Wal72a, Wal70, War75, Wat74, Wat70, Whi70a, Whi70b, Whi72, WS75, Wil74, Wil71, Wil72, Win71, Wit68, Wri74, Yoh72, van70a, BJ71, BL71a, CR79, Cha72, CO76b, DKHM76, FLB73, HZ78, KJ85, Mif70, Pat71a, Pat71b, Pee83, Ste73, And76, And79, AV73, Bay73b, BB74, Bou74, Bou76, Bra70d, Bro74, Bro76, Bul74, Byr73b, Col78, Dav76, De 73, Dor75, DH73, Dut76, Ein72b, Ein74, Ell74a, Ell74b, Er85, Fei75, Fre75, FH75, Gai74b, Gau66, GK70b, Ged79, Goo73b, GR70, Gus78, HM72, Han75a, HKM80, Hei72, Hil81a, Hil81b, HP85, HOK73, How73b, Ito76, Jan77, JK72, Ker73, KK74, Kni75, Köl72, KD71, Kop74, KM76b, Lau73, Law73, Lea79, Lon70, eL76, eL79b, Mac72, MC73, Mag75, Mar72b, Mas75, Mis75, Mor76, Ng70, Nie72]. Algorithm [Nik73, Obr73, Pag74b, Pet70, Pie74c, Pie74b, PM75, PS76, Pom76, Pro74, Rad71, RG73, Rav74, Red70, Red71, Rob72, Roy73c, Sal72a, Sal73b, Sch78a, Sch72a, Sch72b, Sch72d, Sch72c, She74, Sip77, Sko75b, Sko75a, Sko78, Ste70d, Sul72, vSdH79, Sys75, Ten77, Tes74, Vei77, Vos73, Wat73, Whe73, WGL76, Wil73a, Wil73c, Wit74]. Algorithmic [AA71]. Algorithms [Agn76, BS77, BJ74, Buz73, Cla76c, CR78, Deu72, Ehr73a, FGG78, GM74, Hir73, Hir78, HT73, Ive76, KR81, Lew73, Lon70, Mal72, McM78, McM79, Red71, Roy73b, SD76, Sla79, Ten78, VD75b, Ver72, KJ74, Knu72a, Knu76, Lin71, Pro71, Roy73a, Fri72, VD75a]. Aliasing [Cro77]. Allocation [BB78, Bur76, Fer71, Fre74, GK74, GB77, HMT79, Hir73, IGK71, LSWL75, MPY79, Mor74, Nie70, Nie77, SP74a, Sho75, Sho77, Ham73, MD70, SP75]. Allowable [Blo70]. Allowing [Fen71b, WM92]. Almost [Kub73, IM70]. Alphabetic [Byr73b, Yoh72, Byr73a]. Alphard [SWL77]. Alston [Wil75b]. Alternating [SS73]. Alternative [Kes77a, WS75, Bal74]. Alternatives [MS77b]. Altran [Hal71]. Alway [MW78]. Ambiguous [AJU75, KJ73, Mau70]. American [Ano70c, Ano70d, DD77a, DK78]. AMESPLOT [Hir70]. Among [Boy79, LPW79]. Analyses [CR78]. Analysis [Abe74, AC76, Bar78, BHH78, BC75, Car75, CT77, CP78b, CO76a, CO76b, Dam70, Fis77, GW75, GW76, Hod70, HW73, KM76a, Lea79, Lin73, Lum73, MS77b, McM78, McM79, Nie71a, NKHS76, OC71b, Ric72b, Ros77, Sch77, Sed72, SF79, SBB72, SR73, TP77, TP72, Wad76, Wal76, Weg75, Woo70b, de 69, Lyo74, Pro71]. Analytic [LS71, Ray76, Rok73, Str74a]. analyzer [Wil75a]. anatomy [KC79b]. Ancient [Knu72a, Knu76]. AND/OR [Hal73]. Angle [Kli71, Kno70, Sch72a]. Animation [BW76, TCCH71]. Animator [TCCH71]. Anomalies [FGG78]. Anomalous [Sho77]. Anomaly [Dem70, Pom70, Wil76]. ANSI [Ame66]. Answer [Nie70]. Answerer [SBS70]. Answering [Sim70, Wal78]. APL [GJ79]. Application [Alt79, BS70, BW74, FM78a, Led72, LST78, Mar76, SBGN70, SL71, TCB78, BF71a, van77]. Applications [BER76, HHKW77, Hea71, Hor75, Ros76]. Applied [GR73, Gon76, HW79, PM78]. Approach [BCLS74, Chr72, Fra73, KFM71, Lyo78b, MV72, MS77a, Mil77, ML72, Mul77, OG76, RF77, SS76, Shw74, Ski74, WAG77, BYG92, Bal74, Fol71, Lyo74, van77]. Approaches [BK73b, Hun70]. Approximate [RS72, RS74, War78, Ste73]. Approximating [Pat72, Yao77]. Approximation [GS71, NW79, RR73a, Wat70, You74, Dun72, LR70, Mon70]. Approximations [Phi71, Wil74, Wil70b]. Arbitrary [Abd70, FS75, Sou75, Sal72a]. Archaeology [Alt72, BW74]. Architecture [BFS78, CP78a, Fos72b, SS72a, Tan78]. Architectures [Lun77, PG74]. Arcs [Bre77]. Area [FS75]. Argument [GZ75, Mor76, Wil70c]. argument/function [Wil70c]. Arguments [Bur74, Gau65, Jan77, Pre74]. Arising [KS77]. Arithmetic [DG73, GM74, How71, How73b, LM73, Mal72, RL71, Ehr70b, RG71, Sto73].

ARPANET [KNO76a]. Array [KBS75, MR73, Mul77, PA72]. Arrays [DEL78, Deu72, LK72, Rey79, Sit70]. Arrow [Cra74]. Art [Knu73, Knu74a]. arts [LRT77]. aspect [BF73]. aspects [Est73]. Assembling [Szy78]. Assertions [MW78]. Assessment [SC78, WG79]. Assigned [O’R72]. Assignment [BL71b, Wil71, BL71a]. assist [Fra70]. Assisted [Win71]. Associated [Cle69, Fri72]. Associative [Ben75, SB77]. Assumptions [Wag70]. Asymmetric [RS74, Wil73b]. Atlas [Lav78]. Atoms [Les79]. Attitudes [Nol77]. Attribute [JOR75, Lum70, RL74, WC71]. Authentication [EKW74, NS78]. author [FLB73, KJ74]. authored [Dwy72]. Automata [Arb72]. Automated [Kli78a, OC71a]. Automatic [Dam70, DB76, Gea71b, Jef71, Low78, Low70, MW71a, MM78c, ML72, Pat73, Ric72b, Rok73, SWY75, Sch75, SF71, SN73]. Automating [Alt79]. Automation [AKA71, Zd77]. Automaton [Bav70]. Automorphisms [Bav70]. Availability [CP78b]. available [Sto73]. Average [DEL78, FM71, FM72, Jon72, ST74]. AVL [Fos73, Tan72]. Avoidance [Dev77]. Avoiding [Fra73]. Axiomatic [Hun70, Mac73, OG76].

B [HS78, Sny78a, Sny78b]. B* [McC77]. B*-Trees [McC77]. B-Trees [Sny78a, HS78, Sny78b]. Babylonian [Knu76, Knu72a]. Back [CHI^{+}74]. Back-end [CHI^{+}74]. Backtrack [BR75]. Balanced [Hir76, KFSK76, Kos78, RZ79, ZM78]. Balancing [BS77, Pie75]. Band [Cra73]. Band-Symmetric [Cra73]. BANDSOLVE [Sch72d, Thu63]. Banks [Cod70]. Base [CHI^{+}74, Car75, Gol75, OC71b, Shn73, SC75, Sou75, Wal78, Bae70]. Based [Aki72, Aki74a, And76, And79, Fab74, Gai72, Lau75, Pag72, Rey70, RL74, Wil74, WC71, Aki74b, Low70, Mor70a, Bro77]. Bases [DPR70]. Basic [Chr78, Gib75a, Gib75b, Lie76, May79]. Basis [KW75, Mac73]. Batch [BM76, Wen78]. Be [Bac78, Par72a, Gri77b, McC75]. Behavior [RR73b, Sho77, TS77, Ber70]. Bel [HZ78]. Bel-Petrov [HZ78]. Believing [Gri77b]. Bell [LB70b]. Bessel [Bra70d, Sko75a, Bur74, Gau64b]. Best [AA71, Bay77, BK73b, Phi71, Sha77, Sho75, Van74]. Best-Fit [Sho75, Bay77]. Best-Match [BK73b, Sha77, Van74]. Beta [BB74, Gau64a, Lud63, PS76, Che78]. Better [Fos72a, MW78, Led74]. Between [ER78, Hal73, Hec78, Lin75, Fra72, GC72]. Beyond [Win79]. Bibliographic [AC75a]. Biharmonic [Ehr73b, GS72]. Binary [Ben75, BER76, Cha71b, Cha73, DL78, FM71, FM72, Gla71, HKLS70b, Hur71, Jon72, Kno77, Lau73, Lev72, MN72, PM78, RZ79, Wal71, War75, de 69, Ehr70b, Rad70b]. Binding [Bak78b, Ham76]. Biomedical [Hod70]. Bit [IGK71]. Bivariate [Aki74a, And79, Don73, Aki74b, BP73]. BLISS [WRH71]. Block [Kha72, Sit78]. Block-Transfer [Sit78]. Blockage [BC75]. Blocks [BL79, Fle66, Gim72, Hin75, IGK71, Ito76, Pat71a, Pat71b]. BMD [Fra76]. BMDP [Fra76]. BNF [Led74]. Bonus [Mor72b]. Boolean [GH79, Lai79, Han77, Mar72c, Pin70, de 70]. Bound [Bra70b]. Boundary [Kim70, Hos73]. Bounded [Bra75, Fis75, Lin74, Loe70, Rob77]. Bounded-Context [Loe70]. Bounds [FR75b, Mai77, Yap76]. Boyer [Gal79]. Braille [RF72]. Brains [Arb72]. Breaking [PR79]. Brélaz [Pee83]. Brent [FLB73]. Broadcast [DM78]. Bromwich [Pie73]. Brother [OSW78]. Brown [Pee83]. Bubble [BL79]. Buddy [Bur76, CT75, Hin75, PN77, SP74a, SP75].

Buffer [Fer71, Gel73]. Buffers [BC75]. Burnett [Sto74]. Business [Alt79, MT71, Zd77, LML74]. Byte [Lam75].

C [GR70, Smi75b, Smi75a]. C1 [Bro73]. C2 [JT72, Squ72]. C5 [FSC73, LS71, Rad71]. C6 [Pie74a]. CAI [Dwy72, Kof72]. Calculation [Ein72a, Ein72b, Jon70, Pie74c, SS73, ST74]. Calculi [BW70, BW72b]. Calendar [Lam79, Ano70e]. Can [Bac78, Led74, Wir77b]. Canonical [WC71]. Capabilities [PKT73]. Capability [Fab74]. Capability-Based [Fab74]. Cardiography [Hor75]. Care [Was75]. Carlo [Gal73]. Carotid [SKK76]. Case [Car77, Gal79, MT79]. Cauchy [MR72]. caused [Par79]. CCD [Sit78]. celestial [Jef71]. Cell [BJ74, Fen71b, JB74]. Cellular [LK72, NW79]. Center [CH76]. Central [Sla79]. centralization [OH72]. Certain [ET75, FJ70, Ham71b, BP73, GK70a]. Certification [AV73, Bra70d, DD77b, HM72, Kni75, Köl72, Kop74, Lei69, Lon70, Mac72, Ng70, Nik73, PM75, RG73, Red71, Sul72]. Chain [Chi78, Bay73c]. Chained [Hu72, Ken72]. Channels [Mer78]. Character [Gim74, KW75]. Characteristic [Cle69, Fri72]. Characteristics [Boy74, LST78, MB76a, Fai79]. Chargeout [Nol77]. Cheap [Kli79b]. Chebychev [Sch71]. Chebyshev [Bro73, GS71, PM75]. Checking [Led72]. Checkpoint [You74]. Chemical [Les79]. Chen [Smi75b, Smi75a]. Cheney [Fen71a, Wal72a]. Chess [KC73a]. Chi [Gol73, HP67, HP85, eL76, Kni75]. Chi-Square [Gol73, Kni75]. Chi-Squared [HP67, HP85, eL76]. Choice [Sch75, Sha77, Kli78b]. Chosen [KZ75]. Chromatic [Bro72c]. Cigarette [Par75]. Cipher [McC75]. Ciphers [PR79]. Circle [Kan72]. Circles [KBS75]. Circuit [BS70]. Circuits [Sys73, Sys75, Tie70, Har71a].

Circular [Bre77, CT71, CR79]. Circularity [JOR75]. Clarification [ANS69, ANS71]. Class [BH73, HMT79, Hir73, KS77, Smi75c, MP74]. Class-Teacher [Smi75c]. Classification [Sla79, HZ78]. Clenshaw [Ged79, Gen72a, Gen72b, Gen72c, Goo73a, Goo73b]. Client [Kli78a]. Client-Tracking [Kli78a]. Clipping [SH74]. Cliques [BK73a, KSW78]. Clocks [Lam78]. Closed [Buz73, FS75, RK75]. Closure [War75]. CLU [LSAS77]. Cluster [Lin73]. Clustering [Bel70, Mag75, Pag74a, WGL76]. COBOL [van77, LC76, Lem79, LD76]. Code [BK70, Bel70, Bel73, BER76, Car77, Dew75, LB70b, Mye78, Sam78, Szy78, Sal73c, Sto73]. Coded [Bay73d]. Codes [CRSW74, Ham71b, O’R72]. Coding [Blo70, Byr73b, Flo76, Wir69, Wir70, Yoh72, Byr73a]. Coefficients [Gus71a, LS71, Pie74a]. Coffman [Sto74]. Cognition [Fin77]. COKO [KC73a]. Cold [EF78]. Cold-Start [EF78]. Collection [Bae72, Bar77, DLM^{+}78, Mor79, Ste75, Ste76, Wad76]. Collector [DB76]. Colleges [AE73, LRT77]. Collision [LPW79]. Collision-Free [LPW79]. Color [Bré79]. Coloring [NT74, SR73, KJ85, Pee83]. colouring [Kor70]. Columbia [Ano70a]. Combinations [Cha70a, Kno74, Kno76b, LT73]. Combinatorial [Ehr73a, Sto74]. Combined [Lum70, Mul71, ZC77]. Combining [Shw75]. Comm [GH79, Lai79]. command [Gra70]. Commands [Dij75]. Commas [ER78]. Comment [Ban77, Ber70, Dem70, FLB73, Fen71a, Hu72, Hun70, Jon72, Lai79, LB70b, Law77, PHH72, Pol71, Red71, Smi70a, Wex73, de 70, van72, Est73]. Comments [AA79b, AA79a, Hee70, Hol71, KJ73, KJ74, SM70a, Smi75b, Smi75a, Smo72, EM72]. Commitee [McC74]. Committee [ABB^{+}79, MDB79, Ash72]. Common [Bra70a, Bra70b, Hir75, HS77, RG73, Wag73b]. Communicating [GC72, Hab72, Hoa78b, Hoa78a, Pag73]. Communication [HL77, Fra77, Wal72b]. Communications [FJ70, Mer78, Pac73, Fra72]. Community [Pag72, Pre74]. Compact [AG74]. Compactifying [Ste75, Ste76]. Compacting [Che70, Wal72a]. Compaction [Fen71a, KW75, Mor78a]. Comparative [Fox72, Lie76, TP72, Fra70]. Comparison [ACD74, BS77, BS72a, Bay77, Gre72, Hil70c, Mor79, Sch70b, Ste78, SB77, VD75b, Wil76, FM78b, Jon86, VD75a]. Compilation [Sha74]. Compile [Bar77]. Compiler [CW73, Dit76, Sit70, Wal76]. Compilers [Car77, MS73, Atk73, Bal74]. Compiling [BRT79, BG72a, Ber72, EC72, Gla71]. Complement [Mau70, SM71, War77, Ehr70b]. Complete [BW70, Mer62, Sko78]. Completeness [Rey70]. Complex [AV73, Bra70c, Bur74, Dun70, Ehr70a, Gau69, JT72, Kuk72a, Kuk72b, LT71, Pag73, RK73, RL71, She74, Wit74, Str73, Köl72]. Complexity [HSU75, JOR75, Poh72, Rab77, Rab78, FW78a]. Components [HCS79, de 69, Pin70]. compose [MM70]. Composed [Lev76]. Composing [Rad74]. Composition [Moo72, BLSW71, McM73, Smo72, van72]. Compressed [Mal76, MS73]. Compressing [AA71]. Compression [Hah74, Rub76]. Computation [Chi78, ES74, FP79, FG74, Gus71a, Gus71b, Gus71c, MV72, Pie74a, Ric72a, DKHM76, GK70a, Mar72c]. Computational [Buz73, de 69, McM73]. Computations [Mol72b, Rab77, Rab78, Tho70]. Computer [AT77, AD72, AK73, Ama71, And78, Ard76, AE73, Aus77, ABE77, ABB^{+}79, BM76, Bae72, Bak78a, Bal73, BS76, BFS78, BN76, Bou70, BS70, BB78, CGM79, CHI^{+}74, CP75, CP78b, Chr78, CH76, Cro77, Eld70, EKW74, Fag76a, Fag76b, FS79, Fos72b, Fox72, Fra76, FJ70, Gai72, GR73, HOR72, HWBC77, HMS72, IC78, Kan72, Kan73, KJ73, Knu73, Knu74a, Lew73, Lin73, Mam77, MM78a, MDB79, MT71, Mer73, MB76b, Mil75a, Mil75b, MM76, Moo72, ML77, Mul77, MR70, NGMM70, NS76, Nie70, Nie77, Nol73, Nol71, NKHS76, O’R72, Pac73, Pag72, Pho75, Pro74, Rad74, Rav72, Rav74, Rus78, SG75, Sch78b, She73, Shw71, Shw74, Spr72, SB77, Tad78a, Tad78b, Taj77, TC77a, TC77b, TK77, Tra72, Wag70, Wal74, Wex73, Win71, Wir69, Wir70]. Computer [WAG77, Ash72, BF73, Col79, Est73, Fai79, Fis76, Fol71, Kim72, LRT77, LML74, McM73, Pol71, RB70, Ros76, Smo72, Smo76, TC76, TC79, Wal72b, WG75, WG79, Knu72b]. Computer-Aided [BS70, CP78b, NKHS76]. Computer-Based [Pag72]. Computer-Communications [Pac73]. Computer-Generated [Cro77]. Computer-Oriented [MT71]. Computerized [Ste74, Ste79]. Computers [BG72b, DD77a, DK78, HCS79, HMS72, KD78, McC74, NS78, SW77a, SW77b, Stu72, Tei71, Was75, Zin72, Alt72, Ehr70b, Fra72, Ham71a, Sto73]. Computing [CL79, Chr72, Den79, ET75, Fel79, Fin75, Gen72c, Hir75, HCS79, HS77, Kli78a, Min74, Nie71b, Wei72, WD78, FW78a, Law77, LS74, OH72]. Concept [Bri78, Gai72, Hoa74, Lum73, PS75, Rey70]. Concepts [BCD72]. Conceptual [Ama71, Gol75, Kes77b]. Concerning [McC74, McM73, Was75]. Concurrent [Bri78, CHP71, KD78, Lam74a, Lam77, EM72]. Condition [Rok70]. Conditional [FW78b, Par75]. Conditions [FM78a, NT74]. Conference [Hor77]. Configurations [Teo78, CR79]. Confinement [Lam73]. Conflict [Gri70]. Conflicts [KSW78, Kli78b]. Conjugate [Hes75]. Connected [HCS79, TK77].

Connections [Lin75]. connectivity [Low70]. Consecutive [Gho72, Gho75]. Considerations [Lee77]. Considered [Smo74, Dij68]. consistency [EGLT76]. Constant [AHS76, Cla76a]. Constants [Tod75a, Tod75b]. Constrained [RK73, She74]. Constraint [Fre78]. Constraints [JL78]. construct [Sha70]. Constructing [MS73]. Construction [Bro71, Loe70, Pre74, SS74b, SK74, Kni70]. Constructs [Hen72]. Consulting [SK75]. Consumer [Ste79]. Content [Dam70, Smo74, Hil73b]. Contention [OS72]. Context [Ear70, Hal73, Loe70, Mau70, SB77, Weg76, Woo70a, Lyo74]. Context-Free [Ear70, Hal73, Mau70, Lyo74]. Context-Sensitive [Woo70a]. Contingency [Bou74, Bou76, Han75a, Mar72a]. Continued [Wil74]. Continuous [GS71]. Contour [AG74, Mar76, ML72, War78, Mon70]. Contours [FKU77, Mer73]. Contraction [Sha74]. Contradictory [LS77]. Control [CHP71, DPR70, Dij74, HMT79, Kow79, Kuk72a, Kuk72b, LM75, LC76, McG73, Mil73, MPT77, Sal74a, BJ71, EM72]. Controlled [Mye78]. controller [RF72]. Conventions [Gan77]. Conversational [Wil70a]. Conversion [BJ74, Cav74, Dat72, GR73, JB73, KJ73, Lew78, MR70, Rob72, SS76, SHL75, Shw71, Sou75, SLSH75, Sto70, Ver72, KJ74, Pol71]. CONVERT} [SLSH75, SHL75]. Converting [Shw74, Kai70, Sal73d]. Convex [PH77, Pre79]. Cooper [KC73a]. Cooper-Koz [KC73a]. Cooperation [DLM^{+}78]. Copying [Cla78, Fis75, Lin74, Rob77]. Correct [Gri77b, Gri77a, Gri78]. Correction [Ehr70b, JP73, Wag74, Mor70b, Pee83]. Correctness [Gel78, Lam79, LCW74, MW78, Sam78, Taj77]. Correlated [HK72, Pag74b]. Corrigenda [HH70a, Wir70]. Corrigendum [AA79b, CO76b, DS73, Ers72a, Fag76a, FM72, GH72, Gri78, Hoa78b, McM79, Mil75a, Pat71b, Rab78, SW77a, Sed79, She72, SP75, Ste76, TC77a, Tod75a, VD75a, van75a]. COSERS [Ard76]. Cosine [Lin72, Gen72c]. Cost [BED77, Lew78, LSWL75, OH72]. Cost/utilization [BED77]. Costs [Lip79, Sha74]. Cotes [Hil70c]. Coulomb [Gun67, Vos73]. Counter [SS74b]. Counterintuitive [Fag76a, Fag76b]. Counting [Mor78b]. Counts [Fre74]. Coupled [BB78]. Course [AE73, CH76, HMS72]. Covering [KSW78]. CPI [Ano70c, Ano70d]. CPU [SBB72, Teo78, TCB78]. CPU-I [Teo78]. CPU-I/O [Teo78]. Cray [Rus78]. Cray-1 [Rus78]. Creation [DPR70]. Crisis [MDB79]. Criteria [Lew78, Par72a]. Criterion [McC75, Lin71]. Critical [Fer74]. critique [DH76]. CRT [ID72, Luc74]. Cryptographic [Ham71b]. Cryptosystems [RSA78]. crystallography [Les72]. CSIEVE [Dun70]. cubatures [BP73]. Cubic [BH73, FS73, Hos73, Nil70]. Cumulative [Pom74, Pom76]. Current [Lip79, MM78a]. Curricula [MT71]. Curriculum [Ama71, ABB^{+}79, Cou73, MN73, Ash72, ABE77]. Curtis [Ged79, Gen72a, Gen72b, Gen72c, Goo73a, Goo73b]. Curve [Aki72, And76, Cli74a, Cli74b, FS73, FS75, LS77, Sch71, SA77]. Curves [Mon71, DH72]. cut [Pat71a]. cutnodes [Pat71b]. Cycle [Gib75a]. Cycles [Gib75b]. Cyclic [Fis75, Rob77, SF79, Low70]. cyclical [PW70]. Cycling [KS77].

D [Cla76b]. D1

[Bol72b, Bol72c, Bol72d, Ein72a, Gal73, Gen72a, Gus71b, Lin72, Pat73, Pie73]. D2 [BH73, Gea71a]. D3 [SS73]. D5 [KM76b]. Dangerous [Smo74]. Data [AC76, AA71, Bar78, BL79, BCKH75, BW74, CHI^{+}74, Car75, CH75, Cod70, DPR70, Ear71, Fle66, FM77, Gan77, Gel78, Ger75, Gim72, GG77, Gut77, GHM78, Hah74, HHKW77, Han75b, Hir70, Hol77, Ito76, JL78, LT77, Low78, LSWL75, ML77, Mul77, OC71b, PF78, Ros77, Sch75, Shn73, SS74a, SHL75, ST73, Sil76, SC75, Spä70a, SLSH75, Vui78, Wal78, Weg74b, Ano70e, Bae70, Cot75, Fra72, GJ79, Rad70b, DH76]. Database [Min76, Sil76, SS77, Sto88, Yao77, EGLT76, Han79, SC78]. Databases [Mil77]. Date [Rob72, Sto70, Ano70e]. day [Sto73]. DBTG [Ger75]. Deadlock [Dev77, How73a, PHH72]. Deadlocks [Fra73, Hol71, Bae70]. dealing [SC78, WG79]. Decade [Mos71b]. decentralized [CR79]. Decimal [CH75, Kai70]. Decision [Cav74, CR76, Dat72, Dia70, GR73, IR78, KJ73, Lew78, Low73, Mar72b, MM78b, MR70, SS76, Shw71, Shw74, Shw75, Ver72, KJ74, Pol71, Str73]. Decisions [Fab73]. Decomposability [Cou75]. Decomposing [Par72a, Pin70]. DECsystem [BKHH78]. Deductive [SBS70]. Defining [SWL77]. Definite [Hil70c]. Definition [SHL75, ST73, SLSH75]. definitions [Wir77b]. Deletion [ZM78]. Deletions [Kos78]. Demand [Boy74, Mor72a, Sal74b, Smi70a]. Denotational [Ten76]. Dense [FM71, FM72]. Dependent [Szy78]. Depth [TS77]. Derivation [Dij75, Lou76]. Derivatives [GK70b, GK70a, Gau66]. Derived [Hen72]. Descendants [Loe74]. Design [Alt79, Bal73, BCD72, BS70, BS79, Cas73, CW73, Ger75, GCD73, Hor77, LS76, Lis72, Lyo78b, MS77b, Mil77, Mis73, NKHS76, PS75, Pfe75, SLR78, WDT76, WS74, Fol71, RRD73]. Designing [Cla76b, CRML73]. Detailed [SMMH77]. detect [DH72]. Detecting [SR78]. Detection [Boy79, Hor75, IR78, Les79, Mar76, Mon71, ZC77, SA77].

Determinants [Chr78]. Determination [FS76, Lew73]. Determining [Chi78, FS75, Ric72b]. Deterministic [AJU75, CR78, KD78, Sch74]. Development [Gut77, IC78, Ivi77, Smo76, Wir71, Ben72b, Knu72b, Str73, Zin72]. developments [Kli78b]. Deviate [HD73]. Deviates [Kno73]. Deviation [Han75b, Cot75, Nel79]. Deviations [CL79]. Device [Mor72b]. Devices [Mor74]. Diagnosis [LD76]. Diagnostic [CW73, WDT76]. Diagram [FG74]. difference [GS72]. Differences [Hec78]. Differential [BH73, CK70, CB74, Cle69, Fox72, Fri72, Gea71a, Gea71b, Nik73, Wil74, Lin71]. Difficulties [Ste79]. DIFSUB [Gea71a, Nik73]. Digital [Bre77, Cla75, RSA78, SR78]. digitized [Mon70]. Digits [Tsa74]. digraphs [Han79]. Dijkstra [EM72, Lam74a]. Dilogarithm [GZ75, Mor76]. Dimensional [Bou70, BF71c, Fei75, Les79, Wat74, TCCH71]. Dimensions [PH77]. Dipole [Kno73]. Direct [BCLS74, Ehr73b, Bay73c]. direct-access [Bay73c]. Directed [Mil70, PM72, Lyo74]. Direction [SS73]. directions [Zin72]. Discipline [SF73, Wir77a]. Disciplines [Ful74, SC79]. Discourse [SS72b]. Discrete [Gim73, RR73a, Sch71, Sha75, Ulr78, Wym75]. Discretized [Pan70]. Discrimination [WG75]. Discussion [Man79, BF73]. Disk [Ben72a, Bur76, Ful74, GK74, MS76, Mor74, TP72, Wil76]. Disks [Pie75]. dispatcher [RRD73]. Display [BJ74, Bre77, yC74, JB74, New71, Nol71, LR70]. Distributed [Bri78, BS79, Dij74, Fel79, Lam78, MS76, MB76b, Taj77]. Distribution [Don73, Gel73, Hil70a, Hur71, Kno73, Pom74, Pom76, Sch70a, TS77, Hil81a, eL79b, Fis76]. Distributions [AD72, FJ70, SC79, Tsa74]. District [Ano70a]. diversity [Wir77b]. Divide [Wil70b]. Division [AHS76, MB73, SM71, WS74, Mif70]. Divisor [Bra70a, Bra70b, RG73]. Divisors [AHS76, MR75]. do [Led74, Wir77b, Lam74b]. doctoral [Fai79]. Document [Sal72b, Van74]. Documentation [Mil70]. Dominoes [Smi73]. Double [Hu72]. Double-Chained [Hu72]. Doubly [Ken72, Whi70a]. Doubly-Chained [Ken72]. Drawing [Lev76]. Drawings [JB74]. Driven [BRT79, BS71, Gon76, SBB72, WDT76, Zel71]. Drum [Ful74, SF73]. Dutch [McM79, McM78]. Dynamic [BCPW74, Fab73, GK74, Hir73, IGK71, Mam77, Nie77, PA72, Sal72b, SP74a, SP75, Sho77, Ten78, TF71]. Dynamics [BW76, Cla75].

E1 [Gus71a, HR73, LS74]. E2 [And76, Aki72, Aki74a, Cli74a, ES74, GS71, RR73a, Sch71]. E4 [Dav76, MM73, RK73]. Early [GMS77]. Economic [CP75]. Economical [ER78]. Economy [Kli78a]. Edge [Mar76, Wis76]. Edges [KSW78, SR78]. edit [MM70]. Editing [FB73, ID72, TP77]. Editor [Ben72a, Wil75b]. Education [ABE77, CP75, Don76, FMP74, Lon79, Lyo78b, Pay73, Tei71, WD78, Ash72, BF73, Est73, Ham71a, NGMM70]. Effective [Low73, YS77]. Effects [Nol77, Pac73]. Efficiency [CZ74, Whe73, van70a, van70b]. Efficient [AC75a, BER76, Cas73, CH75, Cla76a, De 73, DB76, Ear70, FM77, GR70, Ham73, HT73, IGK71, JH75, KR81, Kro74a, Mer73, Mor78a, Pet70, Rad70b, Sin69a, Sin69b, Tie70, Kro74b]. Effort [WD78]. EFT [Kli79a, KC79a, KC79b, Lon79]. EFTs [Par79]. Eigenproblem [Kha72]. Eigenvalue [Cra73, DWT71]. Eigenvalues [GB68, Nie72, Rei73b, Ste70b, Ste70d, van75b, van75a]. Eigenvectors [GB68, Ste70b, Ste70d, Nie72]. Einschrittcompilers [Dit76]. EL1 [Weg74b]. electronic [Kli78b]. Elementary [MR75, Sys73, Sys75, Tie70]. Elements [Bro76, ET75, FR75a]. Eliminating [Bel70]. Elimination [Bir77b, Mat72, PW75]. Elliptic [Mer62, Sko78]. Emotional [Smo74]. Empirical [CG77, Lun77, NS76, RR73b, Ste79, WD78, Wil76, Jon86]. Employment [BG72b, Fai79, TC77a, TC77b, TC76, TC79, WG75]. Emulation [RFE72]. Encasing [FS75]. Encipherment [FH74]. encoded [Ano70c]. Encoding [BCE77, ER78]. Encryption [DH76, NS78, PM78]. End [FFG73, CHI^{+}74]. Engineering [Ard76, BT79]. English [AC75b, Dit76, MS72, SS72b, Wal78, Wil75a]. Enhancing [BW76]. ENTCAF [LS71]. ENTCRE [LS71]. Entry [FB73, Lew78, Shw71, Shw74, Ver72, KJ74]. Enumerating [GW74, LT73]. Enumeration [Ive76, PM78, Roy77b, KJ85, Roy77a]. Environment [BGM70, DPR70, FP79, McG73, Roc71, RFE72, RL74, Wen78, NR71]. Environments [BW73]. Equality [Sho78]. Equation [BS72b, Cle69, Ehr73b, Ein74, Fri72, FSC73, Mol72a]. Equations [BR74, Bro67, BH73, CK70, CB74, Fox72, FH75, Gea71a, Gea71b, Han72, How71, How73b, MR72, Nik73, Pan70, Rad71, Teo78, Ver70a, Wil74, Igu72, Lin71, Ste73, Ver70b]. Equipment [Pfe75]. Equivalence [Hal73, Sch74, SA77, de 69]. Errata [Knu76, Kro74b, Ver70a]. Error [Gau69, GR75, Kuk72a, Kuk72b, LD76, MM78c, Ric72b, Sch70b, GK70a, Köl72]. Error-Proneness [LD76]. Errorbounds [Rok73]. Errors [Blo70, IR78, LD76, Tsa74, Lyo74, WM92]. ESOPE [BBF^{+}70]. establish [Spa72]. Estimates [FJ70, Sch70b, Nel79, Wes79]. Estimating [CZ74]. Estimation [Gif77]. Etching [AKA71]. Etching-Pattern [AKA71]. Ethernet [MB76b]. Evaluating [Hil70c, Sha70]. Evaluation [Bur74, Car73, DG73, Gan77, GCD73, Han77, KD78, Lie76, Lun77, LS71, RK75, Sil76, Sly74, WD78, vB76, GH79, Kim72, Lai79, RRD73]. Evans [IM70]. Event [BGP75, BRT79, CGL73, FM77, Gon76, Kes77a, KZ75, Ulr78, VD75b, Wym75, FM78b, Jon86, VD75a]. Event-Driven [BRT79]. Event-Node [CGL73]. Event-Scanning [Wym75]. event-set [Jon86]. Eventcounts [RK79]. Events [Lam78, Mor78b, Ulr78]. Evidence [LS77, Pro71]. Evolution [Bac72, BKHH78, BHH78]. Exact [Bou74, Bou76, Han75a, How71, How73b, Mar72a, Pom74, Pom76]. Exactly [Ehr74]. examined [HS78, Sny78a, Sny78b]. Example [Fag76a, Fag76b, Low78, SBGN70, SS74b, SLR78]. examples [Par72b]. Exception [Goo75]. Execution [Boy74, FH74, Kin76, Lam74b, Zel73, RG71, Str74a]. Executive [MDB79]. Exercise [DLM^{+}78, Gri77a, Gri78]. Existence [NT74]. Existing [LYD71, LY72]. Exit [PKT73]. Exits [vB73, Eva74]. Expansion [LS70, Sal73d]. Expected [FR75b]. Experience [GMS77, Iro70, Gen72b]. Experiment [Mye78]. Experimental [FS79, Gan77, MDB79, SMMH77, HW72]. Experimentation [HOW74]. Experiments [OC71a, Rub76, SD70, SF71, SN73, Sla79]. Exponential [AD72, Buz73, CB74, Fri72, Gau73, JOR75, Ng70, Pac70, Red70, Spä70a]. Exponentiation [Ric73]. expressed [Har71a]. Expressing [JL78]. Expressions [BCKH75, DG73, Fre78, FW78b, Han77, LS70, GH79, Lai79, RG71, Sha70, Sto73]. Extended [Lew78, Wen78]. Extended-Entry [Lew78]. Extending [Shw74]. Extensible [Ben72a, Iro70]. Extension [JL78, PF71, BL71a]. External [Sho75]. Extracting [Wag73a]. extrema [CR79]. extrema-finding [CR79]. f [Sal72c]. F1 [Lea79, Bre73a, CT77, Mac72]. F2 [DWT71, Rei73b, Ste70b, Ste70d]. F4 [Gus78, HKM80, Sip77, BR74, BS72b, How71, Kub73, Mol72a]. Faces [Bra75]. facility [Gra70]. Fact [Ski74]. Factor [Ers72b, Ers72a]. Factorization [ES74, EH75]. Factors [CP75, OH72]. Faculties [MM78a]. Failures [GD78]. Family [HFC76]. Fashion [Wag73a]. Fast [AHS76, BM77, Cla78, Dit91, Gel73, GW75, GW76, GS72, Hir78, HS77, KMP77, Lit91, Pea90, Pea91, Sav91, She78, Shn78, Sun90, WM92, MP74]. Faster [Weg76]. Fault [FG74]. FCFS [Wil76]. Feasibility [Eld70]. Feature [OC71b]. Features [Lun77]. Feedback [BB78]. Fen [Mac73]. Fibonacci [CT75]. Field [BW74, Fin75, LM73]. fields [HZ78]. Fifty [Sho77]. Fifty-Percent [Sho77]. File [Alt79, Bab77, BK73b, Car73, CE70, Gho72, GB77, Lum73, RL74, Rub76, Sha77, WC71]. Files [Cla77, Hec78, HH70a, HH70b, LYD71, MS76, MPY79, LY72]. Film [TCCH71]. Find [Sla79, Tie70, Hoa71]. Finding [BK73a, Bro76, FR75a, GM78, KBS75, CR79]. Finishing [BCS74]. Finite [LM73, Pat73, PH77, GS72]. finite-difference [GS72]. Finiteness [Wag70]. First [Bay77, Bra70d, Gau64b, Sho75, Sko75a, SF73, You74, Rev72]. First-Fit [Sho75, Bay77]. Fit [Art63, Sch71, Sch72c, Sho75, Spä70b, Bay77]. Fits [Sha75]. Fitting [Aki72, Aki74a, And76, And79, Cli74a, Cli74b, FS73, Spä70a, Aki74b]. five [Rev72]. Fix [MV72]. Fixed [Ber72, Gla71, BG72a]. Fixed-Point [Ber72, BG72a]. Flag [McM78, McM79]. Floating [GM74, Gre72, KC73b, Lin70, Mal71, Mal72]. Floating-Point [Lin70, Mal71, Mal72]. Flow [AC76, Bar78, Den76, DD77b, GW75, GW76, Ros77]. flowchart [RB70]. Flowcharts [SMMH77, Ver72, KJ74]. Floyd [IM70]. Floyd-Evans [IM70]. Fly [DLM^{+}78]. FOCUS [EL79a]. forerunner [BW72a]. Foreword [Wil75b]. Formal [Bro71, Dij75, Gre77, HH70a, HH70b, Kel76, PG74, PF71]. formalism [ES70]. Formatted [LYD71, LY72]. Formula [Har72]. Formulas [Gus71a, Gus71c, Lin75, Pie73, Str74b]. Formulation [SO74]. Forsythe [Her72, Knu72b]. Fortran [Ame66, ANS69, ANS71, Bra78, DPR70, Han74, Mil73, Mol72b, Pay70]. Forum [McC75]. Forwarding [DM78]. Forword [BFS78]. Foster [Tan72]. Foundation [FS79]. Four [Ehr73a, Ive76]. Fourier [Ein72a, Ein72b, Lin72, Pie74c]. fourth [Hos73]. fourth-order [Hos73]. Fraction [Wil74]. Fragmentation [Sho75]. Frame [BW76]. Framework [Ama71, Kes77b]. Free [Ear70, EB77, Hal73, LPW79, Mau70, Roy73b, Roy77b, Lyo74, Roy73a, Roy77a]. Frequency [Hil73a]. Frequently [She78]. Fresnel [Act74]. Front [FFG73]. Front-End [FFG73]. Full [Day70, GW74, Lam75]. Full-Time [GW74]. Full-Word [Lam75]. Function [AV73, Bul74, Dav76, Fie70, Ful72, Gau64a, Gau69, GZ75, Her61, HKLS70a, KK74, Kuk72a, Kuk72b, LT71, LS71, MM73, Mor76, Sch78a, Hal71, Köl72, Pin70, Wil70c, Col78]. Functional [Bac78, BH73]. Functions [AA79b, AA79a, Bra70d, Bur74, CE70, Gau64b, Gau65, GS71, Gun67, HKLS70b, HKLS70c, Jan77, Lag74, Mos72, Rok73, Sko75a, Spr77, Vos73, War77, DKHM76, Mar72c, MM70]. Fundamental [LYD71, LY72]. funds [Kli78b]. Furniture [Pfe75]. Further [EM72, Nel79, Pro71]. Future [Ric72a, Sam72, Lee77].

G1 [Gol73]. G2 [Bou76, Mar72a]. G5 [Bre74, HK72, Kno73, Pro72, Sul72]. G6 [Lei69, Ehr73a, LT73, Roy73c]. G7 [Ehr74, MR74]. Game [Dev77, Fin77, SL71]. Games [Cha72]. Gaming [Lyo78b].

Gamma [AV73, Ful72, Kuk72a, Kuk72b, LT71, Sch78a, Tad78a, Tad78b, Wal74, Fis76]. Gamma-distribution [Fis76]. Garbage [Bae72, Bar77, DB76, DLM^{+}78, Mor78a, Mor79, Ste75, Ste76, Wad76]. Gauss [Bol72c, Bol72d, HMR72, PW75, Str74b]. Gaussian [Bre74, Pie73]. GEDANKEN [Rey70]. Genealogy [LM75]. General [BF71a, But70, Gus71c, Lew78, Lum73, Mos72, RK75, SKK76, Teo78, Eas72, Nie72, Pro72]. Generalization [Fos73]. Generalized [Cra73, DWT71, DS78, OC71b, Sit70, KJ85, Wil70c]. Generated [Bav70, BN76, Cro77, Lew73, Lin73, Pho75, Hil73b]. Generating [Che78, Cro72, Luc74, MS73, RS72, RS74, SS72b, Wil70b, Gra70, IM70, Wil70c]. Generation [BER76, Bou70, Car77, CRSW74, Gib75a, Gib75b, Har71a, HK72, Lei69, Mai77, OS68, Pag74b, PG74, RH72, Roy73c, Tad78a, Tad78b, Wal74, Sto73]. generational [Ben72b]. generative [Kof72]. Generator [But70, De 73, Ehr74, GB68, Sch70a, Sin69b, Whi70b, Bre74, Pay70, Pro72, Sha70]. Generators [BS72a, SWL77]. Genie [Sit70]. Gentleman [Kar75]. Geometric [BF71c, Cla76c]. Geometry [Ray76]. George [Her72, Knu72b]. Gill [Tho70]. Given [Kli71]. Givens [HW79]. Global [GW75, GW76, MR79]. Glut [FMP74, Pay73]. GLYPNIR [LLBR75]. Go [Dij68, Knu74b, Eva74]. Goodman [Kim70]. Goodman-Lance [Kim70]. Gotlieb [SS74b]. GOTO [vB73]. Governments [DK78, DD77a]. Gradients [Hes75]. Graduate [CH76, FMP74, MN73, Pay73, Ash72, LML74, NGMM70]. Grammar [MR73]. Grammars [AJU75, Cro72, DeR71, Hal73, JOR75, LaL77, SG75, Woo70b, IM70]. Grammatical [CRML73]. Graph [Bré79, BK73a, Čer71, Gib75b, HT73, LK72, Min74, NT74, PF71, PM72, SO74, Sys73, Sys75, Tie70, Wat73, Cha72, Kor70, KJ85, Law77, Pat71a, Pat71b]. graph-colouring [Kor70]. Graphic [yC74]. Graphical [Wen78, RB70]. Graphics [Bou70, Nol71, SB77, WAG77, Fol71, Smi70b]. Graphs [Hal73, KSW78, TP77, CRM70]. gravitational [HZ78]. Gray [CW78, BER76, Sal73c]. Graycode [Boo64, Er85, Mis75]. GRE [Aus77]. Greatest [Bra70a, Bra70b, RG73]. GROOVE [MM70]. Group [HW73]. Grown [MN72]. Guarded [Dij75]. Guide [Mos71a, SD76]. Guided [MM78b]. Guidelines [Ste74].

H [KD71, BL71b, Cra74, Gib75a, Gib75b, HT73, MW71b, PM72, Rav72, Sys73, Whi72]. H1 [Fia73]. Haåvie [Pie74b]. Håvie [Wal70]. Hadamard [Bir75]. Half [Bou70]. Half-Toned [Bou70]. Hamiltonian [Har71a]. Hanani [GH79, Lai79]. Handling [Goo75, Rad70b]. Handprinted [NT70]. Harada [RH72]. Hardware [Gif77, SS72a, SB77]. harmful [Dij68]. Harmonic [Str74b]. Harrison [Boo73]. Hash [Ack74, Bat75, Bay73d, BK70, Bel70, Blo70, Bob75, HP78, LB70b]. Hash-Coded [Bay73d]. Hashed [Mul72]. Hashing [AA79b, AA79a, CE70, Dit91, Har71b, Lit91, Pea90, Pea91, Sav91, Spr77, Wil71]. Hasp [Str74a]. Having [SN73]. Health [Was75]. Heaps [Gon76, FM78b]. Height [Hir76, KFSK76, Kos78, RZ79, ZM78]. Height-Balanced [Hir76, KFSK76, Kos78, RZ79, ZM78]. Heuristic [Mar76, Pfe75, Ski74, SF71]. Heuristically [MM78b, Sam78]. Heuristics [Riv76]. hexadecimal [Kai70]. Hexagonal [Deu72]. Hidden [Ell74a, Ell74b, Gai74a, Gai74b, MC73, Mat72, Wil72, Wil73c]. Hidden-Line [Ell74a, Ell74b, Gai74b, MC73, Wil72, Wil73c, Gai74a]. Hierarchical [Cla76c, Sal73a, SLR78, vP70]. Hierarchically [PS75, RL77]. Hierarchies [LSWL75, Wil73b]. Hierarchy [HFC76, Wil71]. High [Fel79, GG77, Ham76, HHKW77, HLL73, Pur74, Ros77, Sch75, SHL75, SLSH75]. High-Level [GG77, Ham76, Ros77]. Higher [CP75, Hir70, WD78, Ham71a]. Highly [KD78, Pat72]. Histogram [CW78]. historical [Lav78]. History [BHH78, MT79, Sam72, Spr72]. Homilies [Ros76]. Honoring [Wil75b]. Hopf [Ste73]. Horner [RK75]. Hough [DH72, SA77]. House [Ste74]. Householder [Wil75b]. Hu [Byr73a, Byr73b, Yoh72]. Hu-Tucker [Byr73a, Byr73b, Yoh72]. Hulls [PH77, Pre79]. Human [Ers72b, Ers72a]. Humanities [Sed72]. Humanizing [Ste74]. Humble [Dij72, Ros76]. Hybrid [Sch78b]. HYDRA [WCC^{+}74]. Hyperbolic [Fri72, SM70b]. Hypergeometric [Kop74, Rel63]. Hypothesis [LS77, Nol73].

I. [Smi75b, Smi75a]. I/O [Pie75, Teo78, TCB78]. IBM [Dit76, CP78a, Dit76]. IBM/360 [Dit76, Dit76]. Ideal [SP74b]. Idealized [Teo78]. Ideas [GG77]. Identification [Ano70a]. II [Gen72c, Mor72a, Tad78b]. III [KC73a]. Ill [KZ75]. Ill-Chosen [KZ75]. Illiac [LLBR75, Mil73]. Illumination [Pho75]. Image [HLL78]. Images [BN76, Cro77, Hod70, HB78]. Immediate [PM72]. Immunology [Han72]. Impact [SC79, Col79]. Implementation [AC75b, BW73, CMM72, CW73, Dit76, Fen71c, Gim73, Har71b, HLL73, Kro74b, Kro74a, Lem79, WDT76, Bro77, RRD73, Dit76]. implementations [Jon86]. implemented [Cha72]. Implementing [Gen72b, Gen72c, SS72a, Sed78, Sed79]. Implications [Tan78]. Implicit [SS73, KJ85, Lin71]. Improved [Bra70c, LCW74, Mul72, Wym75, CR79, Wes79]. Improvement [Sch70b, DKHM76]. Improving [Bir77a, Cla77, Fer74, Gal79, Tho70]. In-Situ [CT77, Lea79, Mac72, LB70a]. including [Ano70a]. Incomplete [BB74, Ful72, Gau64a, Lud63, PS76, Sch78a]. Incorporating [Ste70c]. Incorporation [KL78]. Increasing [Whe73, van70a, van70b]. Increment [Ban77, Luc72]. Incremental [Bre77, DB76]. Incrementally [EC72]. Independent [BCS74, Mis73]. Index [BW72b, WS75]. Indexed [Mul72]. Indexed-Sequential [Mul72]. Indexes [Lum70, Mal78, MS77b]. Indexing [BCE77, SWY75]. Indices [Mul71]. Indirect [Dew75]. Individualizing [Kof72]. Induced [de 69]. Inducing [HRM78]. Induction [Mor71b, MW77]. Inductive [MNV73, Ski74]. Industry [POW^{+}72, WG75]. Inference [CRML73, SBS70, Ski74, SN73]. Informal [Wei72]. Information [BCLS74, BS79, CH76, CMM72, Cou73, Den76, DD77b, Flo76, GR73, HH70a, HH70b, Hua73, Luc78, MN73, Mot79, NKHS76, Pag73, Pre74, Sal73a, Sal74a, Shw74, Ste74, Taj77, Tan72, YS77, Ano70b, Ano70a, Ano70c, Ano70d, Ash72, Ben72b, Mor70a]. Information-Theoretic [BCLS74]. Inherently [Mau70]. Initial [ANS69]. Inline [Sch77]. innovation [DD77a]. Input [Eld70, Mil74, RB70]. Inquiry [NS76]. Insecure [Den79, Mer78]. Insertion [Hir76, Hol77, RZ79]. Insertions [Kos78]. Inspections [Mye78]. Instabilities [Cou75]. Instruction [Lun77, Kof72, Sto73]. instructional [OC71a, Zin72]. Instructions [Lam75, Szy78, War77]. instrumentation [SG70]. Integer [CG70, CT71, FM78a, JK72]. Integers [Bra70a, Bra70b, RG73]. Integral [Act74, Fri72, Han72, HP67, Lin72, eL76, Mer62, Ng70, Pac70, Pie73, Red70, Sko78, Wal74, Igu72, Ste73, HP85]. Integrals [Act74, Ein72a, Gau73, Hil70c, Pie74c, Ein72b]. Integrands [Hil70c]. Integrating [Fox72, MS70]. Integration [Bol72a, Bol72b, Bol72c, Bol72d, Gea71b, HMR72, Kli78a, Mos71b, Pat73, Pie74b, Sch70b, Wal70, Wil74, Lin71]. integrity [Han79]. Intellectual [Wag70]. intelligent [Wil75a]. Intentional [Min76]. intentionally [Par79]. interactions [ES70]. Interactive [AT77, BS70, BW76, yC74, FB73, Luc78, Pag72, SW77a, SW77b, SK75, WDT76, WAG77, Gra70, LR70, RB70, Smi70b]. Interchange [Ano70b, Fle66, Ito76, Ano70a, Ano70c, Ano70d, Ano70e]. Interface [SC75, Fra72]. Interference [BS76, Boy79, GC72, HRM78, Smi77]. Interleaved [BS76, BC75]. Intermittent [GD78, MW78]. Internal [DPR70]. international [Smo76]. Interpolating [LS74]. Interpolation [Aki72, Aki74a, And76, And79, Gus71a, Gus71c, HR73, Lag74, vdN79, PIA78, Str74b, Aki74b]. Interpolatory [Gus71b]. Interpretation [Dev77]. Interpreter [Man79]. Interprocedural [Bar78]. interprocess [Wal72b]. Interrupt [Zel71, Mor70a]. interrupts [Hil73b]. Intersection [KSW78]. Interval [Pat73, RR73a, RL71, You74]. Intrinsically [JOR75]. Introduction [Bir77a, Hor77, Kli79a, LT77, OC71b, Poo76]. Inventory [SBGN70]. Inverse [ET75]. Inversion [Ehr70a, KM76b, Ste70a, Vei74, Vei77]. Inverted [Car75]. Investigation [Ste79]. Investigations [SMMH77]. Isolating [Hec78]. Isolation [Wag70]. ISPL [Bal73]. Issue [BFS78, Wil75b]. Issues [Goo75, McC74, Was75]. items [Bay73c]. Iteration [AW77, SWL77]. IV [Mil73, Han74, LLBR75].

J6 [Dut76, Fei75, Čer71, Fre75, Lew73, Wat74, Wil72, Wri74]. Jacobi [Sko75b, Wit68]. Job [FB73]. Joining [BM76]. Jordan [PW75]. Jump [Shn78].

K. [RH72]. K2 [Ito76]. Keep [Ric71]. Kernel [Mil76, WCC^{+}74]. kernels [BP73]. Key [BW76, LYD71, Lum73, RSA78, LY72]. Key-to-Address [LYD71, Lum73, LY72]. Keys [She78]. Keyword [KSW78, Fra77]. Killed [Lyn70]. Kind [Gau64b, Lon70, Sko75a, Igu72, Bra70d]. Kolmogorov [Pom76, Pom74]. Konrad [BW72a]. Koz [KC73a]. Kutta [Tho70].

L [Sim74]. Label [Fen71c]. labels [Eva74]. laboratory [HW72]. Lance [Kim70]. Language [AW77, BF71c, Bro72b, CK70, Dit76, Gre77, HHKW77, HLL73, Hen72, HOW74, Hor77, JL78, Kes77b, PF71, Ray76, Rey70, SG75, Sch77, SBS70, Sch75, Sed72, SK75, Sha74, SHL75, ST73, Sim70, SLSH75, Wal78, Woo70b, WRH71, Zd77, CRM70, Iro70, LLBR75, Lee77, SP74b, AC75b]. Languages [CZ74, CRML73, EC72, GG77, KL78, Mau70, Mor73, Nau75, PA72, Ros72, Sam72, Sam76, Sch74, Sco77, Ten76, Wag74, Win79, BW72a, Lyo74]. Laplace [KM76b, Vei77, Ste70a, Vei74]. Large [Cod70, Lau75, LYD71, Mor78b, NS78, Ulr78, Wal78, LY72]. Larger [Gau65, Jan77]. Latency [GK74, SF73]. Lattice [Den76]. Lawler [de 70]. Layout [AKA71]. Layouts [Pfe75]. Learning [May79, SF71, Smi73]. Least [Art63, FS73, Kro74b, Kro74a, Sch72c, Sha75, Spä70b, Lyo74]. least-errors [Lyo74]. Left [vB76]. legal [Smo76]. Legendre [Bol72c, Bol72d, HMR72, Jan77, Gau65, Pie74a]. Lemniscate [Tod75a, Tod75b]. Length [Dit91, FM71, FM72, FS76, Jon72, Lit91, McC77, Pea90, Pea91, Sav91, Mon70]. Letter [Ral72]. Level [CW78, Fel79, GG77, Ham76, HHKW77, HLL73, Hir70, MPT77, Ros77, Sch75, She73, SHL75, SLSH75, Fai79]. Levels [Bro72b]. Lexicographic [Lei69, OS68, Roy73c]. LG [Ray76]. liberal [LRT77]. Liberated [Bac78]. Lifetimes [BB77]. like [Pan70]. Limited [Shw71, Shw74, Ver72, KJ74].

Limited-Entry

[Shw71, Shw74, Ver72, KJ74]. Line

[AG74, Chr72, Eld70, Ell74a, Ell74b, Gai74b, Hod70, JB74, KNO76a, Kli79b, MC73, TCCH71, Wil72, Wil73c, Gai74a]. Linear [Ban77, BK70, Bre77, CRSW74, yC74, Fia73, Fie70, Fis75, GW75, GW76, GM78, Hir75, HOK73, How71, How73b, KD71, KS77, Kub73, Lin75, Luc72, Mol72a, Obr73, Pro74, Rav72, Rav74, RR73a, SS68, Sal74b, DKHM76, LR70, MP74, Mar72c]. Lines [Bur77, Mat72, DH72]. linguist [Fra70]. linguistics [McM73]. Linkers [Ham76]. Linking [Bob75]. Lions [SS74b]. LISP [Roc71, Bak78b, CG77]. List [ACD74, Bak78a, Bob72, Che70, Cla76a, CG77, Cla78, Fen71a, Fen71b, Fis75, Lin74, McC74, Rei73a, Smi75c, VD75b, Wal72a, Was75, Sha70, VD75a]. List-Compacting [Wal72a]. List-Compaction [Fen71a]. Lists [FM71, FM72, Wis76]. Literate [Knu92]. Literature [ABE77]. Local [Aki72, Aki74a, And76, And79, DK78, MB76b, TP77, Aki74b, DD77a]. Localities [MB76a]. Locality [Fer74, CR72]. Localization [DH73, Squ72, Wil73a]. Locally [Mil74]. Locally-Organized [Mil74]. Locating [Hin75]. Locations [ML77]. locks [EGLT76]. Log [PIA78, Pur74]. Log-in [Pur74]. Logic [DK79, Kow79, Sco77]. ’logical [Ehr70b, IR78, KM76a, War77]. London [Red71]. Longest [HS77]. Lookup [Abr74, GP73, She78]. Loop [Weg74a, vB73, Eva74]. Loops [Lam74b]. losses [Par79]. Low [Ham76]. Low-Level [Ham76]. Lowe [SM70a]. LP [HW79]. LR [DeR71, HSU75, MM78c]. LRU [TS77]. LU [EH75]. Lucid [AW77].

M [GH79, Lai79, SD70]. M. [Sim74]. M1 [Bro76, Bro72a, Cha71a, FR75a, GR70, Lon70, Pet70]. Machine [Ben72a, Fin77, HLL73, Ivi77, KFM71, OS72, Tan78, Wil70a, Wir69, Wir70, Ano70e]. machine-to-machine [Ano70e]. machines [SP74b, Szy78]. Made [Kli79b]. magnetic [Ano70c, Ano70d]. Maintenance [LST78, Smi75c, Taj77]. majority [PW70]. Man [KFM71]. Man-Machine [KFM71]. Management [BW74, BBC77, CHI^{+}74, CH76, DK78, OC71b, SW72, Ash72, BBF^{+}70, Mor70a, NGMM70]. Manager [Nol77]. Managing [Fra73, Nol73]. Manchester [Lav78]. Maniac [Mor72a]. Manipulating [Vui78]. Manipulation [Ars79, Bro73, HT73, PM75, Ulr78, Hal71, Hea71, Jef71, Mor71a]. manipulative [BF71a]. Many [Bra75, Bur77]. Many-Sided [Bur77]. Map [AA71]. Mapping [ST73]. Maps [ML72, NW79, War78]. Mark [Lav78]. Markov [CO76b, CO76a, Ste78]. Marovich [Kar75]. Marriage [MW71b, MW71c]. Martin [DKHM76]. Mask [Dat72]. Masked [Fei75, Wat74]. Masking [GK74]. Mason [Hee70]. Mass [Bla70]. Match [BK73b, Sha77, Van74]. Matching [AC75a, CW78, Gal79, HRM78, KR81, KMP77, SA77]. Math [Mil75a, Mil75b]. Mathematical [Fos72a, HW79, CR72]. Mathematics [KC75, Nau75, Tra72]. Mathieu [Cle69, Fri72]. Matrices [BW70, Bir75, BL71b, Kha72, PM78, Ste70c, van75b, BL71a, Rok70, van75a]. Matrix [BS72b, BW72b, Bre73a, Chi78, DWT71, Ehr70a, ET75, FP79, GB68, Gus78, HKM80, Hei72, JH75, Kub73, LB70a, Law73, Lea79, Mac72, McN71, Mol72b, MK73, Rei73b, Sal73b, Sip77, Ste70b, Ste70d, EH75, Mar72c, Nie72]. Maxflow [Bay68, Bay73a, Bay73b]. Maximal [Hir75]. Maximum [OH72]. Mean [BCS74, Han75b, Cot75, Nel79, Wes79]. Meaning [Fra71]. measure [BED77, FW78a]. Measurement [AT77, NR71]. Measurements [BJW70]. Measures [CMM72]. Mechanical [Gus71c, Weg75]. mechanics [Jef71]. Mechanism [HG78]. Mechanisms [Gon76, LSAS77, Wym75]. Median [She78]. Medical [Luc78]. Meetings [Gri70]. Memories [SB77]. Memory [Bae72, BCPW74, BS76, BCD72, BL79, BBC77, BC75, Gif77, Her72, HMT79, Hir73, Mal78, MS77b, Nie77, OS72, RL74, Sho77, Sit78, Smi77, Wil71, Wil73b, Hil73b, Spa72]. memory-content-generated [Hil73b]. Merge [Bro72a, Bro74, Fer71]. Merge-Sorting [Fer71]. Merging [Bla70, Gav75]. Merit [Nie71b]. Mesa [GMS77]. Mesh [TK77]. Mesh-Connected [TK77]. Meshes [Nil70]. Method [AA71, AA79b, AA79a, Bat75, Bel70, BCE77, Byr73b, Cav74, CB74, Dat72, EC72, Ehr73b, ES74, HW79, JH75, Kim70, LB70b, Lip75, Mag75, MR72, Mul72, Pag74a, Pan70, PHH72, Rad74, RS72, RS74, RK73, RSA78, SP74a, SP75, She74, SS74b, Spr77, WGL76, WS75, Yoh72, ZM78, Aki74b, BF71b, Byr73a, Igu72, Kai70, Lin71, MP74, Mar72c, MD70, Tho70, Wes79, Wil70c]. Methodology [AT77, BS79, Gen72b, WD78]. Methods [AD72, Bré79, Gre72, HLL78, Lum73, MNV73, Mar76, SR78, OC71a]. Microcomputer [EL79a]. Microprogramming [RFE72, TF71]. Minicomputer [FFG73]. Minimal [CGL73, Ful74, GR70, Ker73, Mag75, Pag74a, Pet70, Sin69a, WGL76, Whi72, Mon70]. Minimal-Total-Processing [Ful74]. Minimax [Dun72, Wat70]. Minimization [Bul74, Dav76, Fie70, Gim74, KK74, MM73, de 70]. Minimizing [GBK73]. Minimum [Byr73b, FS75, Wag73b, Yoh72, Byr73a]. Minimum-Area [FS75]. Minimum-Space [Wag73b]. Minit [HOK73, KD71, Obr73, SS68]. Minorities [MM78a, MM76]. Miss [EF78]. Mixed [How73a]. mobile [Wai70]. Model [Alt79, Bab77, BW73, Car73, CO76a, Cod70, DS72, Den76, GK74, Led72, Man79, MS72, OS72, PF78, SWY75, Sal74b, SBGN70, She73, Sil76, Wil76, Ber70, CO76b, CR72, Str74a, DS73]. Modeling [SBB72, Ste78]. Models [Cla76c, FJ70, HB78, KS77, Sch78b, TCB78, Kim72, WG79]. Modification [Shw71, War75, BF71b, Pee83]. Modified [Bur74, Ful72, HW79, Pie74b, Sch78a, Wal70, Wen78]. Modularization [HFC76]. module [Par72b]. Modules [Par72a]. monitor [Str74a]. Monitors [Hoa74, How76, Kes77a]. Monotone [CW78]. Monotonic [Sal72c]. Monte [Gal73]. Moore [Gal79]. Moorer [Smo72]. Moratorium [Pre74]. Motion [BW76]. Moving [Cla76a, Rei73a]. MR [GH79, Lai79]. MU5 [IC78]. Muenster [Dit76]. Multi [HMT79, Lum70, Man79]. Multi-Attribute [Lum70]. Multi-Class [HMT79]. Multi-Interpreter [Man79]. multiaccess [BBF^{+}70]. Multiclass [BM76]. Multics [BCD72, SG70, Sal74a]. Multidimensional [Ben75, Gal73, Mul77]. Multifield [BCE77]. Multilevel [de 70]. Multiple [BW73, Eva74, Fen71b, Lam75, McG73, MB73, SC79, vB73, Mif70]. Multiple-Precision [MB73, Mif70]. Multiplexed [Gim74]. Multiplexing [Pac73]. Multiplication [Mot79]. Multiplications [Ber72, Gla71, BG72a]. Multipliers [Bra70a, RG73]. Multiply [BS73]. Multiply-Restricted [BS73]. Multipriority [BM76]. Multiprocessing [Ste75, Ste76]. Multiprocessor [BS76, BB78, Smi77, WCC^{+}74]. Multiprogrammed [She73]. Multiprogramming [BB78, Bri70, Bri72, Cou75, SBB72, Wir69, Wir70, Ham73, Smi70a]. Multiprogramming-Multiprocessor [BB78]. Multipurpose [SF71]. Multiset [DM79]. Multistep [Lin75, Lin71]. multistore [vP70]. Multivariate [BS72a]. Munkres [BL71a]. Münster [Dit76]. MUSE [MS72]. Music [Moo72, Rad74, Wex73, McM73, Smo72]. Mutual [Atk73, Bal74]. MUX [Chr72].

N [PIA78, SD70]. Names [Fra71]. National [Ano70c, Ano70d, McM78, McM79, FS79]. Natural [FW72, HR73, LS74, McK70b, Nau75, SG75, SBS70, SK75, Sim70, Woo70b]. Navigator [Bac73]. Near [Chi78, SF73, Ver72]. Near-Optimal [Chi78, Ver72]. Near-Optimality [SF73]. NEATER2 [CS70]. needs [Ham71a]. Negative [Bro71]. neither [Eva74]. Nested [EC72]. Nets [Mis73]. Network [CGL73, Cra74, Den79, Taj77, Woo70b, Wal72b, WG79]. Networks [Buz73, CP75, DK79, Kas63, Mam77, MB76b, ML77, NS78, Nie71b, RK75, SS72b, Ten77]. Neumann [Col78, Bac78, Her61]. NEWBASIC [Dwy72]. Newton [Hil70c, Lin71, Pan70]. Newton-like [Pan70]. Next [Bay77, Ral72]. Next-fit [Bay77]. No [Wil70b, GH79, Lai79]. Node [CGL73]. nodes [Pat71a]. Noise [Lyn70]. Noisy [Mon71]. Non [Wal74, van75b, van75a]. Non-Integral [Wal74]. Non-Normal [van75b, van75a]. Nondeterminacy [Dij75]. Nondeterministic [MPT77]. nonintegral [Che78]. Nonlinear [Bro67, Kro74b, Kro74a, Pan70, Rad71, Shr72, Wat70, Dun72, Igu72, Lin71, SW72]. Nonprocedural [AW77, Kes77b]. Nonrecursive [Che70, Rei73a, Wal72a]. Nonsingular [PM78]. nor [Eva74]. Nordsieck [BF71b]. Norm [BR74, FH75]. Normal [AD72, BS72a, Don73, HD73, HK72, Mot79, Pag74b, van75b, van75a]. Normalization [NT70]. Normalized [LS71]. Norms [Bau75]. Notation [Goo75, Fra77, Wir77b]. Note [Bob75, DG73, Dit91, FW78b, Gla71, Hua73, Ken72, KW75, Lam73, Lit91, Mal78, Mau70, Pea91, Phi71, Pom70, RH72, Roy77b, Sav91, Wal72a, Bae70, Bay73c, BF71b, EH75, GH79, Mon70, Roy77a, ST74, Sto74]. Notes [Bir77b]. notions [EGLT76]. NRZI [Ano70d]. Nucleus [Bri70]. Number [Bat75, Bav70, BS73, But70, De 73, EL79a, KC73b, Sin69b, Whi70a, WS74, Yao77, Bre74, Bro72c, Pag70, Pay70, Pro72]. Numbering [Kno74, Kno76b, Kno77]. Numbers [Bir75, GM78, Kli79b, Mai77, Mor78b, PH65, Sch70a, Sul72, Ulr78, Rok70]. Numerals [NT70]. Numerical [Abe74, BJ71, KM76b, Pat73, Ste70a, Ste78, Tra72, Vei74, Vei77, OC71a, Smi70b].

O [Pie75, Teo78, TCB78]. O1 [Ged79]. Oakford [GJ75, KM75a, KM75b]. Object [Mat72]. Objectives [Cav74]. Objects [Cha70a, Lev76, LT73]. Obstacles [LPW79]. Obtaining [RSA78]. Occurrences [KS77, ZC77]. Occurring [She78]. off [BF71b, Tho70]. off-step [BF71b]. offerings [LRT77]. offs [Blo70]. On-Line [Chr72, Eld70, Hod70, TCCH71]. On-Site [BW74]. On-the-Fly [DLM^{+}78]. One [Dit76, Gau65, Hir76, Kos78, Phi71, RZ79, Spä70a, ZM78, Pro71, Sal73d, Jan77]. One-Sided [Hir76, Kos78, Phi71, RZ79, ZM78]. One-Step [Dit76]. Online [Han77, GH79, Lai79]. Only [Pat72, Rip78]. Operating [CMMS79, Fab73, Gai72, HFC76, HRU76, Hoa74, LS76, Lau75, Lis72, Lyn72, PF78, WCC^{+}74, DR72]. Operations [HSU77, Sit70]. Operator [CK77]. Opportunity [Smi75c]. Optimal [CW78, Chi78, FKU77, Han77, Ken72, Lew78, MS77a, MS76, MPY79, Mil77, Mon71, Mor74, ML77, PBH78, Pie75, Pre79, PF76, RZ79, Sit78, Ver72, Wag73a, Wil70b, You74, ZM78, ADU71, BJ71, GH79, IM70, Lai79, Sto73]. Optimality [FM78a, Har72, SF73]. Optimization [MR79, RK73, She74, Wal76, BCR72]. Optimized [Sam78, WS74]. Optimizing [MM78b, MN72, She71, She72, SC75].

Optimum [Bla70, SS73, Shn73, Fol71]. Order [Chi78, Lei69, McK70b, OS68, Roy73c, SF79, Wag74, You74, Hos73]. Order- [Wag74]. Ordered [FM71, FM72, HKLS70b, HKLS70c, Sal73c]. Ordering [DG73, Kas63, Lam78, Sal72c, SN73, Ten77, RG71]. Orderings [DM79]. Orderly [PM78]. Ordinary [CB74, Gea71a, Gea71b, Nik73, Wil74, Lin71]. Organization [Car73, Gho72, Hua73, RL74, Sil76, Smi77, TF71, Wir69, Wir70, WC71, Mor70a]. Organizations [Tei71, Yao77]. Organized [BJ74, JB74, Mil74]. Organizing [Riv76]. Orientation [And78]. Oriented [HL77, LSWL75, MT71]. Origin [Ste70c]. orthogonal [Sal73d]. Other [KS77]. Output [Fis77]. Overa [LM73]. Overdetermined [BR74, FH75]. Overflow [Mul72, Bay73c]. Overhead [Bar77, KNO76a]. Overlap [Teo78, TCB78]. overlays [Spa72]. Overview [Bal73, KC79a, Low78].

P.D.E. [Fri72, SM70b]. P.E. [Ano70d]. Package [Gus78, HKM80, Law73, McN71, Sal73b, Sip77, WS74]. Packed [Lyo78a]. Packet [MB76b]. Packets [DM78]. Page [Boy74, FG74, Pri73, PF76, ADU71, Smi70a]. Page-On-Demand [Boy74]. Paged [BBMT72, RL74]. Pagination [McC77]. Paging [BGM70, Dem70, Fag76a, Fag76b, FP79, FGG78, Mol72b, Mor72a, OS72, Pom70, Roc71, Sal74b, TS77]. Paper [KM75a, Sim74, Hee70, SM70a, Smi75b, Smi75a]. Papers [DR72, Hor77, Mor71a, Poo76]. Parabolic [MR72]. Paradigms [Flo79]. Parallel [ACD74, Gav75, GC72, Gri77a, Hir78, HCS79, Kel76, Lam74b, Lip75, OG76, Pat72, TK77, TCB78, War77, Gri78]. Parallelism [Eve74]. Parameter [Pri73, Fra77]. Parameters [Mil77, SS73, Wal74, Che78, Kni70].

Parametric [Lev76]. Paraphrasing [Gol75]. Parser [MS73, Mil74, vP70]. Parser-Generating [MS73]. Parsers [Cro72, LaL77, Loe70, MM78c, Rip78]. Parsing [AJU75, Ber77, CR78, Dam70, Ear70, SKK76, Woo70a, Str73]. Part [Gus78, HKM80, LaL77, McN71, Sip77]. Partial [CK70, Cha71a, MR79, SN73]. Partition [FGG78, Whi70b]. Partitioned [GBK73]. Partitioning [Bab77, yC74, CT71, PA72, CR72]. Partitions [BS73, Ehr74, McK70b, Whi70a]. Partnership [Smi73]. PASCAL [Dit76, Dit76]. Password [MT79]. past [Ham71a]. patents [Gal71]. Path [DM78]. Paths [Fox75, LPW79, Min74, Law77]. Patrick [Sim74]. Pattern [AKA71, Cha73, KR81, Kli71, Lau73, SL71, Sla79, SKK76, HW72, KMP77, Les72]. Pattern-Matching [KR81]. Patterns [BF71c, CGM79, Cha71b, Gim73, Les79, Lev72, Bro77]. Payments [Lip79]. PDEL [CK70]. PDP [BBMT72]. PDP-10 [BBMT72]. Peak [Hor75]. Peaked [Hil70c]. pedagogic [SP74b]. Pei [Rok70]. Pencils [DWT71, Hei72]. Percent [Sho77]. Perfect [AA79b, AA79a, Spr77]. perforated [Ano70b]. Performance [CGM79, Car75, GD78, KFSK76, KD78, Loe74, LYD71, Lum73, Lyn72, Sal74b, SC75, BCR72, BED77, Kim72, LY72]. Performed [Wil71]. Permutation [BL79, Ive76, Roy77b, Roy77a]. Permutations [Cha70b, GP74, Kno76b, Lei69, OS68, RH72, Roy73c, Roy73b, Har71a, Pag70, Roy73a]. Perplexed [Mos71a]. Personal [Den79]. perspective [Ben72b, Lav78]. PERT [Ten77, Kas63]. Perturbations [van75a, van75b]. Petri [Mis73]. Petrov [HZ78]. PFF [CO76b, CO76a]. Ph.D [Pay73]. Ph.D. [FMP74, TC76, TC77a, TC77b]. Ph.D.s [TC79]. phase [Ano70c, Les72].

Phenomena [KS77]. Philosophy [Don76]. Phone [AG74]. Photocomposition [AK73]. Phrases [Wag73a, Wag73b]. physics [Hea71]. Pictorial [Les72]. Picture [Lev72, TP77]. Pictures [GH71, GH72, Lev76, Mon71, Pho75, SR78, DH72]. Piecewise [FS73]. Pivoting [PW75]. PL [CS70, CW73, Fra70]. PL/ [Fra70]. PL/I [CS70, CW73]. PL360 [Mil70]. Place [Bre73a, Lea79]. Planar [Cli74b, FKU77, Pre79]. Planar-Valued [Cli74b]. Plankalkül [BW72a]. Plankton [Cla75]. Planning [Ama71, Eas70, LPW79, WD78, Eas72]. Plays [Smi73]. Plot [Fei75, Wat74]. Plots [AG74, Lew73]. Plotter [Čer71, Wat73]. Plotting [Dor75, Dut76, Ell74a, Ell74b, Fre75, Gai74b, Hir70, MC73, Mas75, vSdH79, War78, Wil72, Wil73c, Wri74, Gai74a]. Point [Ber72, Bol72c, GM74, Gla71, Gre72, Kim70, KC73b, Lin70, Mal71, Mal72, MV72, BG72a, BF71b, Bol72d, HMR72]. Points [PH77, Sha77, Shn73]. Poisson [Sch70a]. Poker [Fin77]. police [Col79]. Policies [BM76, BB78, TP72]. Policy [BS71, KC79b, McC74, Was75]. Policy-Driven [BS71]. Political [CP75, HOR72, Kli78a]. Pollack [Shw71]. Polygon [SH74]. Polygonal [Bur77, NW79, Mon70]. Polygons [Bur77]. Polyhedral [LPW79]. Polynomial [JT72, Pat72, Sha75, Squ72, Wil73a, Wit74, DH73]. Polynomials [McK70a, Sch72b, Sko75b, Wit68]. Polyphase [She71, She72]. Population [Cla75]. Portable [Bro72b, CMMS79]. Position [MDB79]. Positive [Sal72c]. Positivity [Bau75]. Power [LP78, OH72]. Powers [Bro71]. Practical [Bar78, Fra73, GR75, BF73, Est73, Lyo74]. practicality [TT82]. Practice [Mil76]. Practitioner [SD76]. Precedence [Ber77, CGL73, Cra74, Rip78, DKHM76, IM70, Mar72c]. Precise [Abe74]. Precision [Abd70, MB73, Ric72b, Ric73, Mif70, Sal72a]. predicate [EGLT76]. Predicates [Weg74a]. Prediction [Mam77]. Predominators [PM72]. Preliminary [Eas72, Ger75]. Present [Ric72a, Sto73]. present-day [Sto73]. Presentations [Bou70]. Preserving [DEL78]. President [Ral72]. Prevention [Hol71, PHH72]. Price [CGM79]. Price/Performance [CGM79]. Pricing [Nie70]. Primary [Gel73, Gif77]. Prime [Bat75, De 73, GM78, Mai77, Sin69b]. Primes [Bra70c, Dun70]. Principle [Mor71b, Rey70]. Principles [ADU71, DR72]. Priority [Vui78, Jon86]. priority-queue [Jon86]. Privacy [Min76]. Probabilities [Bou74, Bou76, Han75a, Mar72a]. Probability [FG74, Hur71]. Probable [McC75]. Probable-Word-Proof [McC75]. Probe [AA79b, AA79a, Spr77]. Problem [BL71b, CG70, Cra73, Cro77, DWT71, Dev77, Fen73, GB77, Gre77, How73a, JOR75, KW75, KFM71, Lam74a, Lam73, McC74, McM79, MW71b, MW71c, MR72, NT74, Pag73, Par75, Pfe75, Poh72, Tes74, Van74, vdBM75, Was75, BL71a, Bro72c, EM72, JK72, Kor70, Les72, OH72, PW70, Pro71, SP74b, Sto74, McM78]. Problem-List [McC74, Was75]. Problem-Reduction [vdBM75]. Problems [CRSW74, Fia73, HW79, Kim70, Kro74a, LK72, Pro74, Rav72, Rav74, Smi75c, GS72, Hos73, Smi70b]. Procedure [AC76, Bou70, HG78, Pur74, Rip78, SC78, WG79]. Procedures [Aki72, Aki74a, And76, HR73, LS74, MW71b, New71, SS73, Aki74b, Kni70, And79]. Proces [Gif77]. Process [BBF^{+}70, FS76, Zin72]. Processes [Bri78, DLP79, GC72, Hab72, Hoa78b, Hoa78a, CR79]. Processing [ACD74, Bak78a, Bob72, Bur77, Eld70, Ful74, HHKW77, Lam75, Mul77, PF71, Sal72b, SW77a, SW77b, TCB78, HW72]. Processor [Lun77, SC79, TF71, Low70, MD70]. Processors [Gav75, Hil73b]. Produce [Bra70c, Dun70]. Produced [Sho75]. Product [Bol72a, Bol72b, Bol72c, Bol72d, HMR72]. Production [Fos72a, TC76, TC77a, TC77b, TC79, Led74]. productions [IM70]. Productivity [Chr78]. Products [Chi78, MR75, MK73]. Professional [CH76, MN73, Ash72]. Professionalism [Fin75]. Program [Abe74, AC76, Alt79, Ars79, AE73, ABB^{+}79, CZ74, Dor75, Dut76, Ell74a, Ell74b, Fei75, FWY70, FG74, Fre75, Gai74b, Gel73, Gel78, Hoa71, JP73, Kin76, KC73a, Lam79, LCW74, Mac73, MC73, MB76a, MW71a, MW78, Mas75, McG73, ML77, Mye78, SD70, SF71, Smi73, Wat74, Weg75, Wil72, Wil73c, Wir71, Wri74, Ber70, BT79, Fra70, Gai74a, Low70, MM70, NR71, vSdH79]. Program-Synthesizing [LCW74]. Programmer [Bac73, Ber71, Dij72, Ivi77]. Programmers [GW74, van77]. Programmiersprache [Dit76]. Programming [Bac78, BR75, Bob72, Bri78, CG70, CRSW74, yC74, Chr78, CRML73, DPR70, Dit76, Ers72b, Ers72a, Fel79, FWY70, Fia73, FM78a, Flo79, Fra71, HHKW77, HW79, Hen72, Hod70, HOK73, HWBC77, Hun70, KL78, Kes77b, Knu73, Knu74a, Knu74b, Knu92, KD71, KS77, Lam74a, Lem79, Low73, MR73, Mil75a, Mil75b, MPT77, Mor73, Nau75, Obr73, Pro74, Rav72, Rav74, RR73a, Ros72, Ros73, SS68, Sam72, Sam76, Sch77, Sco77, SMMH77, Shr72, Tan78, Ten76, TF71, WDT76, Win79, Wir77a, WRH71, Zel71, Zd77, BW72a, EM72, JK72, LR70, LLBR75, Lee77, Pro71, SW72, Wai70, van77]. Programs [Bac78, BB77, Bir77a, Boy74, Cav74, CH76, Cou73, DLP79, DD77b, Dij75, Fox72, Fra76, FH74, GR73, Gri77b, Gri77a, IR78, KM76a, Kel76, KJ73, Lip75, MNV73, MS77a, MN73, MR70, OG76, RL77, Sed78, Sed79, Shw71, Shw74, TCB78, Weg73, Ash72, Gri78, Mor70b, Pol71, Smo76]. Progress [ANS69]. Progressive [Han79]. Project [Ber71, BS70, CGL73]. Projection [BHH78, Kro74b, Kro74a]. Projections [Cha71b, Cha73, GH71, GH72, Lau73]. Proneness [LD76]. Proof [Hoa71, Lam79, Lon70, McC75, RL77, SLR78, Taj77]. Proofs [DLP79]. propagation [GK70a]. Properties [DS72, GM74, Lin75, Lip75, Mal72, MNV73, OG76, BJ71, DS73]. Property [Gho72, SF79]. Proposal [Pag72, Spa72]. Proposed [Ano70c, Ano70d, CH76, Goo75, HMS72, Shw71, DH76, McC75]. Protection [HRU76, Lon79, Min76, Mor73, Sal74a, SS72a, Smo76]. Protocol [Taj77]. Prover [SN73]. providing [BT79]. Proving [DM79, Gel78, Gri77a, HOW74, How76, Lip75, MNV73, MW78, Sam78, vdBM75, Gri78]. Proximity [DEL78]. Pseudo [PH65, Sul72, Bre74, Spa72]. Pseudo-Random [PH65, Sul72, Bre74]. Pseudochaining [HP78]. Pseudoinverses [Hes75]. pseudorandom [Pay70]. Psychology [May79]. Public [Lon79, McC74, RSA78, Was75]. Public-Key [RSA78]. Publications [Wei72]. Pulse [SKK76].

QR [Ste70c]. Quadratic [Ack74, Agn76, Bat75, Bel70, Day70, LB70b, Pro74, Rad70a, Rav72, Rav74, Shr72]. Quadrature [Gal73, Ged79, Gen72a, Goo73b, Gus71b, Gus71c, Hil70c, Lyn70, Pie73, Gen72b, Gen72c, Goo73a]. Quadric [Lev76]. Quantiles [Gol73, Hil70b, Kni75, eL79b, Hil81b]. Queries [BCKH75]. Query [AC75b, Han77, Wal78, GH79, Lai79]. Querying [Cas73]. Question [SBS70, Sim70]. Question-Answering [Sim70]. Questionnaire [Low73]. queue [Jon86]. Queueing [Buz73, RK75, WG79].

Queues [Kes77a, Vui78]. Queuing [She73]. Quicksort [Loe74, van70b, Sed78, Sed79, Whe73, van70a]. Quotient [BK70, Bel70, LB70b].

R. [GR70]. R201 [CH70]. Rademacher [HKLS70a]. Radix [Sou75]. Ramp [Lag74]. Random [But70, Cla77, FJ70, HK72, Kno70, Kno73, Pag74b, PH65, RS72, RS74, Sch70a, Sch72a, Sul72, Tad78a, Tad78b, Wal74, de 69, Bre74, Pro72]. Randomized [KR81]. Ranges [BW72b]. Raphson [Lin71]. Rapid [Bur77, Gus71a, Gus71b, Gus71c]. Raster [BJ74, JB74]. Ratio [BB74, Lud63, PS76]. Rational [Bro71, Hal71]. Ratios [EF78, Gau64a]. ray [Les72]. Re [HS78, Sny78a, Sny78b]. Re-examined [HS78, Sny78a, Sny78b]. Readers [CHP71]. Reading [Lam77]. Readings [Sto88]. Real [Bak78a, CMMS79, Ehr70a, GZ75, GB68, HB78, Mor76, Pat72, Pre79, Rei73b, Ste70b, Ste70d, Wad76, War78, Wir77a, Nie72]. Real-Time [CMMS79, Pre79, Wir77a]. realistic [BT79]. Realizable [War77]. Reallocation [Bay73d]. Reasoning [Rey79, Sho78]. Rebuttal [FMP74]. Rechenanlage [Dit76]. Recognition [SL71, SKK76, HW72, Les72]. Recombination [CT75]. Recommendations [ABB^{+}79, Cou73, MN73, Ash72, Ham71a]. Recommended [MN73]. Reconstruction [Cha71b, Cha73, FKU77, GH71, GH72, HLL78, Lau73]. Record [Bab77]. Recorded [Ano70c]. Records [Gho75, McC77]. Recovery [GD78, GR75, MM78c, Rip78]. Recovery-Only [Rip78]. Rectangle [FS75]. Rectangular [BL71b, Deu72, LB70a, Mac72, BL71a]. Recurrence [Act74, Sal73d]. Recursion [AS78, Bir77a, Bir77b, Mor71b, Wal76, Atk73, Bal74]. Recursive [GK70a, GJ79, MS77a]. Redistricting [HOR72]. Reduce [BCS74, Gri70]. Reduced [JB73]. Reducing [Bre73b, Sch74]. Reduction [CK77, Cra73, JH75, Lip75, Sha74, vdBM75]. Redundancies [MR79]. Redundancy [Byr73b, Gho75, Yoh72, Byr73a]. Reentrant [SH74]. Refection [Roy77b]. Refection-Free [Roy77b]. Reference [DS78, Flo76, Rey70, Sha77]. Referencing [Wis76]. Refinement [Wir71]. Reflected [BER76]. Reflection [BN76, Roy73b, Roy73a, Roy77a]. Reflection-Free [Roy73b, Roy73a, Roy77a]. Reflections [LS76]. Reformatter [CS70]. Regard [KSW78]. Region [yC74, SR73]. Regional [Nie71b]. Regions [Mer73]. Register [Fre74, HB78]. Registers [Mor78b]. Registration [Win71]. Regression [Shr72]. Regular [LaL77, Wag74]. Rejuvenating [FS79]. Related [KW75, Tei71]. Relation [de 69]. Relational [BCKH75, Cod70, SC75, Wal78]. Relations [Act74, CGL73, HSU77, War75]. Relationships [vdBM75]. relativity [BF71a]. Relaxation [HLL78, PR79]. Relevant [Gho75]. Reliable [Hor77]. Remark [And76, And79, AV73, Bay73b, BB74, Bou74, Bou76, Bro74, Bro76, Bul74, Byr73b, Col78, Cot75, Dav76, De 73, Dor75, DH73, Dut76, Ein72b, Ein74, Ell74a, Ell74b, Er85, Fei75, Fre75, Fri72, FH75, Gai74b, GK70b, Ged79, Goo73b, GR70, Gus78, Han75a, HKM80, Hei72, Hil81a, Hil81b, HP85, Hol77, HOK73, How73b, Ito76, Jan77, JK72, Ker73, KK74, KD71, Kop74, KM76b, Lau73, Law73, Lea79, eL76, eL79b, MC73, Mag75, Mar72b, Mas75, Mis75, Mor76, Nie72, Obr73, Pag74b, Pet70, Pie74c, Pie74b, PM75, PS76, Pom76, Pro72, Pro74, Rad71, Rav74, Red70, Rob72, Roy73c, Sal72a, Sal73b, Sch78a, Sch72a, Sch72b, Sch72d, Sch72c, She74, Sip77, Sko75b, Sko75a, Sko78, Ste70d, vSdH79, Sys75, Ten77, Tes74, Vei77, Vos73, Wat73, Whe73].

Remark

[WGL76, Wil73a, Wil73c, Wit74, Nel79]. Remarks [Abr74, Fri72]. reminiscences [Alt72]. Remote [FB73]. Removal [AS78]. Reorganization [FP79, MS76, Shn73]. Repacking [BCPW74]. Repeat [PKT73]. Repetitions [Cha70b]. Replacement [CO76a, PF76, ADU71, CO76b]. Reply [Kar75, FLB73, KJ74]. Report [ANS71, ABB^{+}79, FS79, Luc74, McC74, Ste74, Ash72, Eas72, Gal71]. Representation [AG74, Ano70e, Bur77, CH75, Mer73, WAG77]. Representations [Eas70, HW73]. Representative [SK74]. representing [Han74]. Requests [Pie75]. Required [MK73]. Requirement [Ber77]. Requirements [Bob72, FH74, Gif77, JB73, PG74]. Requiring [EKW74, Ulr78]. Research [Ard76, Luc78, Rad70a, RFE72, KC79b, Zin72]. Residency [Pri73]. Residue [How71, How73b, Rad70a]. Resolution [Min76]. Resource [BB78, Nol73, BBF^{+}70, Ham73, Wal72b]. Resources [Fra73, Nie70]. Response [BBC77, FMP74, HMT79, Mam77, RH72]. Responses [O’R72]. Restricted [BS73, Whi70a, Whi70b, Atk73, Bal74]. Restriction [SG75]. results [LY72]. Retrieval [Bre73b, Gho72, HH70a, HH70b, Luc78, Lum70, Mot79, Mul71, Sal73a, Sil76, Ski74, Tan72, Van74, Weg76, YS77]. Retrieval-Update [Mul71]. Retrieving [AA79b, AA79a, Spr77]. Reveal [GM74, Mal72]. Reverse [DM78]. Reversible [Zel73]. Review [Wei72]. Ricatti [Bra70d]. Right [LaL77, OSW78, vB76]. Rings [SS72a]. Ritz [BJ71]. River [Cla75]. role [Kim72]. Rollback [GD78]. rolls [Ano70b]. Root [Wil70b]. Roots [DWT71, DH73, Hei72, MS70, Squ72, Wil73a]. Rosary [RH72, Roy73b, Har71a, Roy73a]. Rosenbrock [Dav76, Bul74, KK74, MM73].

Roster [Sam76]. Rotating [Mat72]. Rotational [GK74]. Rotations [Fei75, Wat74]. round [Tho70]. round-off [Tho70]. Rounded [Lag74]. Roundoff [Tsa74]. Routine [Pro74, Rav72, Rav74, WS74]. Routines [Hil70c]. Routing [Agn76, Fen73, Tes74]. Rule [Dat72, RK75, Sho77]. Rules [CR76, Gus71b, Gus71c, Shw75, SN73]. Runge [Tho70]. Running [Gal79].

S [Wil75b, LS77]. S-Curve [LS77]. S13 [Gau73]. S14 [eL79b, PS76, Pom76, Sch78a, Ful72, HD73, Hil73a, Kuk72a, Kuk72b, LT71]. S15 [eL76, Don73]. S17 [Sko75a, Bur74, Gau64b]. S21 [Sko78]. S22 [Mor76, GZ75]. Salazar [GJ75, KM75a, KM75b]. Salesman [KFM71]. Samples [Pom74, Pom76]. Sampling [AD72, Fis76, Gal73, SBGN70]. Sard [BP73]. Satisfying [Sch70a]. Saturation [Cou75]. SBA [Zd77]. Scalar [Cli74b]. Scalar- [Cli74b]. SCALE1 [Lew73]. SCALE2 [Lew73]. SCALE3 [Lew73]. Scales [Lew73]. Scan [BJ74, JB73]. Scanned [Nol71]. Scanned-Display [Nol71]. Scanning [ML72, Wym75]. Scatter [Ban77, Bre73b, Day70, Luc72, Lyo78a, FLB73]. Scene [SF79, SR73]. Scheduler [BS71]. Schedules [ACD74]. Scheduling [BCS74, Ful74, Gri70, RF77, SO74, SBB72, SF73, TP72, Wil76, Win71, Bro72c]. Scheme [CT75, EKW74, Sal73d]. School [SO74, SS74b, JH75, LML74, NGMM70]. Schools [MT71]. Science [Ama71, And78, Ard76, AE73, Aus77, ABE77, ABB^{+}79, CH76, FS79, Kan72, Kan73, MM78a, MDB79, MM76, NS76, TC77a, TC77b, Tra72, BF73, Est73, LRT77, SW72, TC76, TC79, Knu72b]. Sciences [Sly74]. Scientific [Pag72, Ric72a]. Scientists [HMS72, Wag70, Fai79]. Search [Ack74, AC75a, Ban77, Ben75, DL78, Fie70, FM71, FM72, Jon72, Luc72, Mar76, MM78b, Mer73, vdN79, NS76, PIA78, RZ79, Riv76, Shn78, Sun90, Ten78, Tie70]. Searching [BCLS74, Ben75, BM77, BK73b, Day70, Knu73, LB70b, Sha77, Shn78, SD70, SL71, BYG92, TT82, WM92]. Second [ANS71, Bra70d, Igu72]. Secondary [Bel70]. Secrecy [EKW74]. Secret [Sha79]. Section [LT77]. Secure [Den76, DD77b, Den79, Mer78]. Security [CMM72, Mil76, MT79, PF78, Pur74]. Seek [Wil76]. Sees [POW^{+}72]. Segment [BJW70, BB77, DS78]. Segmentation [GBK73, Low70]. Segmentations [TP77]. SELECT [Bro76, Pri73, FR75a]. Selected [Hor77]. Selection [AA71, Car73, FR75b, FW72, Low78, PBH78, Yap76]. Selective [DL78]. Self [Dij74, Riv76, SC78, WG79]. Self-Assessment [SC78, WG79]. Self-Organizing [Riv76]. Self-stabilizing [Dij74]. Semantic [DK79, SS72b, vB76]. Semantics [Hen72, Mac73, PF71, Ten76, BLSW71, van72]. Semaphore [Lau75]. Sembegs [Wen78]. semi [CO76b, CO76a]. semi-Markov [CO76b, CO76a]. Semidirect [Ehr73b]. Seminars [Kan73]. Sensitive [Woo70a]. Sentence [Gol75]. separately [Pin70]. Sequence [MK73, Sal73c]. Sequencers [RK79]. Sequency [HKLS70c]. Sequential [Hoa78b, Hoa78a, Mul72, Riv76, SBGN70, Shn78, SL71, Ten78, WS75]. Serial [Bak78a, MPY79]. Series [Abd70, Bro71, Bro73, Fra76, Pie74a, PM75, Sal72a, BHH78]. Servers [Buz73]. Service [AT77, Kli78a]. Set [Cha70b, DS72, EB77, Ehr74, FM77, Gib75b, KW75, LSWL75, Lun77, Pri73, RR73b, Ber70, DS73, FM78b, Jon86, RRD73, Sha70, ST74]. Set-Partitions [Ehr74]. Sets [AA79b, AA79a, DS78, Fer74, Gim74, KW75, MR74, Min74, Mor72a, PH77, Spr77, Sto73]. Settings [PBH78]. Several [Hil70c]. Shaded [Cro77]. Shallow [Bak78b]. Shannon [Cha72]. Shape [Wal74, Che78].

Share [Sha79]. Shared [Cod70, NR71]. Sharing [BS71, BBMT72, McG73, Nie71a, RT74, Sal74a, BBF^{+}70, BCR72, MD70, Wal72b]. SHELLSORT [CH70]. Shift [Sit78]. Shifting [Bar77]. Shifts [Ste70c]. Short [Ric71]. Shortest [Fox75, Min74, SF73, Law77]. Shortest-Latency-Time-First [SF73]. Shrinking [Lev72]. Sided [Bur77, Hir76, Kos78, Phi71, RZ79, ZM78]. Sieve [GM78]. SIGGRAPH [Poo76]. signal [HW72]. Signature [Ham71b]. Signatures [RSA78, TT82]. Significant [BGP75, Tsa74]. Similar [Act74]. Simple [Chr72, DeR71, GP73, MS72, Rad74, Rey70, Rip78, Rok73, Sal74b, Han74]. Simplification [Mos71a]. Simplified [CT75]. simplifying [Pin70]. Simply [Wen78]. Simpson [Lyn70, Bol72b, Bol72c, Bol72d, HMR72]. Simulation [BGP75, Cla75, Fis77, FM77, Ham71b, KD78, Nie77, SBGN70, Sch78b, TS77, VD75b, Wym75, BCR72, FM78b, VD75a]. Simulations [Ten78, Ulr78]. Simultaneous [Bro67, Pan70, Rad71]. Since [ABE77]. Single [AA79b, AA79a, Spr77]. Singleton [GR70]. Singly [Bav70]. SITAR [SW77a, SW77b]. Site [BW74]. Situ [CT77, Lea79, Mac72, LB70a]. situations [Str73]. Six [Cli74a]. Size [Ack74, Bat75, BJW70, ST74]. Sizes [BB77]. Skeleton [BW76]. Skew [HW73]. SL5 [HG78]. Slow [Wil70b]. Small [AE73, Ben72a, Mor78b, Pom74, Pom76, SW77a, SW77b, LRT77]. Smallest [Bro76, FR75a]. Smirnov [Pom76, Pom74]. Smoker [Par75]. Smooth [Aki72, Aki74a, And76, And79, Aki74b]. Smoothing [LS74]. SNOBOL [Bro77]. SNOBOL4 [Gim72, Gim73]. Social [DLP79, Sly74, Kli78b]. Society [HMS72, Stu72]. Socio [CP75].

Socio-Economic-Political [CP75]. Software [BS70, Bro72b, Don76, Fos72a, Gal71, GCD73, GHM78, Hir70, Hor77, Ivi77, LST78, Sly74, SB77, BT79, Par72b]. Solid [Kno70, Lev76, Sch72a]. Solids [Boy79, Bra75]. Solution [BR74, BS72b, Bro67, CB74, Ein74, Fia73, FSC73, FH75, Gea71a, How71, How73b, Kim70, Lam74a, LK72, MR72, Nik73, Par75, Rad71, GS72, Ste73, SP74b]. Solutions [BH73, Cle69, Fri72, How73a, NT74, Sla79, Hos73]. solve [Smi70b]. Solver [Mol72a]. Solving [Ehr73b, GB77, KFM71, Pan70, Pfe75, Igu72, MP74]. Some [Abr74, BW70, BK73b, Chr78, GB77, GG77, Hen72, KJ73, Loe74, LS77, Lun77, Mai77, Nie71a, SR78, Sla79, vdBM75]. Sometime [MW78]. Sometimes [MW78]. Sort [Bro72a, Bro74, She71, She72]. Sorting [BCE77, BGM70, Cha71a, Eve74, Fer71, FW72, GR70, HPSS75, Hir78, Jon70, Knu73, MN72, Pet70, Poh72, Sin69a, TK77]. Source [Ars79, CS70]. SP [HWBC77]. SP/k [HWBC77]. Space [Blo70, BF71c, Eas70, GBK73, Hir75, MS76, Mor74, Mor78a, PF76, SWY75, vdBM75, Wag73a, Wag73b, Eas72]. Space-Efficient [Mor78a]. Space-Optimal [Wag73a]. Space/Time [Blo70]. Span [Szy78]. Span-Dependent [Szy78]. Spanning [Ker73, Mag75, Pag74a, Sep70, WGL76, Whi72]. Sparse [ET75, Gus78, HKM80, HSU77, Law73, McN71, Sal73b, Sip77, TY79]. Spatially [Gim74]. Spatially-Multiplexed [Gim74]. Special [Abd70, BFS78, LT77, Mos72, Sal72a, Wil75b]. Specific [Fis77]. Specification [Gre77, Par72b]. Specifying [BCKH75, SWL77]. Speed [Mis73, Mul71]. Spelling [Mor70b]. Sperry [BHH78]. Spherical [Col78, Her61]. Spite [Dij74]. Spline [BH73, HR73, Hos73]. Splines [Cli74a, Cli74b, ES74, LS74, Nil70]. Split [She78]. Spoken [Mil74]. Squank [HM72, Lyn70]. Square [Ehr73b, Gol73, MS70, Wil70b, BCKH75, Kni75]. Squared [HP67, HP85, eL76]. Squares [Art63, FS73, Kro74a, Sch72c, Sha75, Spä70b, Kro74b]. SSTF [Wil76]. Stability [Lin75, PW75]. stabilizing [Dij74]. Stable [MW71b, MW71c]. Stably [Han75b, Cot75, Nel79]. Stack [BW73, TS77]. Stage [LS77, Nol73, Wai70]. Standard [CL79, Han75b, Cot75, Nel79, Ano70c, Ano70d, DH76]. Standards [ANS69, ANS71, Lee77, Ros76]. Stanley [Ste74]. Start [EF78]. Starting [Wil70b, Igu72]. State [MT71, vdBM75]. State-Space [vdBM75]. Statement [CS70, EC72, Dij68]. Statements [JP73, Knu74b, Par75, PKT73]. States [Ano70a]. Static [AA79b, AA79a, Spr77]. Statistic [Pom74, Pom76]. Statistical [Fra76, KC73b, Sly74]. Statistics [EB77]. Status [MM78a, MM76, Gal71, Ham71a]. Step [Dit76, SR78, BF71b]. Stepwise [Wir71]. Stochastic [Sil76]. stopping [Lin71]. Storage [Bac72, Ben72a, Ber77, Bla70, Bre73b, Bur76, CH75, Day70, FP79, Gel73, Gho75, GR70, Hah74, Hin75, Hua73, IGK71, JB73, Luc78, LSWL75, MPY79, Mil77, Mor74, Mot79, Pet70, Rob77, Sal73a, She73, SP74a, SP75, Sho75, Sin69a, Tan72, Wag73b, Wil71, CR72, FLB73]. Storage-Efficient [CH75]. store [MM70]. Storing [TY79]. Stormy [Mos71b]. Strategic [WD78]. Strategies [Sho75]. Strategy [Sit78, Smi70a]. Stratified [Gal73]. Strength [CK77]. String [AC75a, BM77, Gal79, TT82]. Strings [AA71, DS78, Dit91, ER78, Lit91, Pea90, Pea91, Sav91, Han74, KMP77]. Structural [SKK76]. Structure [CG77, EC72, FM77, Low78, MPT77, Sch75, Vui78, WC71, FM78b]. Structured [Abr74, Bri72, GP73, Hol77, Knu74b, Lem79, PS75, RL77, Sch77, SS74a, Tan78, Weg73, van77, AC75b]. Structures [Bac72, Car75, Cas73, Cla78, CE70, Ear71, Fis75, Ger75, Gut77, Hol77, LM75, Les79, Lin74, Mil73, Rob77, SS74a, GJ79, Low70, vP70]. Structuring [Hoa74]. Student [Hil81a, Hil81b, eL79b, Dwy72, Hil70a, Hil70b, Hil73a]. Students [And78]. Studies [Fin77]. Study [Ard76, Car77, CG77, Fox72, KNO76a, KC73b, LD76, LYD71, CR72, GK70a, LML74, LY72]. Style [Bac78]. Subexpression [DG73, RG71]. Subgoal [MW77]. Sublanguage [BCKH75]. Subprograms [Cli74a, Fra77]. Subroutines [Bro73, PM75]. Subsequences [Hir75, HS77]. Subsets [Ehr74, Dun72]. Substitution [PR79, Sch77]. Substring [Boo73, Har71b, Sun90]. Sufficient [FM78a]. Summation [Abd70, Gre72, Lin70, Mal71, Wal71, Sal72a]. Supervisory [Gai72]. Suppression [MR79]. Surface [Aki74a, And79, Art63, Cla76c, Dor75, Dut76, Fre75, FKU77, HB78, Mas75, Sch72c, vSdH79, Wri74, WAG77, Aki74b]. Surfaces [Boy79, Cla76b, Lev76]. Survey [ABE77, Hal71, LRT77]. Switching [Cha72, MB76b]. symbol [Hea71]. Symbolic [Kin76, LS70, Mos71b, HZ78, Mor71a]. Symbols [NS76]. Symmetric [Cra73, HW73, McK70a, RS72, Rei73b, Sch72b, Ste70b, Ste70c, Ste70d, EH75]. Symposium [Kli79a, KC79a, DR72, Mor71a]. Synchronization [Hab72, Kes77a, RK79]. Synchronizing [Hil73b]. Syntactic [Ars79, GR75, Hor75, Wir77b, vP70]. Syntax [Lyo74, Mil70]. Syntax-Directed [Mil70, Lyo74]. Synthesis [Bra75, CR76, MW71a, Weg74a]. Synthesizing [Fre78, LCW74]. Synthetic [HB78, SS76, SK74]. System [BM76, Bal73, BR74, BS71, BBMT72, Boy74, BS70, BB78, Bri70, BW74, Bur76, Car73, CMMS79, CT75, EL79a, Eld70, Fab73, FB73, FH75, Gai72, Ger75, GS71, GCD73, Han77, HL77, Hin75, Hir70, Hoa74, Hod70, Hol71, HWBC77, HH70a, HH70b, IC78, ID72, KC75, Kno74, Kno76b, Kno77, Lam78, LS76, Lau75, Lip79, Lis72, Luc74, Luc78, Lyn72, MR73, MS73, Pac73, Pan70, Pfe75, RT74, Rus78, SW77a, SW77b, She73, SBB72, SKK76, TCCH71, Wal78, Wen78, WDT76, Win71, WCC^{+}74, Zd77, Ben72b, BBF^{+}70, BED77, Dwy72, Eas72, EGLT76, GH79, Hal71, Kim72, Kli78b, Lai79, OC71a, Pag70, RB70, Smo76, Wai70, Wal72b, CP78a]. System/370 [CP78a]. Systematic [AS78]. Systems [ACD74, Alt79, Bae72, BCLS74, BS76, Bob72, BS79, BC75, CGM79, CP78b, CB74, CH76, CMM72, Cou73, Cou75, Dij74, Fen71b, Fra71, Fri72, GD78, HFC76, HRU76, HMT79, Kub73, KC73b, Lyo78b, Man79, MN73, Min76, NKHS76, Par72a, PS75, PN77, PF78, Ros72, SC79, Sch78b, Sil76, Sim70, Ski74, SM70b, Ste74, Sto88, WRH71, Ash72, BF71a, BCR72, DR72, Fol71, Fra72, Lin71, MP74, Mor70a, Mor70b, SC78, WG79, Led74].

T [Smi75b, Smi75a]. T. [Smi75b, Smi75a]. Tab [PBH78]. Table [Bat75, Day70, Dia70, IR78, IGK71, Mar72b, SS76, Shw75, TY79, WDT76, Bay73c]. Tableless [Rob72, Sto70]. Tables [Ack74, Ban77, Bay73d, Bou74, Bou76, Cav74, Dat72, GR73, HP78, Han75a, KJ73, Lew78, Low73, Luc72, Lyo78a, Mar72a, Mot79, MR70, Shw71, Shw74, Ver72, KJ74, Pol71, Str73]. Tape [Eve74, Ano70b, Ano70c, Ano70d]. Tape-Sorting [Eve74]. task [Str74a]. Tasks [BCS74]. Tausworthe [Pay70]. Taylor [LS71]. Teach [FWY70]. Teacher [Dwy72, Smi75c]. Teacher/student [Dwy72]. Teaching [HWBC77, Lem79, Ros73, SP74b]. Technical [AK73]. Technique [AHS76, BRT79, Boo73, Car77, GP73, Hah74, HRM78, Hec78, Hir76, IGK71, She78, Shn78, SR73, Han74, IM70, Par72b].

Techniques [BR75, Bre73b, BW76, FP79, LYD71, Mor79, NT70, Nie71a, RL77, Roc71, SL71, Ste78, HZ78, LY72]. Technology [CH76, Col79]. Telecommunications [FFG73]. template [SA77]. Ten [Bro73, PM75]. TENEX [BBMT72]. Tension [Cli74a, Cli74b]. Tensor [MR75]. Term [YS77]. Terminal [HL77, RF72]. Terminal-Oriented [HL77]. Terminals [McG73]. Termination [DM79]. Test [Aus77, CRSW74, Gel78, Har71b]. Testing [Boo73, HSU75, Kin76, Mye78]. Tests [Loe74]. Text [AK73, BCLS74, Dit91, FB73, Lit91, PM78, Pea90, Pea91, Rub76, Sav91, SW77a, SW77b, Wag73b, BYG92, TT82, WM92]. Texture [BN76]. th [Bro76, FR75a, Fox75]. Their [Cha71b, FM78a, Gim73, GH71, GH72, KJ73, LaL77]. Theorem [HOW74, SN73, vdBM75]. Theorem-Prover [SN73]. Theorem-Proving [HOW74]. Theorems [DLP79, GB77]. Theoretic [BCLS74]. theoretical [Hea71]. Theory [Arb72, GR73, Gim73, MV72, Mos72, Shw74, WS74]. Thinning [Deu72]. Third [PG74]. Thoth [CMMS79]. Threaded [Bel73, Dew75]. Three [Bol72d, Bou70, Fei75, Les79, MW71b, PH77, Wat74]. Three-Dimensional [Bou70, Fei75, Les79, Wat74]. Three-point [Bol72d]. Throughs [Lem79]. Time [Bak78a, Bar77, BS71, Blo70, BBMT72, Bre73b, BBC77, BCS74, CMMS79, Cla77, Fis75, FR75b, FH74, Ful74, Gal79, GW74, Lam78, Mam77, McG73, Mor78a, MK73, Nie71a, Pre79, RT74, SF73, Wad76, War78, Wir77a, BCR72, MM70, MD70, NR71]. Time- [Mor78a]. time-shared [NR71]. Time-Sharing [BS71, BBMT72, McG73, Nie71a, RT74, BCR72, MD70]. Times [HMT79]. Timetable [NT74, Smi75c]. Timetables [SS74b, JH75]. timing [Low70]. title [Dit76]. TL [FM78b]. today [BW72a].

Toned [Bou70]. tool [Han79]. Tools [Don76]. Topological [Kas63, Ten77]. Topology [Taj77]. Total [Ful74]. Trace [SBB72]. Trace-Driven [SBB72]. Tracing [Fen71b]. track [Ano70d]. track-200 [Ano70d]. Tracking [Kli78a]. Trade [Blo70]. Trade-offs [Blo70]. Tradeoffs [Mul71]. Traditional [Rad74]. Traffic [FJ70]. Training [Ber71, BT79]. Transactions [Ste79]. Transcendental [Ein74, FSC73]. Transfer [Sit78, Kli78b]. Transform [KM76b, LYD71, Vei74, LY72, Vei77]. Transformation [CW78, Cra74, Lum73, DH72, Gen72c]. Transformations [HW79]. Transforms [Ars79, Ste70a]. Transient [EB77]. Transient-Free [EB77]. Transition [FG74, Sch74, Woo70b]. Transitive [War75]. Transitivity [MR74]. Translation [Dia70, Mar72b, SHL75, SLSH75, Ver70b, Ver70a]. translator [ES70]. Transmission [AG74]. Transparency [PS75]. Transposition [Bre73a, CT77, LB70a, Lea79, Mac72, Roy73b, Roy73a]. Trapezoid [Sch70b]. Trapezoidal [Bol72a]. Traveling [KFM71]. Traversal [DL78]. Treating [BF71c, CRM70]. Treatment [Weg74b]. Tree [BS77, Cas73, Hu72, Mag75, Pag74a, Sep70, SD70, SL71, Weg73, WGL76, Whi72, Ker73]. Tree-Balancing [BS77]. Tree-Searching [SD70]. Tree-Structured [Weg73]. Trees [Ben75, DL78, Fos73, Hir76, Hur71, KFSK76, Ken72, Kno77, Kos78, Lou76, LP78, MM78b, MN72, McC77, OSW78, RZ79, She78, Tan72, Weg76, ZM78, HS78, Sny78b, Sny78a]. Treesort [Flo64, Lon70]. Trefftz [BJ71]. Trial [WS74]. Triangular [Deu72]. Tridiagonal [Kha72, Kub73, Rei73b, Ste70c, MP74]. Tridiagonalization [GP74]. Tries [Mal76]. Tucker [Byr73a, Byr73b, Yoh72]. tutor [Kof72]. twenty [Rev72]. twenty-five [Rev72]. Two [Bir75, Bol72c, BF71c, CZ74, Fle66, HMR72, Ito76, Kim70, LT77, MPT77, PH77, Ral72, She73, TCCH71, Ver72, Ehr70b]. Two-Dimensional [BF71c, TCCH71]. Two-Level [MPT77, She73]. Two-Point [Bol72c, Kim70, HMR72]. Two’s [War77]. Two’s-Complement [War77]. Type [Bol72a, Bol72b, Bol72c, Bol72d, Gan77, HMR72, Led72, Gim72]. Typeless [Rey70]. Types [Fen71b, GG77, Gut77, GHM78, Weg74b]. Typesetting [KC75].

U.S [Lip79]. U.S. [CGM79].

Undergraduate [ABB^{+}79, Cou73].

Undergraduates [Kan73]. Understand [MS72]. understander [Wil75a]. Understanding [Ear71]. Undirected [BK73a]. Uniform [Hol77, Kno70, Nil70, Sch72a]. Unifying [RF77]. Unitary [McK70a, Sch72b]. United [Ano70a]. Units [GM74, KL78]. Univac [BHH78]. Universität [Dit76]. University [Dit76]. UNIX [RT74]. unnecessary [Wir77b]. Unrecorded [Ano70d]. Update [Mul71]. Updating [Han75b, HW79, Wes79, Cot75, Nel79]. Upper [Mai77, Yap76]. Urban [Stu72]. Usage [Fre74]. Use [CRML73, DH72, KJ73, KZ75, Low73, Luc78, Mot79, PS75, Rad70a, Tei71, TS77, Col79, LML74, Smi70b]. Used [Ben75, Lyn70, Par72a]. User [EKW74, McG73, NR71, Nol77]. User/Manager [Nol77]. Using [BC75, CP78b, Cla76a, Cli74a, Cli74b, CE70, FFG73, Fin77, Fis75, Ham71a, HW79, HB78, How71, How73b, Lin74, Lum73, Mul71, Mul72, NS78, PR79, Pri73, SKK76, Tan72, TP77, Wil76, YS77, Atk73, Bal74, BF71b, CR72, Dwy72, Eva74, HZ78, RB70]. Usually [GW75, GW76]. Utility [Pre74, SMMH77]. Utilization [DK78, Mor72a, Wil71, BED77].

V [SC78]. Validation [GHM78, Mil76].

Value [And78, Kim70, Kli78b, Hos73]. Valued [Cli74b]. Values [Cle69, Fri72, Hur71, KC79b, Wil70c]. Variable [Dit91, FGG78, FS76, GP73, Kro74b, Kro74a, Lit91, McC77, Pea90, Pea91, PF76, Ric73, Sav91]. Variable-Length [McC77, Pea90]. Variable-Precision [Ric73]. Variable-Space [PF76]. Variables [Abr74, Fen71c, Fia73, FJ70, HK72, Pag74b, RS72, RS74, Tad78a, Tad78b]. variance [Wes79]. Variates [Wal74, Che78]. Variation [Bur76, Jon70, Kim70]. Vector [SWY75]. Vectors [Kno70, Sch72a]. Venus [Lis72]. Verbal [O’R72]. Verhelst [KJ74]. Verification [Fab73, Kel76, PF78, SWL77]. Verifying [OG76]. Versions [Lie76]. Versus [Ehr70a, Ehr73b]. Vertices [Bré79, SF79]. Very [HHKW77, Sch75, Sun90]. VI [WG79]. Via [Fre74, CO76a, CO76b, ES74, SK75, Spa72, Str73]. Vicious [Kan72]. View [Fos72b, Spr72, SW72]. Views [LT77]. Virtual [Bae72, BCD72, Mal78, MS77b, Spa72]. Virtualizable [PG74]. Visible [Cla76c, Dor75, Dut76, Fre75, Mas75, vSdH79, Wri74]. VMIN [PF76]. Voice [Eld70]. Volterra [Igu72]. Vs [EF78]. Vulnerabilities [Par79].

Walk [Lem79]. Walk-Throughs [Lem79]. Walkthroughs [Mye78]. Walkthroughs/Inspections [Mye78]. Wallace [Hee70]. Walsh [HKLS70b, HKLS70c]. Warm [EF78]. Warm-Start [EF78]. Warshall [War75]. Wasted [GBK73]. Wave [Gun67, Vos73]. Waveform [SKK76]. Waveforms [Hor75]. Waves [SKK76]. Way [Low73]. Weighted [Ban77, Lou76, Luc72, SP74a, SP75]. Weights [Gus71b]. Welfare [Kli78a]. Western [Spr72]. Which [Smi73]. While [PKT73]. Width [Kli71]. Wiener [Ste73]. Wijngaarden [Mor72b]. winning [Cha72].

Winograd [Har72]. Within [TCB78, Bay73c]. Without [Par75, Sou75, vB73, Dat72]. Woman [MM78a]. Women [MM76, WG75]. Word [Lam75, McC75, War77, Wil70a, IGK71]. Word-Parallel [War77]. Workbench [Ivi77]. Working [DS72, DS78, EB77, Fer74, Mor72a, Pri73, RR73b, Ber70, DS73, RRD73, ST74]. Working-Set [EB77, DS73]. Workload [SK74]. Workspace [Cla76a, Fis75, Lin74]. Worse [Gal79]. writable [Spa72]. Writers [CHP71]. Writing [Lam77]. WYLBUR [FB73].

x [Sal72c, Les72]. X-ray [Les72]. X3.9 [Ame66]. X3.9-1966 [Ame66].

Years [Ral72, Rev72].

Z [Er85, GH79, Lai79, Mis75, WGL76, Cha73, Lau73, Pag74a, Sch72b, Wag73a, Yoh72]. zero [Pro71]. zero-one [Pro71]. Zeros [JT72, Pat72, Rok73, Wit74]. Zuse [BW72a].

References

Amidon:1971:ASB

[AA71] E. L. Amidon and G. S. Akin. Algorithmic selection of the best method for compressing map data strings. Communications of the ACM, 14 (12):769–774, December 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:1979:CPH

[AA79a] M. R. Anderson and M. G. Anderson. Comments on perfect hashing functions: A single probe retrieving method for static sets. Communications of the ACM, 22(2):104–105, February 1979. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). Corrects errors in [Spr77]. See corrigendum [AA79b].

Anderson:1979:CCP

[AA79b] M. R. Anderson and M. G. Anderson. Corrigendum: “Comments on Perfect Hashing Functions: A Single Probe Retrieving Method for Static Sets”. Communications of the ACM, 22(10):553, October 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Corrects [AA79a].

Austing:1979:CRU

[ABB^{+}79] Richard H. Austing, Bruce H.

Barnes, Della T. Bonnette, Gerald L. Engel, and Gordon Stokes. Curriculum ’78: Recommendations for the undergraduate program in computer science — A report of the ACM curriculum committee on computer science. Communications of the ACM, 22(3):147–166, March 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abdali:1970:AAS

[Abd70] S. K. Abdali. ACM Algorithm 393:

Special series summation with arbitrary precision. Communications of the ACM, 13(9):570, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sal72a].

Aberth:1974:PNA

[Abe74] Oliver Aberth. A precise numerical analysis program. Communications of the ACM, 17(9):509– 513, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Austing:1977:SLC

[ABE77] Richard H. Austing, Bruce H. Barnes, and Gerald L. Engel. A survey of the literature in computer science education since Curriculum ’68. Communications of the ACM, 20(1):13–21, January 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abrahams:1974:SRL

[Abr74] Paul W. Abrahams. Some remarks on lookup of structured variables. Communications of the ACM, 17 (4):209–210, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aho:1975:ESM

[AC75a] Alfred V. Aho and Margaret J. Corasick. Efficient string matching: an aid to bibliographic search. Communications of the ACM, 18 (6):333–340, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Astrahan:1975:ISE

[AC75b] Morton M. Astrahan and Donald D. Chamberlin. Implementation of a Structured English Query Language. Communications of the ACM, 18(10): 580–588, October 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://dblp.uni-trier.de/
                                        db/journals/cacm/AstrahanC75.

html. Also published in/as: 19 ACM SIGMOD Conf. on the Management of Data, King(ed), May.1975.

Allen:1976:PDF

[AC76] Frances E. Allen and John Cocke.

A program data flow analysis procedure. Communications of the ACM, 19(3):137–147, March 1976. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adam:1974:CLS

[ACD74] Thomas L. Adam, K. Mani Chandy, and J. R. Dickson. A comparison of list schedules for parallel processing systems. Communications of the ACM, 17(12):685–690, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ackerman:1974:QSH

[Ack74] A. Frank Ackerman. Quadratic search for hash tables of size p^{n}. Communications of the ACM, 17 (3):164, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Acton:1974:RRF

[Act74] Forman S. Acton. Recurrence relations for the Fresnel integral ∫_{0}^{∞}[exp(−ct)dt/(t)^{1/2}(1 + t^{2})] and similar integrals. Communications of the ACM, 17(8):480–481, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ahrens:1972:CMS

[AD72] Joachim H. Ahrens and Ulrich Dieter. Computer methods for sampling from the exponential and normal distributions. Communications of the ACM, 15(10):873–882, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aho:1971:POP

[ADU71] A. V. Aho, P. J. Denning, and J. D. Ullman. Principles of optimal page replacement. Communications of the ACM, 14(1):80–93, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Austing:1973:CSC

[AE73] Richard H. Austing and Gerald L. Engel. A computer science course program for small colleges. Communications of the ACM, 16(3): 139–147, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adams:1974:CRC

[AG74] John Adams and John Gary. Compact representation of contour plots for phone line transmission. Communications of the ACM, 17(6): 333–336, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agnew:1976:QAR

[Agn76] Carson E. Agnew. On quadratic adaptive routing algorithms. Communications of the ACM, 19(1): 18–22, January 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Artzy:1976:FDT

[AHS76] Ehud Artzy, James A. Hinds, and Harry J. Saal. A fast division technique for constant divisors. Communications of the ACM, 19(2): 98–101, February 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aho:1975:DPA

[AJU75] Alfred V. Aho, Stephen C. Johnson, and Jeffrey D. Ullman. Deterministic parsing of ambiguous grammars. Communications of the ACM, 18 (8):441–452, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Alt:1973:CPT

[AK73] Franz L. Alt and Judith Yuni Kirk. Computer photocomposition of technical text. Communications of the ACM, 16(6):386–391, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Describes the typesetting system used by the American Institute of Physics.

Aramaki:1971:AEP

[AKA71] Itaru Aramaki, Tomokazu Kawabata, and Kazuhiko Arimoto. Automation of etching-pattern layout. Communications of the ACM, 14 (11):720–730, November 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Akima:1972:AAI

[Aki72] Hiroshi Akima. ACM Algorithm 433: Interpolation and smooth curve fitting based on local procedures [E2]. Communications of the ACM, 15(10):914–918, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [And76].

Akima:1974:AAB

[Aki74a] Hiroshi Akima. ACM Algorithm 474: Bivariate interpolation and smooth surface fitting based on local procedures [E2]. Communications of the ACM, 17(1):26–31, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [And79].

Akima:1974:MBI

[Aki74b] Hiroshi Akima. A method of bivariate interpolation and smooth surface fitting based on local procedures. Communications of the ACM, 17(1):18–20, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Alt:1972:ACR

[Alt72] Franz L. Alt. Archaeology of computers — reminiscences, 1945–1947. Communications of the ACM, 15 (7):693–694, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Alter:1979:MAF

[Alt79] Steven Alter. Model for automating file and program design in business application systems. Communications of the ACM, 22(6):345– 353, June 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Amarel:1971:CSC

[Ama71] Saul Amarel. Computer science:

A conceptual framework for curriculum planning. Communications of the ACM, 14(6):391–401, June 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

ANSI:1966:AF

[Ame66] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. ANSI Fortran X3.9-1966, 1966. ??–?? pp. Approved March 7, 1966 (also known as Fortran 66). See also subsequent clarifications [ANS69] and [ANS71].

Anderson:1976:RIS

[And76] Michael R. Anderson. Remark on “Algorithm 433: Interpolation and Smooth Curve Fitting Based on Local Procedures [E2]”. ACM Transactions on Mathematical Software, 2(2):208, June 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Aki72].

Anderson:1978:VOC

[And78] Ronald E. Anderson. Value orientation of computer science students. Communications of the ACM, 21 (3):219–225, March 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:1979:RBI

[And79] M. R. Anderson. Remark on “Algorithm 474: Bivariate Interpolation and Smooth Surface Fitting Based on Local Procedures”. ACM Transactions on Mathematical Software, 5(2):241, June 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Aki74a].

Anonymous:1970:ISU

[Ano70a] Anonymous. Identification of States of the United States (including the District of Columbia) for information interchange. Communications of the ACM, 13(8):514–515, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1970:IRP

[Ano70b] Anonymous. Interchange rolls of perforated tape for information interchange. Communications of the ACM, 13(1):56, January 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1970:PANa

[Ano70c] Anonymous. Proposed American National Standard. recorded magnetic tape for information interchange (1600 CPI, phase encoded). Communications of the ACM, 13 (11):679–685, November 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1970:PANb

[Ano70d] Anonymous. Proposed American National Standard. unrecorded magnetic tape for information interchange (9 track-200 and 800 CPI, NRZI and 1600 CPI, P.E.). Communications of the ACM, 13(11): 686–692, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1970:RCD

[Ano70e] Anonymous. Representation for calendar date for machine-to-machine data interchange. Communications of the ACM, 13(1):55, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

ANSI:1969:CFS

[ANS69] ANSI Subcommittee X3J3. Clarification of Fortran standards — initial progress. Communications of the ACM, 12(5):289–294, May 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66].

ANSI:1971:CFS

[ANS71] ANSI Subcommittee X3J3. Clarification of Fortran standards — second report. Communications of the ACM, 14(10):628–642, October 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66].

Arbib:1972:TAT

[Arb72] Michael A. Arbib. Toward an automata theory of brains. Communications of the ACM, 15(7):521– 527, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Arden:1976:CSE

[Ard76] Bruce W. Arden. The computer science and engineering research study (COSERS). Communications of the ACM, 19(12):670–673, December 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arsac:1979:SSS

[Ars79] Jacques J. Arsac. Syntactic source to source transforms and program manipulation. Communications of the ACM, 22(1):43–54, January 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arthurs:1963:AAL

[Art63] T. D. Arthurs. ACM Algorithm 176: Least squares surface fit. Communications of the ACM, 6(6):313, June 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sch72c].

Auslander:1978:SRR

[AS78] Marc A. Auslander and H. Raymond Strong. Systematic recursion removal. Communications of the ACM, 21(2):127–134, February 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ashenhurst:1972:RAC

[Ash72] Robert L. Ashenhurst. A report of the ACM curriculum committee on computer education for management. curriculum recommendations for graduate professional programs in information systems. Communications of the ACM, 15(5):363– 398, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abrams:1977:MIC

[AT77] Marshall D. Abrams and Siegfried Treu. A methodology for interactive computer service measurement. Communications of the ACM, 20 (12):936–944, December 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Atkins:1973:MRA

[Atk73] M. Stella Atkins. Mutual recursion in Algol 60 using restricted compilers. Communications of the ACM, 16(1):47–48, January 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Austing:1977:GAT

[Aus77] Richard H. Austing. The GRE advanced test in computer science. Communications of the ACM, 20 (9):642–645, September 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andrejkova:1973:RCA

[AV73] G. Andrejkova and J. Vinar. Remark and certification on “Algorithm 404: Complex gamma function”. Communications of the ACM, 16(8):489, August 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ashcroft:1977:LNL

[AW77] Edward A. Ashcroft and William W.

Wadge. Lucid, a nonprocedural language with iteration. Communications of the ACM, 20(7):519– 526, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Babad:1977:RFP

[Bab77] Jair M. Babad. A record and file partitioning model. Communications of the ACM, 20(1):22–30, January 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bachman:1972:ESS

[Bac72] Charles W. Bachman. The evolution of storage structures. Communications of the ACM, 15(7):628– 634, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bachman:1973:PN

[Bac73] Charles W. Bachman. The programmer as navigator. Communications of the ACM, 16(11):653– 658, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This is the 1973 ACM Turing Award Lecture.

Backus:1978:CPL

[Bac78] John Backus. Can programming be liberated from the von Neumann style? A functional style and its algebra of programs. Communications of the ACM, 21(8):613–641, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.stanford.edu/class/cs242/
readings/backus.pdf.  Repro-

duced in Selected Reprints on Dataflow and Reduction Architectures, ed. S. S. Thakkar, IEEE, 1987, pp. 215–243, and in ACM Turing Award Lectures: The First Twenty Years, ACM Press, 1987, pp. 63–130.

Baecker:1970:NDB

[Bae70] H. D. Baecker. A note on data base deadlocks. Communications of the ACM, 13(7):451, July 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baecker:1972:GCV

[Bae72] H. D. Baecker. Garbage collection for virtual memory computer systems. Communications of the ACM, 15(11):981–986, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baker:1978:LPR

[Bak78a] Henry G. Baker, Jr. List processing in real time on a serial computer. Communications of the ACM, 21 (4):280–294, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Originally appeared as MIT Artificial Intelligence Laboratory Working Paper No. 39, February 1977.

Baker:1978:SBL

[Bak78b] Henry Givens Baker, Jr. Shallow binding in Lisp 1.5. Communications of the ACM, 21(7):565– 569, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Balzer:1973:OIC

[Bal73] Robert M. Balzer. An overview of the ISPL computer system design. Communications of the ACM, 16 (2):117–122, February 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Balfour:1974:AAM

[Bal74] A. Balfour. An alternative approach to mutual recursion in Algol 60 using restricted compilers. Communications of the ACM, 17(4):210, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bandypadhyay:1977:CWI

[Ban77] S. K. Bandypadhyay. Comment on weighted increment linear search for scatter tables. Communications of the ACM, 20(4):262–263, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barth:1977:SGC

[Bar77] Jeffrey M. Barth. Shifting garbage collection overhead to compile time. Communications of the ACM, 20 (7):513–518, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barth:1978:PID

[Bar78] Jeffrey M. Barth. A practical interprocedural data flow analysis algorithm. Communications of the ACM, 21(9):724–736, September 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Batagelj:1975:QHM

[Bat75] Vladimir Batagelj. Quadratic hash method when the table size is not a prime number. Communications of the ACM, 18(4):216–217, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bauer:1975:PN

[Bau75] Friedrich L. Bauer. Positivity and norms. Communications of the ACM, 18(1):9–13, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder.

Bavel:1970:NAS

[Bav70] Z. Bavel. On the number of automorphisms of a singly generated automaton. Communications of the ACM, 13(9):574–575, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bayer:1968:AAM

[Bay68] G. Bayer. ACM Algorithm 324:

Maxflow. Communications of the ACM, 11(2):117–118, February 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Bay73b].

Bayer:1973:M

[Bay73a] G. Bayer. Maxflow. Communications of the ACM, 16(5):309, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bayer:1973:RAM

[Bay73b] G. Bayer. Remark on “Algorithm 324: Maxflow”. Communications of the ACM, 16(5):309–??, 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Bay68].

Bays:1973:NWC

[Bay73c] Carter Bays. A note on when to chain overflow items within a directaccess table. Communications of the ACM, 16(1):46–47, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bays:1973:RHC

[Bay73d] Carter Bays. The reallocation of hash-coded tables. Communications of the ACM, 16(1):11–14, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bays:1977:CNF

[Bay77] Carter Bays. A comparison of nextfit, first-fit, and best-fit. Communications of the ACM, 20(3):191–192, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bosten:1974:RAI

[BB74] Nancy E. Bosten and E. L. Battiste. Remark on “Algorithm 179: Incomplete beta ratio”. Communications of the ACM, 17(3):156–157, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Lud63].

Batson:1977:SSL

[BB77] Alan P. Batson and Robert E. Brundage. Segment sizes and lifetimes in Algol 60 programs. Communications of the ACM, 20(1): 36–44, January 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brice:1978:FCR

[BB78] Richard S. Brice and James C.

Browne. Feedback coupled resource allocation policies in the multiprogramming-multiprocessor computer system. Communications of the ACM, 21(8):678–686, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:1977:MMR

[BBC77] R. M. Brown, James C. Browne, and K. Mani Chandy. Memory management and response time. Communications of the ACM, 20 (3):153–165, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Betourne:1970:PMR

[BBF^{+}70] C. Betourne, J. Boulenger, J. Ferrie, C. Kaiser, S. Krakowiak, and J. Mossiere. Process management and resource sharing in the multiaccess system ESOPE. Communications of the ACM, 13(12): 727–733, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bobrow:1972:TPT

[BBMT72] Daniel G. Bobrow, Jerry D.

Burchfiel, Daniel L. Murphy, and Raymond S. Tomlinson. TENEX, A paged time-sharing system for the PDP-10. Communications of the ACM, 15(3):135–143, March 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burnett:1975:AIM

[BC75] G. J. Burnett and Edward G. Coffman, Jr. Analysis of interleaved memory systems using blockage buffers. Communications of the ACM, 18(2):91–95, February 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bensoussan:1972:MVM

[BCD72] A. Bensoussan, C. T. Clingen, and Robert C. Daley. The Multics virtual memory: Concepts and design. Communications of the ACM, 15 (5):308–318, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blasgen:1977:EMM

[BCE77] Michael W. Blasgen, Richard G. Casey, and Kapali P. Eswaran. An encoding method for multifield sorting and indexing. Communications of the ACM, 20(11):874– 878, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also published in/as: IBM, Res.R., RJ 1753, IBM San Jose, Res.Lab, Mar.1976.

Boyce:1975:SQR

[BCKH75] Raymond F. Boyce, Donald D. Chamberlin, W. Frank King III, and Michael M. Hammer. Specifying queries as relational expressions: The SQUARE data sublanguage. Communications of the ACM, 18 (11):621–628, November 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://dblp.uni-trier.de/
db/journals/cacm/BoyceCKH75.

html. Also published in/as: in Douque74 (IFIP TC-2), 1974. Also published in/as: ACM SIGPLAN-SIGIR 1974, Nance (ed), pp. 31– 47. Also published in/as: IBM, San Jose Research Report RJ1291, October 1973.

Barton:1974:ITA

[BCLS74] Ian J. Barton, Susan E. Creasey, Michael F. Lynch, and Michael J. Snell. An information-theoretic approach to text searching in direct access systems. Communications of the ACM, 17(6):345–350, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Balkovich:1974:DMR

[BCPW74] E. Balkovich, W. Chiu, Leon Presser, and R. Wood. Dynamic memory repacking. Communications of the ACM, 17(3):133–138, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blatny:1972:OPT

[BCR72] J. Blatny, S. R. Clark, and T. A.

Rourke. On the optimization of performance of time-sharing systems by simulation. Communications of the ACM, 15(6):411–420, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bruno:1974:SIT

[BCS74] John L. Bruno, Edward G. Coffman, Jr., and Ravi Sethi. Scheduling independent tasks to reduce mean finishing time. Communications of the ACM, 17(7):382–387, July 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Borovits:1977:CUM

[BED77] Israel Borovits and Phillip Ein-Dor. Cost/utilization: A measure of system performance. Communications of the ACM, 20(3):185–191, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:1970:QQM

[Bel70] James R. Bell. The quadratic quotient method: A hash code eliminating secondary clustering. Communications of the ACM, 13(2): 107–109, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:1973:TC

[Bel73] James R. Bell. Threaded code. Communications of the ACM, 16 (6):370–372, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Benjamin:1972:EES

[Ben72a] Arthur J. Benjamin. Extensible editor for a small machine with disk storage. Communications of the ACM, 15(8):742–747, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Benjamin:1972:GPI

[Ben72b] Robert I. Benjamin. A generational perspective of information system development. Communications of the ACM, 15(7):640–643, July 1972.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1975:MBS

[Ben75] Jon Louis Bentley. Multidimensional binary search trees used for associative searching. Communications of the ACM, 18(9):509– 517, September 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://dblp.uni-trier.de/db/
journals/cacm/Bentley75.html.

Bernstein:1970:CWS

[Ber70] A. Bernstein. Comment on the working set model for program behavior. Communications of the ACM, 13(11):698–699, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bernstein:1971:PTP

[Ber71] Arthur J. Bernstein. A programmer training project. Communications of the ACM, 14(7):449–452, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bernstein:1972:CFP

[Ber72] Herbert J. Bernstein. Compiling fixed-point multiplications. Communications of the ACM, 15(8): 772–??, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bitner:1976:EGB

[BER76] James R. Bitner, Gideon Ehrlich, and Edward M. Reingold. Efficient generation of the binary reflected Gray code and its applications. Communications of the ACM, 19 (9):517–521, September 1976. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bertsch:1977:SRP

[Ber77] Eberhard Bertsch. The storage requirement in precedence parsing. Communications of the ACM, 20 (3):192–194, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barton:1971:GRA

[BF71a] David Barton and John P. Fitch. General relativity and the application of algebraic manipulative systems. Communications of the ACM, 14(8):542–547, August 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blumberg:1971:NMN

[BF71b] John W. Blumberg and Clinton R. Foulk. A note on ‘a modification of Nordsieck’s method using an “off-step” point’. Communications of the ACM, 14(12):796, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bracchi:1971:LTG

[BF71c] Giampio Bracchi and Domenico Ferrari. Language for treating geometric patterns in a twodimensional space. Communications of the ACM, 14(1):26–32, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blount:1973:PAC

[BF73] Sumner E. Blount and Louis Fein.

The practical aspect of computer science education — discussion. Communications of the ACM, 16 (1):45–46, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:1978:FSI

[BFS78] Gordon Bell, Samuel H. Fuller, and Daniel P. Siewiorek. Forword to the special issue on computer architecture. Communications of the ACM, 21(1):3–??, 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bernstein:1972:CFM

[BG72a] H. J. Bernstein and H. T. Gladwin. Compiling fixed-point multiplications. Communications of the ACM, 15(8):772, August 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Borodin:1972:CE

[BG72b] Allan Borodin and C. C. Gotlieb.

Computers and employment. Communications of the ACM, 15(7): 695–702, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brawn:1970:SPE

[BGM70] B. S. Brawn, F. G. Gustavson, and E. S. Mankin. Sorting in a paging environment. Communications of the ACM, 13(8):483–494, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Babich:1975:SES

[BGP75] Alan F. Babich, John Grason, and David Lorge Parnas. Significant event simulation. Communications of the ACM, 18(6):323–329, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burkowski:1973:AAC

[BH73] Forbes J. Burkowski and W. D. Hoskins. ACM Algorithm 461: Cubic spline solutions to a class of functional differential equations [D2]. Communications of the ACM, 16(10):635–637, October 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Borgerson:1978:ESU

[BHH78] B. R. Borgerson, M. L. Hanson, and P. A. Hartley. Evolution of the Sperry Univac 1100 Series: a history, analysis, and projection. Communications of the ACM, 21 (1):25–43, January 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Birkhoff:1975:THN

[Bir75] Garrett Birkhoff. Two Hadamard numbers for matrices. Communications of the ACM, 18(1):25–29, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder.

Bird:1977:IPI

[Bir77a] Richard S. Bird. Improving programs by the introduction of recursion. Communications of the ACM, 20(11):856–863, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bird:1977:NRE

[Bir77b] Richard S. Bird. Notes on recursion elimination. Communications of the ACM, 20(6):434–439, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bosarge:1971:NPR

[BJ71] W. E. Bosarge, Jr. and O. G. Johnson. Numerical properties of the Ritz–Trefftz algorithm for optimal control. Communications of the ACM, 14(6):402–406, June 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barrett:1974:SCA

[BJ74] R. C. Barrett and B. W. Jordan, Jr.

Scan conversion algorithms for a cell organized raster display. Communications of the ACM, 17(3):157–163, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Batson:1970:MSS

[BJW70] A. Batson, Shy-Ming M. Ju, and D. C. Wood. Measurements of segment size. Communications of the ACM, 13(3):155–159, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:1970:LQH

[BK70] James R. Bell and Charles H. Kaman. The linear quotient hash code. Communications of the ACM, 13 (11):675–677, November 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Independent discovery of technique of secondary hash functions first proposed by [de 69].

Bron:1973:AAF

[BK73a] Coen Bron and Joep Kerbosch. ACM Algorithm 457: Finding all cliques of an undirected graph. Communications of the ACM, 16 (9):575–577, September 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burkhard:1973:SAB

[BK73b] Walter A. Burkhard and Robert M. Keller. Some approaches to bestmatch file searching. Communications of the ACM, 16(4):230–236, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:1978:ED

[BKHH78] C. Gordon Bell, A. Kotok, Thomas N. Hastings, and R. Hill. The evolution of the DECsystem 10. Communications of the ACM, 21 (1):44–63, January 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bourgeois:1971:EMA

[BL71a] François Bourgeois and Jean-Claude Lassalle. An extension of the Munkres algorithm for the assignment problem to rectangular matrices. Communications of the ACM, 14(12):802–804, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bourgeois:1971:AAA

[BL71b] François Bourgeois and John-Claude C. Lassalle. ACM Algorithm 415: Algorithm for the assignment problem (rectangular matrices) [H]. Communications of the ACM, 14(12):805–806, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bongiovanni:1979:PDB

[BL79] Gian Carlo Bongiovanni and Fabrizio Luccio. Permutation of data blocks in a bubble memory. Communications of the ACM, 22(1): 21–25, January 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Black:1970:OMM

[Bla70] N. A. Black. Optimum merging from mass storage. Communications of the ACM, 13(12):745– 749, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bloom:1970:STT

[Blo70] Burton H. Bloom. Space/time trade-offs in hash coding with allowable errors. Communications of the ACM, 13(7):422–426, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://dblp.uni-trier.
de/db/journals/cacm/Bloom70.
html.

Branquart:1971:CSA

[BLSW71] P. Branquart, J. Lewi, Michel Sintzoff, and P. L. Wodon. The composition of semantics in Algol 68. Communications of the ACM, 14 (11):697–708, November 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Babad:1976:JPM

[BM76] Jair M. Babad and Mario M. Modiano. Joining policies in a multipriority multiclass batch computer system. Communications of the ACM, 19(3):127–136, March 1976. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boyer:1977:FSS

[BM77] Robert S. Boyer and J. Strother Moore. A fast string searching algorithm. Communications of the ACM, 20(10):762–772, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [KMP77, Sun90, BYG92].

Blinn:1976:TRC

[BN76] James F. Blinn and Martin E. Newell. Texture and reflection in computer generated images. Communications of the ACM, 19(10): 542–547, October 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bobrow:1972:RAP

[Bob72] Daniel G. Bobrow. Requirements for advanced programming systems for list processing. Communications of the ACM, 15(7):618–627, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bobrow:1975:NHL

[Bob75] Daniel G. Bobrow. A note on hash linking. Communications of the ACM, 18(7):413–415, July 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boland:1972:AAPa

[Bol72a] W. Robert Boland. ACM Algorithm 436: Product type trapezoidal integration. Communications of the ACM, 15(12):1070, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boland:1972:AAPb

[Bol72b] W. Robert Boland. ACM Algorithm 437: Product type Simpson’s integration [D1]. Communications of the ACM, 15(12):1070–1071, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boland:1972:AAPc

[Bol72c] W. Robert Boland. ACM Algorithm 438: Product type twopoint Gauss–Legendre–Simpson’s integration [D1]. Communications of the ACM, 15(12):1071–??, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boland:1972:AAPd

[Bol72d] W. Robert Boland. ACM Algorithm 439: Product type threepoint Gauss–Legendre–Simpson’s integration [D1]. Communications of the ACM, 15(12):1072–??, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boothroyd:1964:AAG

[Boo64] J. Boothroyd. ACM Algorithm 246: Graycode. Communications of the ACM, 7(12):701, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Mis75, Er85].

Bookstein:1973:HST

[Boo73] Abraham Bookstein. On Harrison’s substring testing technique. Communications of the ACM, 16 (3):180–181, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bouknight:1970:PGT

[Bou70] W. Jack Bouknight. A procedure for generation of three-dimensional half-toned computer graphics presentations. Communications of the ACM, 13(9):527–536, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boulton:1974:RAE

[Bou74] D. M. Boulton. Remark on “Algorithm 434: Exact probabilities for R × C contingency tables”. Communications of the ACM, 17(6):326, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Mar72a, Han75a].

Boulton:1976:REP

[Bou76] D. M. Boulton. Remark on “Algorithm 434: Exact Probabilities for R × C Contingency Tables [G2]”. ACM Transactions on Mathematical Software, 2(1):108, March 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Mar72a].

Boyse:1974:ECP

[Boy74] John W. Boyse. Execution characteristics of programs in a page-on-demand system. Communications of the ACM, 17(4):192–196, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boyse:1979:IDA

[Boy79] John W. Boyse. Interference detection among solids and surfaces. Communications of the ACM, 22 (1):3–9, January 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barnhill:1973:SKC

[BP73] Robert E. Barnhill and David T.

Pilcher. Sard kernels for certain bivariate cubatures. Communications of the ACM, 16(9):567– 570, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barrodale:1974:AAS

[BR74] I. Barrodale and F. D. K. Roberts.

ACM Algorithm 478: Solution of an overdetermined system of equations in the L_{1} norm [F4]. Communications of the ACM, 17(6):319– 320, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [FH75].

Bitner:1975:BPT

[BR75] James R. Bitner and Edward M. Reingold. Backtrack programming techniques. Communications of the ACM, 18(11):651–656, November 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bradley:1970:AAG

[Bra70a] G. H. Bradley. ACM Algorithm 386: Greatest common divisor of n integers and multipliers. Communications of the ACM, 13(7): 447–448, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [RG73].

Bradley:1970:ABG

[Bra70b] Gordon H. Bradley. Algorithm and bound for the greatest common divisor of n integers. Communications of the ACM, 13(7):433–436, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bratley:1970:AAI

[Bra70c] P. Bratley. ACM Algorithm 401: An improved algorithm to produce complex primes. Communications of the ACM, 13(11):693, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bray:1970:CAR

[Bra70d] T. Bray. Certification of Algorithm 22, Ricatti–Bessel Functions of First and Second Kind. Communications of the ACM, 13(7):448, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Braid:1975:SSB

[Bra75] I. C. Braid. The synthesis of solids bounded by many faces. Communications of the ACM, 18(4):209–216, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brainerd:1978:F

[Bra78] Walt Brainerd. Fortran 77. Communications of the ACM, 21(10): 806–820, October 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brenner:1973:AAM

[Bre73a] Norman Brenner. ACM Algorithm 467: Matrix transposition in place [F1]. Communications of the ACM, 16(11):692–694, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Lea79].

Brent:1973:RRT

[Bre73b] Richard P. Brent. Reducing the retrieval time of scatter storage techniques. Communications of the ACM, 16(2):105–109, February 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Modification of open addressing with double hashing to reduce the average number of probes for a successful search.

Brent:1974:AAG

[Bre74] Richard P. Brent. ACM Algorithm 488: A Gaussian pseudo-random number generator [G5]. Communications of the ACM, 17(12):704– 706, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bresenham:1977:LAI

[Bre77] Jack Bresenham. A linear algorithm for incremental digital display of circular arcs. Communications of the ACM, 20(2):100–106, February 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brelaz:1979:NMC

[Bré79] Daniel Brélaz. New methods to color the vertices of a graph. Communications of the ACM, 22(4): 251–256, April 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pee83, KJ85].

BrinchHansen:1970:NMS

[Bri70] P. Brinch Hansen. The nucleus of a multiprogramming system. Communications of the ACM, 13(4): 238–241, 250, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

BrinchHansen:1972:SM

[Bri72] Per Brinch Hansen. Structured multiprogramming. Communications of the ACM, 15(7):574–578, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

BrinchHansen:1978:DPC

[Bri78] Per Brinch Hansen. Distributed processes: A concurrent programming concept. Communications of the ACM, 21(11):934–941, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:1967:AAS

[Bro67] K. M. Brown. ACM Algorithm 316:

Solution of simultaneous nonlinear equations. Communications of the ACM, 10(11):728–729, November 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Rad71].

Broucke:1971:CRN

[Bro71] R. A. Broucke. Construction of rational and negative powers of a formal series. Communications of the ACM, 14(1):32–35, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bron:1972:AAM

[Bro72a] Coenraad Bron. ACM Algorithm 426: Merge sort algorithm [M1]. Communications of the ACM, 15 (5):357–358, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Bro74].

Brown:1972:LLP

[Bro72b] P. J. Brown. Levels of language for portable software. Communications of the ACM, 15(12):1059–1062, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:1972:CSC

[Bro72c] R. J. Brown. Chromatic scheduling and the chromatic number problem. Management Science, 19:451– 463, 1972. CODEN MSCIAM. ISSN 0025-1909 (print), 1526-5501 (electronic). See [KJ85].

Broucke:1973:AAT

[Bro73] R. Broucke. ACM Algorithm 446: Ten subroutines for the manipulation of Chebyshev series [C1]. Communications of the ACM, 16 (4):254–256, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark and certification [PM75].

Bron:1974:RAM

[Bro74] C. Bron. Remark on “Algorithm 426: Merge sort algorithm”. Communications of the ACM, 17(12): 706–??, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Bro72a].

Brown:1976:RAS

[Bro76] Theodore Brown. Remark on “Algorithm 489: The Algorithm SELECT — for Finding the ith Smallest of n Elements [M1]”. ACM Transactions on Mathematical Software, 2(3):301–304, September 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [FR75a].

Brownlee:1977:ABI

[Bro77] J. Nevil Brownlee. An Algolbased implementation of SNOBOL 4 patterns. Communications of the ACM, 20(7):527–529, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Banatre:1979:EDC

[BRT79] Jean-Pierre P. Banâtre, Jean-Paul P. Routeau, and L. Trilling. An event-driven compiling technique. Communications of the ACM, 22(1):34–42, January 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bracchi:1970:ISS

[BS70] G. Bracchi and M. Somalvico.

Interactive software system for computer-aided design. an application to circuit project. Communications of the ACM, 13(9):537– 545, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bernstein:1971:PDS

[BS71] Arthur J. Bernstein and J. C.

Sharp. A policy-driven scheduler for a time-sharing system. Communications of the ACM, 14(2):74–78, February 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barr:1972:CMN

[BS72a] Donald R. Barr and Norman L.

Sezak. A comparison of multivariate normal generators. Communications of the ACM, 15(12):1048– 1049, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bartels:1972:AAS

[BS72b] Richard H. Bartels and G. W. Stewart. ACM Algorithm 432: Solution of the matrix equation ax + xb = c [F4]. Communications of the ACM, 15(9):820–826, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beyer:1973:AAN

[BS73] Terry Beyer and D. F. Swinehart. ACM Algorithm 448: Number of multiply-restricted partitions [A1]. Communications of the ACM, 16(6):379, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baskett:1976:IMC

[BS76] Forest Baskett and Alan Jay Smith. Interference in multiprocessor computer systems with interleaved memory. Communications of the ACM, 19(6):327–334, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baer:1977:CTB

[BS77] Jean-Loup Baer and B. Schwab. A comparison of tree-balancing algorithms. Communications of the ACM, 20(5):322–330, May 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bucci:1979:MDD

[BS79] Giacomo Bucci and Donald N. Streeter. Methodology for the design of distributed information systems. Communications of the ACM, 22(4):233–245, April 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Busenberg:1979:APP

[BT79] Stavros N. Busenberg and Wing C. Tam. An academic program providing realistic training in software engineering. Communications of the ACM, 22(6):341–345, June 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bultheel:1974:RAR

[Bul74] Adhemar Bultheel. Remark on “Algorithm 450: Rosenbrock function minimization”. Communications of the ACM, 17(8):470–??, 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [MM73, KK74].

Burrell:1974:AAE

[Bur74] Keith H. Burrell. ACM Algorithm 484: Evaluation of the modified Bessel functions K_{0}(z) and K_{1}(z) for complex arguments [S17]. Communications of the ACM, 17(9): 524–526, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burton:1976:BSV

[Bur76] Warren Burton. A buddy system variation for disk storage allocation. Communications of the ACM, 19 (7):416–417, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burton:1977:RMS

[Bur77] Warren Burton. Representation of many-sided polygons and polygonal lines for rapid processing. Communications of the ACM, 20(3): 166–171, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Butler:1970:AAG

[But70] E. L. Butler. ACM Algorithm 370: General random number generator. Communications of the ACM, 13 (1):49–52, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Pro72].

Buzen:1973:CAC

[Buz73] Jeffrey P. Buzen. Computational algorithms for closed queueing networks with exponential servers. Communications of the ACM, 16 (9):527–531, September 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bayer:1970:SCC

[BW70] Rudolf Bayer and Christoph Witzgall. Some complete calculi for matrices. Communications of the ACM, 13(4):223–237, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bauer:1972:PKZ

[BW72a] Friedrich L. Bauer and H. Wössner. The “Plankalkül” of Konrad Zuse: a forerunner of today’s programming languages. Communications of the ACM, 15(7):678–685, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Bayer:1972:IRM

[BW72b] Rudolf Bayer and C. Witzgall. Index ranges for matrix calculi. Communications of the ACM, 15(12):

1033–1039, December 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bobrow:1973:MSI

[BW73] Daniel G. Bobrow and Ben Wegbreit. A model and stack implementation of multiple environments. Communications of the ACM, 16 (10):591–603, October 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:1974:SDM

[BW74] J. A. Brown and Bernard Werner.

An on-site data management system application in field archaeology. Communications of the ACM, 17 (11):644–646, November 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burtnyk:1976:IST

[BW76] N. Burtnyk and M. Wein. Interactive skeleton techniques for enhancing motion dynamics in key frame animation. Communications of the ACM, 19(10):564–569, October 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baeza-Yates:1992:NAT

[BYG92] Ricardo Baeza-Yates and Gaston H.

Gonnet. A new approach to text searching. Communications of the ACM, 35(10):74–82, October 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0001-
0782/135243.html. This paper

describes a new linear-time string search algorithm that can handle limited regular-expression pattern matching without backtracking. See also [KMP77], [BM77], [KR81], [Sun90], and [WM92].

Byrne:1973:HMR

[Byr73a] J. G. Byrne. Hu-Tucker minimum redundancy alphabetic coding method. Communications of the ACM, 16(8):490, August 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Byrne:1973:RAH

[Byr73b] J. G. Byrne. Remark on “Algorithm 428: Hu-Tucker minimum redundancy alphabetic coding method”. Communications of the ACM, 16(8):490–??, 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Yoh72].

Cardenas:1973:ESF

[Car73] Alfonso F. Cardenas. Evaluation and selection of file organization — A model and system. Communications of the ACM, 16(9):540– 548, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cardenas:1975:API

[Car75] Alfonso F. Cardenas. Analysis and performance of inverted data base structures. Communications of the ACM, 18(5):253–263, May 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://dblp.uni-trier.de/
                                        db/journals/cacm/Cardenas75.
                                        html.

Carter:1977:CSN

[Car77] J. Lawrence Carter. A case study of a new code generation technique for compilers. Communications of the ACM, 20(12):914–920, December 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Casey:1973:DTS

[Cas73] Richard G. Casey. Design of tree structures for efficient querying. Communications of the ACM, 16 (9):549–556, September 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cavouras:1974:CPD

[Cav74] John C. Cavouras. On the conversion of programs to decision tables: Method and objectives. Communications of the ACM, 17(8):456–462, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chu:1974:EMS

[CB74] Sherwood C. Chu and Mones Berman. An exponential method for the solution of systems of ordinary differential equations. Communications of the ACM, 17(12): 699–702, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coffman:1970:FSU

[CE70] E. G. Coffman, Jr. and J. Eve. File structures using hashing functions. Communications of the ACM, 13 (7):427–432, 436, July 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cermak:1971:AAG

[Čer71] J. Čermák. ACM Algorithm 412: Graph plotter [J6]. Communications of the ACM, 14(7):492–493, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Wat73].

Chang:1970:AAI

[CG70] S. K. Chang and A. Gill. ACM Algorithm 397: An integer programming problem. Communications of the ACM, 13(10):620–621, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [JK72].

Clark:1977:ESL

[CG77] Douglas W. Clark and C. Cordell Green. An empirical study of list structure in Lisp. Communications of the ACM, 20(2):78–87, February 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Corneil:1973:MEN

[CGL73] Derek G. Corneil, C. C. Gotlieb, and Y. M. Lee. Minimal event-node network of project precedence relations. Communications of the ACM, 16(5):296–298, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cale:1979:PPP

[CGM79] E. G. Cale, Lee L. Gremillion, and J. L. McKenney. Price/performance patterns of U.S. computer systems. Communications of the ACM, 22 (4):225–232, April 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chandler:1970:RS

[CH70] J. P. Chandler and W. C. Harrison. R201–SHELLSORT. Communications of the ACM, 13(6):373– 374, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chen:1975:SER

[CH75] Tien Chi Chen and Irving T. Ho.

Storage-efficient representation of decimal data. Communications of the ACM, 18(1):49–52, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder. See comment [Smi75b].

Conte:1976:TCC

[CH76] Samuel D. Conte and Maurice H.

Halstead. Technology of computer center management: a proposed course for graduate professional programs in computer science or in information systems. Communications of the ACM, 19(7):369– 370, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chase:1970:AAC

[Cha70a] P. J. Chase. ACM Algorithm 382:

Combinations of M out of N objects. Communications of the ACM, 13(6):368, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chase:1970:AAP

[Cha70b] P. J. Chase. ACM Algorithm 383: Permutations of a set with repetitions. Communications of the ACM, 13(6):368–369, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chambers:1971:AAP

[Cha71a] J. M. Chambers. ACM Algorithm 410: Partial sorting [M1]. Communications of the ACM, 14(5):357– 358, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chang:1971:RBP

[Cha71b] Shi-Kuo Chang. The reconstruction of binary patterns from their projections. Communications of the ACM, 14(1):21–25, January 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chase:1972:IGA

[Cha72] Stephen M. Chase. An implemented graph algorithm for winning Shannon Switching Games. Communications of the ACM, 15(4):253–256, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chang:1973:AAB

[Cha73] Shi-Kuo Chang. ACM Algorithm 445: Binary pattern reconstruction from projections [Z]. Communications of the ACM, 16(3):185–186, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Lau73].

Cheney:1970:NLC

[Che70] C. J. Cheney. A nonrecursive list compacting algorithm. Communications of the ACM, 13(11): 677–678, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cheng:1978:GBV

[Che78] R. C. H. Cheng. Generating beta variates with nonintegral shape parameters. Communications of the ACM, 21(4):317–322, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Canaday:1974:BEC

[CHI^{+}74] R. H. Canaday, R. D. Harrison, Evan L. Ivie, J. L. Ryder, and L. A. Wehr. A back-end computer for data base management. Communications of the ACM, 17(10): 575–582, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chin:1978:ADN

[Chi78] Francis Y. Chin. An O(n) algorithm for determining a nearoptimal computation order of matrix chain products. Communications of the ACM, 21(7):544–549, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Courtois:1971:CCR

[CHP71] P. J. Courtois, F. Heymans, and David Lorge Parnas. Concurrent control with “readers” and “writers”. Communications of the ACM, 14(10):667–668, October 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Christman:1972:MSA

[Chr72] Ronald D. Christman. MUX, a simple approach to on-line computing. Communications of the ACM, 15 (5):319–329, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chrysler:1978:SBD

[Chr78] Earl Chrysler. Some basic determinants of computer programming productivity. Communications of the ACM, 21(6):472–483, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cardenas:1970:PLP

[CK70] A. F. Cardenas and W. J. Karplus. PDEL — a language for partial differential equations. Communications of the ACM, 13(3):184–191, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cocke:1977:ARO

[CK77] John Cocke and Ken Kennedy. An algorithm for reduction of operator strength. Communications of the ACM, 20(11):850–856, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chan:1979:CSD

[CL79] Tony F. Chan and John Gregg Lewis. Computing standard deviations: Accuracy. Communications of the ACM, 22(9):526– 531, September 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Claudson:1975:DSR

[Cla75] R. Mark Claudson. Digital simulation of river plankton population dynamics. Communications of the ACM, 18(9):517–523, September 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clark:1976:ELM

[Cla76a] Douglas W. Clark. An efficient list moving algorithm using constant workspace. Communications of the ACM, 19(6):352–354, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clark:1976:DSD

[Cla76b] James H. Clark. Designing surfaces in 3-D. Communications of the ACM, 19(8):454–460, August 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clark:1976:HGM

[Cla76c] James H. Clark. Hierarchical geometric models for visible surface algorithms. Communications of the ACM, 19(10):547–554, October 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clapson:1977:IAT

[Cla77] Philip Clapson. Improving the access time for random access files. Communications of the ACM, 20 (3):127–135, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clark:1978:FAC

[Cla78] Douglas W. Clark. A fast algorithm for copying list structures. Communications of the ACM, 21(5):351– 357, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clemm:1969:AAC

[Cle69] D. S. Clemm. ACM Algorithm 352: Characteristic values and associated solutions of Mathieu’s differential equation. Communications of the ACM, 12(7):399–407 (or 399–406??), July 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Fri72].

Cline:1974:AAS

[Cli74a] A. K. Cline. ACM Algorithm 476: Six subprograms for curve fitting using splines under tension [E2]. Communications of the ACM, 17 (4):220–223, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cline:1974:SPV

[Cli74b] A. K. Cline. Scalar- and planarvalued curve fitting using splines under tension. I. Communications of the ACM, 17(4):218–220, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Conway:1972:ISM

[CMM72] Richard W. Conway, William L.

Maxwell, and Howard L. Morgan. On the implementation of security measures in information systems. Communications of the ACM, 15 (4):211–220, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cheriton:1979:TPR

[CMMS79] David R. Cheriton, Michael A.

Malcolm, Lawrence S. Melen, and Gary R. Sager. Thoth, a portable real-time operating system. Communications of the ACM, 22(2): 105–115, February 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chu:1976:APR

[CO76a] Wesley W. Chu and Holger Opderbeck. Analysis of the PFF replacement algorithm via a Semi-Markov model. Communications of the ACM, 19(5):298–304, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Papers from the Fifth ACM Symposium on Operating Systems Principles (Univ. Texas, Austin, Tex., 1975). See corrigendum [CO76b].

Chu:1976:CAP

[CO76b] Wesley W. Chu and Holger Opderbeck. Corrigendum: “Analysis of the PFF replacement algorithm via a semi-Markov model”. Communications of the ACM, 19(9), September 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [CO76a].

Codd:1970:RMD

[Cod70] E. F. Codd. A relational model of data for large shared data banks. Communications of the ACM, 13(6):377–387, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://dblp.uni-trier.de/
                                        db/journals/cacm/Codd70.html.

Reprinted in [Sto88, pp. 5–15].

Coleman:1978:RAS

[Col78] John P. Coleman. Remark on “Algorithm 49: Spherical Neumann Function”. ACM Transactions on Mathematical Software, 4(3):295, September 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Her61].

Colton:1979:IUC

[Col79] Kent W. Colton. The impact and use of computer technology by the police. Communications of the ACM, 22(1):10–20, January 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cotton:1975:RSU

[Cot75] Ira W. Cotton. Remark on stably updating mean and standard deviation of data. Communications of the ACM, 18(8):458, August 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Han75b].

Couger:1973:CRU

[Cou73] J. Daniel Couger. Curriculum recommendations for undergraduate programs in information systems. Communications of the ACM, 16 (12):727–749, December 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Courtois:1975:DIS

[Cou75] P. J. Courtois. Decomposability, instabilities, and saturation in multiprogramming systems. Communications of the ACM, 18(7):371– 377, July 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chambers:1975:CNH

[CP75] Jack A. Chambers and Ray V.

Poore. Computer networks in higher education: Socio-economic-political factors. Communications of the ACM, 18(4):193–199, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Case:1978:AIS

[CP78a] Richard P. Case and Andris Padegs.

Architecture of the IBM System/ 370. Communications of the ACM, 21(1):73–96, January 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chattergy:1978:AAC

[CP78b] Rahul Chattergy and Udo W.

Pooch. Analysis of the availability of computer systems using computer-aided algebra. Communications of the ACM, 21(7):586– 591, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coffman:1972:SSP

[CR72] Edward G. Coffman, Jr. and T. A. Ryan, Jr. A study of storage partitioning using a mathematical model of locality. Communications of the ACM, 15(3):185–190, March 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cheng:1976:SDR

[CR76] Cheng-Wen Cheng and Jonas Rabin. Synthesis of decision rules. Communications of the ACM, 19 (7):404–406, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:1978:ADP

[CR78] Jacques Cohen and Martin S. Roth. Analyses of deterministic parsing algorithms. Communications of the ACM, 21(6):448–458, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chang:1979:IAD

[CR79] Ernest J. H. Chang and Rosemary Roberts. An improved algorithm for decentralized extremafinding in circular configurations of processes. Communications of the ACM, 22(5):281–283, May 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). They present a deterministic distributed algorithm for finding the largest of a set of n uniquely numbered processes in a ring. The algorithm uses O(n log n) messages on the average and O(n^{2}) messages in the worst case, and does not assume that n is known a priori.

Crawford:1973:RBS

[Cra73] C. R. Crawford. Reduction of a band-symmetric generalized eigenvalue problem. Communications of the ACM, 16(1):41–44, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crandall:1974:AAA

[Cra74] Keith C. Crandall. ACM Algorithm 481: Arrow to precedence network transformation [H]. Communications of the ACM, 17(8):467–469, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crespi-Reghizzi:1970:LTG

[CRM70] S. Crespi-Reghizzi and R. Morpurgo. A language for treating graphs. Communications of the ACM, 13(5):319–323, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crespi-Reghizzi:1973:UGI

[CRML73] Stefano Crespi-Reghizzi, Michel A.

Melkanoff, and L. Lichten. The use of grammatical inference for designing programming languages. Communications of the ACM, 16 (2):83–90, February 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crowe:1972:GPA

[Cro72] David Crowe. Generating parsers for affix grammars. Communications of the ACM, 15(8):728–734, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Crow:1977:APC

[Cro77] Franklin C. Crow. The aliasing problem in computer-generated shaded images. Communications of the ACM, 20(11):799–805, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Charnes:1974:GTP

[CRSW74] A. Charnes, W. M. Raike, J. D. Stutz, and A. S. Walters. On generation of test problems for linear programming codes. Communications of the ACM, 17(10):583–586, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Conrow:1970:NPS

[CS70] K. Conrow and R. G. Smith. NEATER2: A PL/I source statement reformatter. Communications of the ACM, 13(11):669– 675, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coleman:1971:AAC

[CT71] M. W. Coleman and M. S. Taylor. ACM Algorithm 403: Circular integer partitioning [A1]. Communications of the ACM, 14(1):48, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cranston:1975:SRS

[CT75] Ben Cranston and Rick Thomas.

Simplified recombination scheme for the Fibonacci buddy system. Communications of the ACM, 18(6): 331–332, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cate:1977:AAS

[CT77] Esko G. Cate and David W. Twigg.

Algorithm 513: Analysis of in-situ transposition [F1]. ACM Transactions on Mathematical Software, 3 (1):104–110, March 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Lea79].

Conway:1973:DID

[CW73] Richard W. Conway and Thomas R.

Wilcox. Design and implementation of a diagnostic compiler for PL/I. Communications of the ACM, 16 (3):169–179, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chang:1978:OHM

[CW78] Shi-Kuo Chang and Yin-Wah Wong. Optimal histogram matching by monotone gray level transformation. Communications of the ACM, 21(10):835–840, October 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:1974:TLE

[CZ74] Jacques Cohen and Carl Zuckerman. Two languages for estimating program efficiency. Communications of the ACM, 17(6):301– 308, June 1974. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Damerau:1970:APC

[Dam70] F. J. Damerau. Automatic parsing for content analysis. Communications of the ACM, 13(6):356– 360, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dathe:1972:CDT

[Dat72] Gert Dathe. Conversion of decision tables by rule mask method without rule mask. Communications of the ACM, 15(10):906–909, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davies:1976:RAR

[Dav76] Alan M. Davies. Remark on “Algorithm 450: Rosenbrock Function Minimization [E4]”. ACM Transactions on Mathematical Software, 2 (3):300–301, September 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [MM73].

Day:1970:FTQ

[Day70] A. C. Day. Full table quadratic searching for scatter storage. Communications of the ACM, 13(8): 481–482, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Deutsch:1976:EIA

[DB76] L. Peter Deutsch and Daniel G. Bobrow. An efficient, incremental, automatic garbage collector. Communications of the ACM, 19(9):522– 526, September 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Danziger:1977:CIA

[DD77a] James N. Danziger and William H.

Dutton. Computers as an innovation in American local governments. Communications of the ACM, 20 (12):945–956, December 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:1977:CPS

[DD77b] Dorothy E. Denning and Peter J.

Denning. Certification of programs for secure information flow. Communications of the ACM, 20 (7):504–513, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

deBalbine:1969:CAR

[de 69] Guy de Balbine. Computational Analysis of the Random Components Induced by a Binary Equivalence Relation. Ph.d. thesis, California Institute of Technology, Pasadena, CA, USA, 1969. 168 pp. First use of second hash function for computing next hash table location after a collision. See also [BK70].

deVries:1970:CLM

[de 70] Ronald C. de Vries. Comment on Lawler’s multilevel Boolean minimization. Communications of the ACM, 13(4):265–266, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DeMorgan:1973:RAE

[De 73] Richard M. De Morgan. Remark on “Algorithm 357: An Efficient Prime Number Generator [A1]”. Communications of the ACM, 16(8):489, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Sin69b].

DeMillo:1978:PAP

[DEL78] Richard A. DeMillo, Stanley C. Eisenstat, and Richard J. Lipton. Preserving average proximity in arrays. Communications of the ACM, 21(3):228–231, March 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dempster:1970:CPA

[Dem70] J. R. H. Dempster. Comment on a paging anomaly. Communications of the ACM, 13(3):193–194, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:1976:LMS

[Den76] Dorothy E. Denning. A lattice model of secure information flow. Communications of the ACM, 19 (5):236–243, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Papers from the Fifth ACM Symposium on Operating Systems Principles (Univ. Texas, Austin, Tex., 1975).

Denning:1979:SPC

[Den79] Dorothy E. Denning. Secure personal computing in an insecure network. Communications of the ACM, 22(8):476–482, August 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DeRemer:1971:SLK

[DeR71] Franklin L. DeRemer. Simple LR(k)

grammars. Communications of the ACM, 14(7):453–460, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Deutsch:1972:TAR

[Deu72] Edward S. Deutsch. Thinning algorithms on rectangular, hexagonal, and triangular arrays. Communications of the ACM, 15(9):827– 837, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Devillers:1977:GID

[Dev77] Raymond R. Devillers. Game interpretation of the deadlock avoidance problem. Communications of the ACM, 20(10):741–745, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dewar:1975:ITC

[Dew75] Robert B. K. Dewar. Indirect threaded code. Communications of the ACM, 18(6):330–331, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:1973:NSO

[DG73] Peter J. Denning and G. Scott Graham. A note on subexpression ordering in the evaluation of arithmetic expressions. Communications of the ACM, 16(11):700– 702, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Erratum, ibid. 17, 1974, 455.

Duda:1972:UHT

[DH72] Richard O. Duda and Peter E. Hart. Use of the Hough transformation to detect lines and curves in pictures. Communications of the ACM, 15 (1):11–15, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Driessen:1973:RAL

[DH73] H. B. Driessen and E. W. LeM. Hunt. Remark on “Algorithm 429: Localization of the Roots of a Polynomial”. Communications of the ACM, 16(9):579, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Squ72, Wil73a].

Diffie:1976:CPD

[DH76] Whitfield Diffie and Martin E. Hellman. A critique of the proposed Data Encryption Standard. Communications of the ACM, 19(3): 164–165, March 1976. URL https:

//dl.acm.org/doi/pdf/10.1145/
360018.360031.

Dial:1970:AAD

[Dia70] R. B. Dial. ACM Algorithm 394: Decision table translation. Communications of the ACM, 13(9):571– 572, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Mar72b].

Dijkstra:1968:GSC

[Dij68] Edsger Wybe Dijkstra. Go to statement considered harmful. Communications of the ACM, 11(3): 147–148, March 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This paper inspired scores of others, published mainly in SIGPLAN Notices up to the mid-1980s. The best-known is [Knu74b].

Dijkstra:1972:HP

[Dij72] Edsger W. Dijkstra. The humble programmer. Communications of the ACM, 15(10):859–866, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1974:SSS

[Dij74] Edsger W. Dijkstra. Self-stabilizing systems in spite of distributed control. Communications of the ACM, 17(11):643–644, November 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1975:GCN

[Dij75] Edsger W. Dijkstra. Guarded commands, nondeterminacy and formal derivation of programs. Communications of the ACM, 18(8):453–457, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dittmer:1976:IEP

[Dit76] Ingo Dittmer. Implementation eines Einschrittcompilers für die Programmiersprache PASCAL auf der Rechenanlage IBM/360 der Universität Münster. (English title: Implementation of a one-step compiler for the programming language PAS-CAL on the IBM/360 of the University of Muenster). Diplomearbeit, Universität Münster, Münster, Germany, ?? 1976. ?? pp. Diplomearbeit Münster 1976 und doert angegebene Literatur (English: Muenster diploma work 1976 and the literature cited therein). The hashing method was rediscovered fourteen years later by Pearson [Pea90], and then commented on by several authors [Dit91, Sav91, Lit91, Pea91].

Dittmer:1991:NFH

[Dit91] I. Dittmer. Note on fast hashing of variable length text strings. Communications of the ACM, 34 (11):118, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Points out that Pearson’s hashing algorithm [Pea90] was discovered fourteen years earlier by this author [Dit76]. See also comments in [Sav91, Lit91, Pea91].

Dutton:1978:MUC

[DK78] William H. Dutton and Kenneth L. Kraemer. Management utilization of computers in American local governments. Communications of the ACM, 21(3):206–218, March 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Deliyanni:1979:LSN

[DK79] Amaryllis Deliyanni and Robert A. Kowalski. Logic and semantic networks. Communications of the ACM, 22(3):184–192, March 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Duong-Kien:1976:IMA

[DKHM76] C. Duong-Kien, Hans-Jürgen Hoffmann, and D. Muth. An improvement to Martin’s algorithm for computation of linear precedence functions. Communications of the ACM, 19(10):576–577, October 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Driscoll:1978:STA

[DL78] James R. Driscoll and Y. Edmund Lien. A selective traversal algorithm for binary search trees. Communications of the ACM, 21(6):445– 447, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1978:FGC

[DLM^{+}78] Edsger W. Dijkstra, Leslie Lamport, Alain J. Martin, C. S. Scholten, and E. F. M. Steffens. On-the-fly garbage collection: An exercise in cooperation. Communications of the ACM, 21(11):966– 975, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also E. W. Dijkstra Note EWD496, June 1975.

DeMillo:1979:SPP

[DLP79] Richard A. De Millo, Richard J.

Lipton, and Alan J. Perlis. Social processes and proofs of theorems and programs. Communications of the ACM, 22(5):271–280, May 1979.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dalal:1978:RPF

[DM78] Yogen K. Dalal and Robert M. Metcalfe. Reverse path forwarding of broadcast packets. Communications of the ACM, 21(12): 1040–1048, December 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dershowitz:1979:PTM

[DM79] Nachum Dershowitz and Zohar Manna. Proving termination with multiset orderings. Communications of the ACM, 22(8):465–476, August 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Donnelly:1973:AAB

[Don73] Thomas G. Donnelly. ACM Algorithm 462: Bivariate normal distribution [S15]. Communications of the ACM, 16(10):638, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Donovan:1976:TPS

[Don76] John J. Donovan. Tools and philosophy for software education. Communications of the ACM, 19(8): 430–436, August 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Doran:1975:RAV

[Dor75] C. J. Doran. Remark on “Algorithm 475: Visible surface plotting program”. Communications of the ACM, 18(5):277–??, 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wri74, Fre75, Mas75].

DeSalvio:1970:CCI

[DPR70] A. J. DeSalvio, J. G. Purdy, and J. Rau. Creation and control of internal data bases under a Fortran programming environment. Communications of the ACM, 13(4): 211–215, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:1972:PAS

[DR72] Peter J. Denning and Brian Randell. Papers from the 3rd ACM symposium on operating systems principles. Communications of the ACM, 15(3):133–134, March 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:1972:PWS

[DS72] Peter J. Denning and Stuart C.

Schwartz. Properties of the working set model. Communications of the ACM, 15(3):191–198, March 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [DS73].

Denning:1973:CWM

[DS73] Peter J. Denning and Stuart C.

Schwartz. Corrigendum: “Properties of the working-set model”. Communications of the ACM, 15 (2):191–198, February 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [DS72].

Denning:1978:GWS

[DS78] Peter J. Denning and Donald R. Slutz. Generalized working sets for segment reference strings. Communications of the ACM, 21(9):750– 759, September 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dunham:1970:AAA

[Dun70] K. B. Dunham. ACM Algorithm 372: An algorithm to produce complex primes CSIEVE. Communications of the ACM, 13(1):52–54, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dunham:1972:MNA

[Dun72] Charles B. Dunham. Minimax nonlinear approximation by approximation on subsets. Communications of the ACM, 15(5):351, May 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Duta:1976:RVS

[Dut76] Lucian D. Duta. Remark on “Algorithm 475: Visible Surface Plotting Program [J6]”. ACM Transactions on Mathematical Software, 2 (1):109–110, March 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Wri74].

Dell:1971:AAR

[DWT71] Alice M. Dell, Roman L. Weil, and Gerald L. Thompson. ACM Algorithm 405: Roots of matrix pencils: The generalized eigenvalue problem [F2]. Communications of the ACM, 14(2):113–117, February 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Hei72].

Dwyer:1972:TSA

[Dwy72] T. A. Dwyer. Teacher/student authored CAI using the NEWBASIC system. Communications of the ACM, 15(1):21–28, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Earley:1970:ECF

[Ear70] Jay Earley. An efficient context-free parsing algorithm. Communications of the ACM, 13(2):94–102, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Earley:1971:TUD

[Ear71] Jay Earley. Toward an understanding of data structures. Communications of the ACM, 14(10):617–627, October 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eastman:1970:RSP

[Eas70] C. M. Eastman. Representations for space planning. Communications of the ACM, 13(4):242–250, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eastman:1972:PRS

[Eas72] Charles M. Eastman. Preliminary report on a system for general space planning. Communications of the ACM, 15(2):76–87, February 1972.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Easton:1977:TFW

[EB77] Malcolm C. Easton and B. T. Bennett. Transient-free working-set statistics. Communications of the ACM, 20(2):93–99, February 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Earley:1972:MIC

[EC72] Jay Earley and Paul Caizergues. Method for incrementally compiling languages with nested statement structure. Communications of the ACM, 15(12):1040–1044, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Easton:1978:CSV

[EF78] Malcolm C. Easton and Ronald Fagin. Cold-start vs. warm-start miss ratios. Communications of the ACM, 21(10):866–872, October 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eswaran:1976:NCP

[EGLT76] Kapali P. Eswaran, Jim N. Gray, Raymond A. Lorie, and Irving L. Traiger. The notions of consistency and predicate locks in a database system. Communications of the ACM, 19(11):624– 633, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://dblp.uni-trier.de/db/
journals/cacm/EswarranGLT76.

html. Also published in/as: IBM Research Report RJ1487, San Jose, CA, December, 1974.

Evans:1975:NLF

[EH75] D. J. Evans and Michael Hatzopoulos. A note on the LU factorization of a symmetric matrix. Communications of the ACM, 18(5):278– 279, May 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ehrlich:1970:CMI

[Ehr70a] L. W. Ehrlich. Complex matrix inversion versus real. Communications of the ACM, 13(9):561– 562, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ehrman:1970:CAC

[Ehr70b] J. R. Ehrman. Correction to ’logical’ arithmetic on computers with two’s complement binary arithmetic. Communications of the ACM, 13(11):697–698, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ehrlich:1973:AAF

[Ehr73a] Gideon Ehrlich. ACM Algorithm 466: Four combinatorial algorithms [G6]. Communications of the ACM, 16(11):690–691, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ehrlich:1973:SBE

[Ehr73b] Louis W. Ehrlich. Solving the biharmonic equation in a square: a direct versus a semidirect method. Communications of the ACM, 16(11):

711–714, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ehrlich:1974:AAG

[Ehr74] Gideon Ehrlich. ACM Algorithm 477: Generator of set-partitions to exactly R subsets [G7]. Communications of the ACM, 17(4):224–225, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Einarsson:1972:AAC

[Ein72a] Bo Einarsson. ACM Algorithm 418: Calculation of Fourier integrals [D1]. Communications of the ACM, 15(1):47–48, January 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Ein72b, Pie74c].

Einarsson:1972:RAC

[Ein72b] Bo Einarsson. Remark on “Algorithm 418: Calculation of Fourier integrals”. Communications of the ACM, 15(6):469, June 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ein72a, Pie74c].

Einarsson:1974:STE

[Ein74] Bo Einarsson. Remark on “Algorithm 443: Solution of the transcendental equation w exp(w) = x”. Communications of the ACM, 17(4):225, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [FSC73].

Evans:1974:UAS

[EKW74] Arthur Evans, Jr., William Kantrowitz, and Edwin Weiss. A user authentication scheme not requiring secrecy in the computer. Communications of the ACM, 17(8):437–442, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lozy:1976:RCS

[eL76] Mohamed el Lozy. Remark on “Algorithm 299: Chi-Squared Integral [S15]”. ACM Transactions on Mathematical Software, 2(4): 393–395, December 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HP67, HP85].

Edgar:1979:FMN

[EL79a] Albert D. Edgar and Samuel C.

Lee. FOCUS microcomputer number system. Communications of the ACM, 22(3):166–177, March 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lozy:1979:RSD

[eL79b] Mohamed el Lozy. Remark on “Algorithm 395: Student’s t-Distribution” and remark on “Algorithm 396: Student’s Quantiles [S14]”. ACM Transactions on Mathematical Software, 5(2):238– 239, June 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Hil70a, Hil70b, Hil81a, HP85].

Elder:1970:FVI

[Eld70] H. A. Elder. On the feasibility of voice input to an on-line computer processing system. Communications of the ACM, 13(6):339– 346, June 1970. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Ellis:1974:RAHa

[Ell74a] T. M. R. Ellis. Remark on “Algorithm 420: Hidden-line plotting program”. Communications of the ACM, 17(6):324–325, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Wil72, MC73, Wil73c, Ell74b, Gai74b, MM73].

Ellis:1974:RAHb

[Ell74b] T. M. R. Ellis. Remark on “Algorithm 420: Hidden-line plotting program”. Communications of the ACM, 17(12):706–??, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [Wil72, MC73, Wil73c, Ell74a, Gai74b, Gau64a].

Eisenberg:1972:FCD

[EM72] Murray A. Eisenberg and Michael R. McGuire. Further comments on Dijkstra’s concurrent programming control problem. Communications of the ACM, 15(11):999, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Even:1978:EEC

[ER78] Shimon Even and Michael Rodeh. Economical encoding of commas between strings. Communications of the ACM, 21(4):315–317, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Er:1985:RG

[Er85] M. C. Er. Remark on “Algorithm 246: Graycode [Z]”. ACM Transactions on Mathematical Software, 11 (4):441–443, December 1985. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Boo64, Mis75].

Ershov:1972:CAH

[Ers72a] A. P. Ershov. Corrigendum: “Aesthetics and the Human Factor in Programming”. Communications of the ACM, 15(10):913, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ers72b].

Ershov:1972:AHF

[Ers72b] Andrei P. Ershov. Aesthetics and the human factor in programming. Communications of the ACM, 15 (7):501–505, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Ers72a].

Earley:1970:FTI

[ES70] J. Earley and H. Sturgis. A formalism for translator interactions. Communications of the ACM, 13 (10):607–617, October 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eidson:1974:AAC

[ES74] Harold D. Eidson and Larry L.

Schumaker. ACM Algorithm 485: Computation of g-splines via a factorization method [E2]. Communications of the ACM, 17(9):526– 530, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Estell:1973:CPA

[Est73] Robert G. Estell. A comment on the practical aspects of computer science education. Communications of the ACM, 16(5):314–315, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erisman:1975:CCE

[ET75] A. M. Erisman and W. F. Tinney. On computing certain elements of the inverse of a sparse matrix. Communications of the ACM, 18 (3):177–179, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Evans:1974:MEL

[Eva74] R. V. Evans. Multiple exits from a loop using neither GO TO nor labels. Communications of the ACM, 17(11):650, November 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Even:1974:PTS

[Eve74] Shimon Even. Parallelism in tapesorting. Communications of the ACM, 17(4):202–204, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fabry:1973:DVO

[Fab73] Robert S. Fabry. Dynamic verification of operating system decisions. Communications of the ACM, 16 (11):659–668, November 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fabry:1974:CBA

[Fab74] Robert S. Fabry. Capability-based addressing. Communications of the ACM, 17(7):403–412, July 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fagin:1976:CCE

[Fag76a] Ronald Fagin. Corrigendum: “A Counterintuitive Example of Computer Paging”. Communications of the ACM, 19(4):187, April 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Fag76b].

Fagin:1976:CEC

[Fag76b] Ronald Fagin. A counterintuitive example of computer paging. Communications of the ACM, 19(2): 96–97, February 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Fag76a].

Fairley:1979:ECD

[Fai79] Richard E. Fairley. Employment characteristics of doctoral level computer scientists. Communications of the ACM, 22(2):77–78, February 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fajman:1973:WIT

[FB73] Roger Fajman and John Borgelt.

WYLBUR, an interactive text editing and remote job entry system. Communications of the ACM, 16 (5):314–322, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feinstein:1975:RMT

[Fei75] Robert Feinstein. Remark on “Algorithm 483: Masked Three-Dimensional Plot Program with Rotations [J6]”. ACM Transactions on Mathematical Software, 1(9): 285, September 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Wat74].

Feldman:1979:HLP

[Fel79] Jerome A. Feldman. High level programming for distributed computing. Communications of the ACM, 22(6):353–368, June 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fenichel:1971:CCL

[Fen71a] Robert R. Fenichel. Comment on Cheney’s list-compaction algorithm. Communications of the ACM, 14(9):603–604, September 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fenichel:1971:LTS

[Fen71b] Robert R. Fenichel. List tracing in systems allowing multiple cell types. Communications of the ACM, 14 (8):522–526, August 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fenichel:1971:ILV

[Fen71c] Robert R. Fenichel. On implementation of label variables. Communications of the ACM, 14(5):349– 350, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fencl:1973:AAR

[Fen73] Zdeněk Fencl. ACM Algorithm 456: Routing problem. Communications of the ACM, 16(9):572– 574, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Tes74].

Ferguson:1971:BAM

[Fer71] David E. Ferguson. Buffer allocation in merge-sorting. Communications of the ACM, 14(7):476– 478, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ferrari:1974:ILC

[Fer74] Domenico Ferrari. Improving locality by critical working sets. Communications of the ACM, 17(11): 614–620, November 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feinroth:1973:TUF

[FFG73] Y. Feinroth, E. Franceschini, and M. Goldstein. Telecommunications using a front-end minicomputer. Communications of the ACM, 16 (3):153–160, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Franklin:1974:CPF

[FG74] Mark A. Franklin and R. K. Gupta.

Computation of page fault probability from program transition diagram. Communications of the ACM, 17(4):186–191, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Franklin:1978:AVP

[FGG78] Mark A. Franklin, G. Scott Graham, and R. K. Gupta. Anomalies with variable partition paging algorithms. Communications of the ACM, 21(3):232–236, March 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Friedman:1974:ETR

[FH74] Theodore D. Friedman and Lance J. Hoffman. Execution time requirements for encipherment programs. Communications of the ACM, 17 (8):445–449, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [McC75].

Fritsch:1975:RAS

[FH75] Fred N. Fritsch and Alan C. Hindmarsh. Remark on “Algorithm 478: Solution of an overdetermined system of equations in the L_{1} norm”. Communications of the ACM, 18(5):277–??, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [BR74].

Fiala:1973:AAS

[Fia73] Frantisek Fiala. ACM Algorithm 449: Solution of linear programming problems in 0-1 variables [H1]. Communications of the ACM, 16 (7):445–447, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fielding:1970:AAF

[Fie70] K. Fielding. ACM Algorithm 387: Function minimization and linear search. Communications of the ACM, 13(8):509–510, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Finerman:1975:PCF

[Fin75] Aaron Finerman. Professionalism in the computing field. Communications of the ACM, 18(1):4–9, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Findler:1977:SMC

[Fin77] Nicholas V. Findler. Studies in machine cognition using the game of poker. Communications of the ACM, 20(4):230–245, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fisher:1975:CCL

[Fis75] David A. Fisher. Copying cyclic list structures in linear time using bounded workspace. Communications of the ACM, 18(5):251– 252, May 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fishman:1976:SGD

[Fis76] George S. Fishman. Sampling from the Gamma-distribution on a computer. Communications of the ACM, 19(7):407–409, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fishman:1977:ASA

[Fis77] George S. Fishman. Achieving specific accuracy in simulation output analysis. Communications of the ACM, 20(5):310–315, May 1977.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fuchs:1970:EDR

[FJ70] E. A. Fuchs and P. E. Jackson. Estimates of distributions of random variables for certain computer communications traffic models. Communications of the ACM, 13(12): 752–757, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reproduced in Advances in Computer Commun., Chu, W. W., (Ed (1974), 2-7; in Computer Commun., Green, P. E., and Lucky, R. W. (Eds.), (1975), 577-582).

Fuchs:1977:OSR

[FKU77] H. Fuchs, Zvi M. Kedem, and Samuel P. Uselton. Optimal surface reconstruction from planar contours. Communications of the ACM, 20(10):693–702, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:1973:CBS

[FLB73] Jerome A. Feldman, James R. Low, and R. P. Brent. Comment on Brent’s scatter storage algorithm (and author’s reply). Communications of the ACM, 16(11): 703, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fletcher:1966:AAI

[Fle66] W. Fletcher. ACM Algorithm 284: Interchange of two blocks of data. Communications of the ACM, 9(5): 326, May 1966. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ito76].

Floyd:1964:AAT

[Flo64] R. W. Floyd. ACM Algorithm 245:

Treesort 3. Communications of the ACM, 7(12):701, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Lon70].

Florentin:1976:IRC

[Flo76] J. J. Florentin. Information reference coding. Communications of the ACM, 19(1):29–33, January 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Floyd:1979:PP

[Flo79] Robert W. Floyd. The paradigms of programming. Communications of the ACM, 22(8):455–460, August 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Flores:1971:ABS

[FM71] Ivan Flores and George Madpis. Average binary search length for dense ordered lists. Communications of the ACM, 14(9):602–603, September 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [FM72].

Flores:1972:CAB

[FM72] Ivan Flores and George Madpis.

Corrigendum: “Average Binary Search Length for Dense Ordered Lists”. Communications of the ACM, 15(2):113, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [FM71].

Franta:1977:EDS

[FM77] William R. Franta and Kurt Maly. An efficient data structure for the simulation event set. Communications of the ACM, 20(8):596–602, August 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fleisher:1978:NSO

[FM78a] J. M. Fleisher and R. R. Meyer. New sufficient optimality conditions for integer programming and their application. Communications of the ACM, 21(5):411–418, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Franta:1978:CHT

[FM78b] William R. Franta and Kurt Maly. A comparison of heaps and the TL structure for the simulation event set. Communications of the ACM, 21(10):873–875, October 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freeman:1974:GEP

[FMP74] Peter Freeman, Michael A. Malcolm, and William H. Payne. Graduate education: The Ph.D. glut: Response and rebuttal. Communications of the ACM, 17(4):206–207, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pay73].

Foley:1971:AOD

[Fol71] James D. Foley. An approach to the optimum design of computer graphics systems. Communications of the ACM, 14(6):380–390, June 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fosdick:1972:PBM

[Fos72a] Lloyd D. Fosdick. The production of better mathematical software. Communications of the ACM, 15(7):611–617, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Foster:1972:VCA

[Fos72b] Caxton C. Foster. A view of computer architecture. Communications of the ACM, 15(7):557–565, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Foster:1973:GAT

[Fos73] Caxton C. Foster. A generalization of AVL trees. Communications of the ACM, 16(8):513–517, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:1972:CSC

[Fox72] Phyllis Fox. Comparative study of computer programs for integrating differential equations. Communications of the ACM, 15(11): 941–948, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:1975:MKS

[Fox75] B. L. Fox. More on kth shortest paths. Communications of the ACM, 18(5):279, May 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fischer:1979:SRT

[FP79] Patrick C. Fischer and Robert L. Probert. Storage reorganization techniques for matrix computation in a paging environment. Communications of the ACM, 22(7):405– 415, July 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Floyd:1975:AAA

[FR75a] Robert W. Floyd and Ronald L. Rivest. ACM Algorithm 489: The algorithm SELECT — for finding the ith smallest of n elements [M1]. Communications of the ACM, 18(3):173, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Bro76].

Floyd:1975:ETB

[FR75b] Robert W. Floyd and Ronald L. Rivest. Expected time bounds for selection. Communications of the ACM, 18(3):165–172, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frantz:1970:PPA

[Fra70] D. G. Frantz. A PL/ 1 program to assist the comparative linguist. Communications of the ACM, 13 (6):353–356, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fraser:1971:MNP

[Fra71] A. G. Fraser. On the meaning of names in programming systems. Communications of the ACM, 14 (6):409–416, June 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fraser:1972:IBC

[Fra72] A. G. Fraser. On the interface between computers and data communications systems. Communications of the ACM, 15(7):566–573, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frailey:1973:PAM

[Fra73] Dennis J. Frailey. A practical approach to managing resources and avoiding deadlocks. Communications of the ACM, 16(5):323–329, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frane:1976:BBS

[Fra76] James W. Frane. The BMD and BMDP series of statistical computer programs. Communications of the ACM, 19(10):570–576, October 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Francez:1977:AAK

[Fra77] Nissim Francez. Another advantage of keyword notation for parameter communication with subprograms. Communications of the ACM, 20 (8):604–605, August 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freiburghouse:1974:RAU

[Fre74] R. A. Freiburghouse. Register allocation via usage counts. Communications of the ACM, 17(11): 638–642, November 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frederick:1975:RA

[Fre75] Lawrence W. Frederick. Remark on “Algorithm 475”: Visible surface plotting program [J6]. Communications of the ACM, 18(4):202–??, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wri74, Dor75, Mas75].

Freuder:1978:SCE

[Fre78] Eugene C. Freuder. Synthesizing constraint expressions. Communications of the ACM, 21(11): 958–966, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also published as MIT AI MEMO 370, Cambridge, MA, USA, 1976.

Frisch:1972:RAR

[Fri72] Michael J. Frisch. Remark on “Algorithms 352, 385, 392: Remarks on characteristic values and associated solutions of Mathieu’s differential equation, exponential integral, and systems of hyperbolic P.D.E.”. Communications of the ACM, 15(12):1074–??, 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Cle69, Pac70, SM70b].

Ferguson:1973:LSP

[FS73] J. Ferguson and P. A. Staley. Least squares piecewise cubic curve fitting. Communications of the ACM, 16(6):380–382, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freeman:1975:DMA

[FS75] Herbert Freeman and Ruth Shapira.

Determining the minimum-area encasing rectangle for an arbitrary closed curve. Communications of the ACM, 18(7):409–413, July 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frieder:1976:PDA

[FS76] Gideon Frieder and Harry J. Saal.

Process for the determination of addresses in variable length addressing. Communications of the ACM, 19(6):335–338, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:1979:REC

[FS79] Jerome A. Feldman and William R.

Sutherland. Rejuvenating experimental computer science — A report to the National Science Foundation and others. Communications of the ACM, 22(9):497– 502, September 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fritsch:1973:AAS

[FSC73] Fred N. Fritsch, R. E. Shafer, and W. P. Crowley. ACM Algorithm 443: Solution of the transcendental equation we^{w} = x [C5]. Communications of the ACM, 16(2):123–124, February 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [Ein74].

Fullerton:1972:AAM

[Ful72] Wayne Fullerton. ACM Algorithm 435: Modified incomplete gamma function [S14]. Communications of the ACM, 15(11):993–995, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Sch78a].

Fuller:1974:MTP

[Ful74] Samuel H. Fuller. Minimal-total-processing time drum and disk scheduling disciplines. Communications of the ACM, 17(7):376–381, July 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frazer:1972:SNS

[FW72] W. D. Frazer and C. K. Wong. Sorting by natural selection. Communications of the ACM, 15(10): 910–913, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fredman:1978:CCM

[FW78a] Michael L. Fredman and Bruce W. Weide. On the complexity of computing the measure of ⋃[a_{i}, b_{i}]. Communications of the ACM, 21 (7):540–544, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Friedman:1978:NCE

[FW78b] Daniel P. Friedman and David S. Wise. A note on conditional expressions. Communications of the ACM, 21(11):931–933, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fenichel:1970:PTP

[FWY70] R. R. Fenichel, J. Weizenbaum, and J. C. Yochelson. A program to teach programming. Communications of the ACM, 13(3):141–146, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gaines:1972:OSB

[Gai72] R. Stockton Gaines. An operating system based on the concept of a supervisory computer. Communications of the ACM, 15(3):150–156, March 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gaither:1974:HPP

[Gai74a] B. Gaither. Hidden-line plotting program. Communications of the ACM, 17(6):324, June 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gaither:1974:RAH

[Gai74b] Blaine Gaither. Remark on “Algorithm 420: Hidden-line plotting program”. Communications of the ACM, 17(6):324–??, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wil72, MC73, Wil73c, Ell74a, Ell74b].

Galbi:1971:SPS

[Gal71] Elmer W. Galbi. Software and patents: a status report. Communications of the ACM, 14(4):274–278, April 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gallaher:1973:AAM

[Gal73] L. J. Gallaher. ACM Algorithm 440: A multidimensional Monte Carlo quadrature with adaptive stratified sampling [D1]. Communications of the ACM, 16(1):49–50, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Galil:1979:IWC

[Gal79] Zvi Galil. On improving the worse case running time of the Boyer– Moore string matching algorithm. Communications of the ACM, 22 (9):505–508, September 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gannon:1977:EED

[Gan77] John D. Gannon. An experimental evaluation of data type conventions. Communications of the ACM, 20(8):584–595, August 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1964:AAI

[Gau64a] W. Gautschi. ACM Algorithm 222: Incomplete beta function ratios. Communications of the ACM, 7(3):143–144, March 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [KD71, Ell74b].

Gautschi:1964:AAB

[Gau64b] W. Gautschi. ACM Algorithm 236: Bessel functions of the first kind [S17]. Communications of the ACM, 7(8):479–480, August 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sko75a].

Gautschi:1965:AAL

[Gau65] W. Gautschi. ACM Algorithm 259:

Legendre functions for arguments larger than one. Communications of the ACM, 8(8):488–492, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Jan77].

Gautschi:1966:AD

[Gau66] Walter Gautschi. Algorithm 282:

Derivatives of e^{x}/x, cos(x)/x, and sin(x)/x. Communications of the ACM, 9(4):272, April 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [GK70b].

Gautschi:1969:AAC

[Gau69] W. Gautschi. ACM Algorithm 363: Complex error function. Communications of the ACM, 12(11): 635, November 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Köl72].

Gautschi:1973:AAE

[Gau73] Walter Gautschi. ACM Algorithm 471: Exponential integrals [S13]. Communications of the ACM, 16 (12):761–763, December 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gavril:1975:MPP

[Gav75] Fanica Gavril. Merging with parallel processors. Communications of the ACM, 18(10):588–591, October 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grad:1968:AAE

[GB68] J. Grad and M. A. Brebner. ACM Algorithm 343: Eigenvalues and eigenvectors of a real generator matrix. Communications of the ACM, 11(12):820–826, December 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Nie72].

Grapa:1977:STA

[GB77] Enrique Grapa and Geneva G. Belford. Some theorems to aid in solving the file allocation problem. Communications of the ACM, 20 (11):878–882, November 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gelenbe:1973:MWS

[GBK73] Erol Gelenbe, J. C. A. Boekhorst, and J. L. W. Kessels. Minimizing wasted space in partitioned segmentation. Communications of the ACM, 16(6):343–349, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gilbert:1972:IBC

[GC72] Philip Gilbert and W. J. Chandler. Interference between communicating parallel processes. Communications of the ACM, 15(6):427– 437, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Graham:1973:SDE

[GCD73] Robert M. Graham, Gerald J. Clancy, Jr., and David B. De-Vaney. A software design and evaluation system. Communications of the ACM, 16(2):110–116, February 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gelenbe:1978:PRR

[GD78] Erol Gelenbe and D. Derochette.

Performance of rollback recovery systems under intermittent failures. Communications of the ACM, 21 (6):493–499, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gear:1971:AAD

[Gea71a] C. William Gear. ACM Algorithm 407: DIFSUB for solution of ordinary differential equations [D2]. Communications of the ACM, 14 (3):185–190, March 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Nik73].

Gear:1971:AIO

[Gea71b] C. William Gear. The automatic integration of ordinary differential equations. Communications of the ACM, 14(3):176–179, March 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Geddes:1979:RCC

[Ged79] K. O. Geddes. Remark on “Algorithm 424: Clenshaw–Curtis Quadrature [O1]”. ACM Transactions on Mathematical Software, 5 (2):240, June 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Gen72a].

Gelenbe:1973:DPP

[Gel73] Erol Gelenbe. Distribution of a program in primary and fast buffer storage. Communications of the ACM, 16(7):431–434, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Geller:1978:TDA

[Gel78] Matthew M. Geller. Test data as an aid in proving program correctness. Communications of the ACM, 21(5):368–375, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gentleman:1972:AAC

[Gen72a] W. Morven Gentleman. ACM Algorithm 424: Clenshaw–Curtis quadrature [D1]. Communications of the ACM, 15(5):353–355, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Goo73b, Ged79].

Gentleman:1972:ICCa

[Gen72b] W. Morven Gentleman. Implementing Clenshaw–Curtis quadrature. I. methodology and experience. Communications of the ACM, 15(5):337–342, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gentleman:1972:ICCb

[Gen72c] W. Morven Gentleman. Implementing Clenshaw–Curtis quadrature. II. computing the cosine transformation. Communications of the ACM, 15(5):343–346, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gerritsen:1975:PSD

[Ger75] Rob Gerritsen. A preliminary system for the design of DBTG data structures. Communications of the ACM, 18(10):551–557, October 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gries:1977:SID

[GG77] David Gries and Narain H. Gehani.

Some ideas on data types in highlevel languages. Communications of the ACM, 20(6):414–420, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gordon:1971:RPT

[GH71] Richard Gordon and Gabor T. Herman. Reconstruction of pictures from their projections. Communications of the ACM, 14(12):759– 768, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [GH72].

Gordon:1972:CRP

[GH72] Richard Gordon and Gabor T. Herman. Corrigendum: “Reconstruction of Pictures from Their Projections”. Communications of the ACM, 15(2):113, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [GH71].

Gudes:1979:NOE

[GH79] Ehud Gudes and Anthony Hoffman.

A note on: “An optimal evaluation of Boolean expressions in an online query system” [Comm. ACM, 20, 1977, no. 5, 344–347, MR 55 #9627] by M. Z. Hanani. Communications of the ACM, 22(10):550–553, October 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guttag:1978:ADT

[GHM78] John V. Guttag, Ellis Horowitz, and David R. Musser. Abstract data types and software validation. Communications of the ACM, 21(12): 1048–1064, December 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ghosh:1972:FOC

[Gho72] Sakti P. Ghosh. File organization — the consecutive retrieval property. Communications of the ACM, 15 (9):802–808, September 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ghosh:1975:CSR

[Gho75] Sakti P. Ghosh. Consecutive storage of relevant records with redundancy. Communications of the ACM, 18 (8):464–471, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gibbs:1975:AAB

[Gib75a] Norman E. Gibbs. ACM Algorithm 491: Basic cycle generation [H]. Communications of the ACM, 18(5):275–276, May 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gibbs:1975:AAG

[Gib75b] Norman E. Gibbs. ACM Algorithm 492: Generation of all the cycles of a graph from a set of basic cycles [H]. Communications of the ACM, 18(6):310, June 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gifford:1977:HEP

[Gif77] David K. Gifford. Hardware estimation of a process’ primary memory requirements. Communications of the ACM, 20(9):655–663, September 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gimpel:1972:BND

[Gim72] James F. Gimpel. Blocks — A new data type for SNOBOL4. Communications of the ACM, 15(6):438– 447, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gimpel:1973:TDP

[Gim73] James F. Gimpel. A theory of discrete patterns and their implementation in SNOBOL4. Communications of the ACM, 16(2):91–100, February 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gimpel:1974:MSM

[Gim74] James F. Gimpel. The minimization of spatially-multiplexed character sets. Communications of the ACM, 17(6):315–318, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garey:1975:SO

[GJ75] Michael R. Garey and David S. Johnson. On Salazar and Oakford. Communications of the ACM, 18(4):240–241, April 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [SO74, KM75b, KM75a].

Gull:1979:RDS

[GJ79] W. E. Gull and Michael A. Jenkins. Recursive data structures in APL. Communications of the ACM, 22(2):79–96, February 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1970:RCC

[GK70a] Walter Gautschi and Bruce J. Klein. Recursive computation of certain derivatives — A study of error propagation. Communications of the ACM, 13(1):7–9, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1970:RAD

[GK70b] Walter Gautschi and Bruce J. Klein. Remark on Algorithm 282, Derivatives of e^{x}/x, cos(x)/x, and sin(x)/x. Communications of the ACM, 13(1):53–54, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gau66].

Gold:1974:MMR

[GK74] D. E. Gold and David J. Kuck. A model for masking rotational latency by dynamic disk allocation. Communications of the ACM, 17 (5):278–288, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gladwin:1971:NCF

[Gla71] H. Tim Gladwin. A note on compiling fixed point binary multiplications. Communications of the ACM, 14(6):407–408, June 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gentleman:1974:MAR

[GM74] W. Morven Gentleman and Scott B.

Marovich. More on algorithms that reveal properties of floating point arithmetic units. Communications of the ACM, 17(5):276–277, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Mal72].

Gries:1978:LSA

[GM78] David Gries and Jayadev Misra. A linear sieve algorithm for finding prime numbers. Communications of the ACM, 21(12):999–1003, December 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Geschke:1977:EEM

[GMS77] Charles M. Geschke, James H. Morris, Jr., and Edwin H. Satterthwaite. Early experience with Mesa. Communications of the ACM, 20 (8):540–553, August 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldstein:1973:AAC

[Gol73] Richard B. Goldstein. ACM Algorithm 451: Chi-square quantiles [G1]. Communications of the ACM, 16(8):483–485, August 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Kni75].

Goldman:1975:SPC

[Gol75] Neil M. Goldman. Sentence paraphrasing from a conceptual base. Communications of the ACM, 18 (2):96–106, February 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gonnet:1976:HAE

[Gon76] Gaston H. Gonnet. Heaps applied to event driven mechanisms. Communications of the ACM, 19 (7):417–418, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Good:1973:CQ

[Goo73a] A. J. Good. Clenshaw–Curtis quadrature. Communications of the ACM, 16(8):490, August 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Good:1973:RAC

[Goo73b] Albert J. Good. Remark on “Algorithm 424: Clenshaw–Curtis quadrature”. Communications of the ACM, 16(8):490–??, 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gen72a].

Goodenough:1975:EHI

[Goo75] John B. Goodenough. Exception handling: Issues and a proposed notation. Communications of the ACM, 18(12):683–696, December 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gates:1973:STS

[GP73] Geoffrey W. Gates and David A.

Poplawski. A simple technique for structured variable lookup. Communications of the ACM, 16(9): 561–565, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gibbs:1974:TP

[GP74] Norman E. Gibbs and William G.

Poole, Jr. Tridiagonalization by permutations. Communications of the ACM, 17(1):20–24, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Griffin:1970:RME

[GR70] R. Griffin and K. A. Redich. Remark on “Algorithm 347: An Efficient Algorithm for Sorting with Minimal Storage [M1], by R. C. Singleton”. Communications of the ACM, 13(1):54, January 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Sin69a, Pet70].

Ganapathy:1973:ITA

[GR73] S. Ganapathy and V. Rajaraman.

Information theory applied to the conversion of decision tables to computer programs. Communications of the ACM, 16(9):532– 539, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Graham:1975:PSE

[GR75] Susan L. Graham and Steven P. Rhodes. Practical syntactic error recovery. Communications of the ACM, 18(11):639–650, November 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grant:1970:ICG

[Gra70] C. A. Grant. An interactive command generating facility. Communications of the ACM, 13(7):403– 406, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gregory:1972:CFP

[Gre72] James Gregory. A comparison of floating point summation methods. Communications of the ACM, 15 (9):838, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greif:1977:LFP

[Gre77] Irene Greif. A language for formal problem specification. Communications of the ACM, 20(12): 931–935, December 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grimes:1970:SRC

[Gri70] J. E. Grimes. Scheduling to reduce conflict in meetings. Communications of the ACM, 13(6):351– 352, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gries:1977:EPP

[Gri77a] David Gries. An exercise in proving parallel programs correct. Communications of the ACM, 20(12): 921–930, December 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Gri78].

Gries:1977:BPB

[Gri77b] David Gries. On believing programs to be correct. Communications of the ACM, 20(1):49–50, January 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wad76].

Gries:1978:CEP

[Gri78] David Gries. Corrigendum: “An exercise in proving parallel programs correct”. Communications of the ACM, 21(12), December 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gri77a].

Golub:1971:AAC

[GS71] G. H. Golub and L. B. Smith. ACM Algorithm 414: Chebyshev approximation of continuous functions by a Chebyshev system of functions [E2]. Communications of the ACM, 14 (11):737–746, November 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greenspan:1972:FFD

[GS72] D. Greenspan and D. Schultz. Fast finite-difference solution of biharmonic problems. Communications of the ACM, 15(5):347–350, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gunn:1967:AAC

[Gun67] J. H. Gunn. ACM Algorithm 300: Coulomb wave functions. Communications of the ACM, 10(4): 244–245, April 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Vos73].

Gustafson:1971:AARa

[Gus71a] Sven-Åke Gustafson. ACM Algorithm 416: Rapid computation of coefficients of interpolation formulas [E1]. Communications of the ACM, 14(12):806–807, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gustafson:1971:AARb

[Gus71b] Sven-Åke Gustafson. ACM Algorithm 417: Rapid computation of weights of interpolatory quadrature rules [D1]. Communications of the ACM, 14(12):807, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gustafson:1971:RCG

[Gus71c] Sven-Åke A. Gustafson. Rapid computation of general interpolation formulas and mechanical quadrature rules. Communications of the ACM, 14(12):797–801, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gustavson:1978:RAS

[Gus78] Fred G. Gustavson. Remark on “Algorithm 408: A Sparse Matrix Package (Part I) [F4]”. ACM Transactions on Mathematical Software, 4(3):295, September 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [McN71].

Guttag:1977:ADT

[Gut77] John V. Guttag. Abstract data types and the development of data structures. Communications of the ACM, 20(6):396–404, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gilchrist:1974:EFT

[GW74] Bruce Gilchrist and Richard E. Weber. Enumerating full-time programmers. Communications of the ACM, 17(10):592–593, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Graham:1975:FUL

[GW75] Susan L. Graham and Mark N.

Wegman. A fast and usually linear algorithm for global flow analysis (abstract). Communications of the ACM, 18(12):716, December 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [GW76] for full paper.

Graham:1976:FUL

[GW76] Susan L. Graham and Mark Wegman. A fast and usually linear algorithm for global flow analysis. Journal of the ACM, 23(1):172–202, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [GW75].

Ginsberg:1975:AAD

[GZ75] E. S. Ginsberg and Dorothy Zaborowski. ACM Algorithm 490: The dilogarithm function of a real argument [S22]. Communications of the ACM, 18(4):200–202, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Mor76].

Habermann:1972:SCP

[Hab72] A. Nico Habermann. Synchronization of communicating processes. Communications of the ACM, 15 (3):171–176, March 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hahn:1974:NTC

[Hah74] Bruce Hahn. A new technique for compression and storage of data. Communications of the ACM, 17 (8):434–436, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hall:1971:ASR

[Hal71] Andrew D. Hall, Jr. The Altran system for rational function manipulation — a survey. Communications of the ACM, 14(8):517–521, August 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hall:1973:EBA

[Hal73] Patrick A. V. Hall. Equivalence between AND/OR graphs and context-free grammars. Communications of the ACM, 16(7):444– 445, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hamblen:1971:UCH

[Ham71a] John W. Hamblen. Using computers in higher education: past recommendations, status, and needs. Communications of the ACM, 14 (11):709–712, November 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hammer:1971:SSC

[Ham71b] Carl Hammer. Signature simulation and certain cryptographic codes. Communications of the ACM, 14(1):3–14, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hamlet:1973:EMR

[Ham73] Richard G. Hamlet. Efficient multiprogramming resource allocation and accounting. Communications of the ACM, 16(6):337–342, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hamlet:1976:HLB

[Ham76] Richard G. Hamlet. High-level binding with low-level linkers. Communications of the ACM, 19(11): 642–644, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hanson:1972:IEI

[Han72] Richard J. Hanson. Integral equations of immunology. Communications of the ACM, 15(10):883–890, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hanson:1974:STR

[Han74] David R. Hanson. A simple technique for representing strings in Fortran IV. Communications of the ACM, 17(11):646–647, November 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hancock:1975:RAE

[Han75a] T. W. Hancock. Remark on “Algorithm 434: Exact probabilities for R×C contingency tables”. Communications of the ACM, 18(2):117– 119, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Bou74, Mar72a].

Hanson:1975:SUM

[Han75b] Richard J. Hanson. Stably updating mean and standard deviation of data. Communications of the ACM, 18(1):57–58, January 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder. See remark [Cot75].

Hanani:1977:OEB

[Han77] Michael Z. Hanani. An optimal evaluation of Boolean expressions in an online query system. Communications of the ACM, 20(5):344– 347, May 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hansen:1979:PAD

[Han79] Wilfred J. Hansen. Progressive acyclic digraphs — a tool for database integrity. Communications of the ACM, 22(9):513–518, September 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harada:1971:GRP

[Har71a] Kazuaki Harada. Generation of rosary permutations expressed in Hamiltonian circuits. Communications of the ACM, 14(6):373–379, June 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harrison:1971:IST

[Har71b] Malcolm C. Harrison. Implementation of the substring test by hashing. Communications of the ACM, 14(12):777–779, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [TT82].

Harter:1972:OWF

[Har72] Richard Harter. The optimality of Winograd’s formula. Communications of the ACM, 15(5):352, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horn:1978:USI

[HB78] Berthold K. P. Horn and Brett L.

Bachman. Using synthetic images to register real images with surface models. Communications of the ACM, 21(11):914–924, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschberg:1979:CCC

[HCS79] Daniel S. Hirschberg, Ashok K. Chandra, and Dilip V. Sarwate. Computing connected components on parallel computers. Communications of the ACM, 22(8):461–464, August 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1973:AAN

[HD73] G. W. Hill and A. W. Davis. ACM Algorithm 442: Normal deviate [S14]. Communications of the ACM, 16(1):51–52, January 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hearn:1971:ASM

[Hea71] Anthony C. Hearn. Applications of symbol manipulation in theoretical physics. Communications of the ACM, 14(8):511–516, August 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heckel:1978:TID

[Hec78] Paul Heckel. A technique for isolating differences between files. Communications of the ACM, 21 (4):264–268, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heess:1970:CPW

[Hee70] W. F. Heess, Jr. Comments on a paper by Wallace and Mason. Communications of the ACM, 13 (4):264–265, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heiberger:1972:RAR

[Hei72] Richard M. Heiberger. Remark on “Algorithm 405: Roots of matrix pencils”. Communications of the ACM, 15(12):1075–??, 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [DWT71].

Henderson:1972:DSS

[Hen72] Peter B. Henderson. Derived semantics for some programming language constructs. Communications of the ACM, 15(11):967– 973, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herndon:1961:AASa

[Her61] J. R. Herndon. ACM Algorithm 49:

Spherical Neumann function. Communications of the ACM, 4(4):179, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Col78].

Herriot:1972:MGF

[Her72] John G. Herriot. In memory of George E. Forsythe. Communications of the ACM, 15(8):719–720, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Hestenes:1975:PCG

[Hes75] Magnus R. Hestenes. Pseudoinverses and conjugate gradients.

Communications of the ACM, 18 (1):40–43, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder.

Habermann:1976:MHF

[HFC76] A. Nico Habermann, Lawrence Flon, and Lee W. Cooprider. Modularization and hierarchy in a family of operating systems. Communications of the ACM, 19(5):266– 272, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hanson:1978:SPM

[HG78] David R. Hanson and Ralph E. Griswold. The SL5 procedure mechanism. Communications of the ACM, 21(5):392–400, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hsiao:1970:CFS

[HH70a] D. K. Hsiao and F. D. Harary. Corrigenda: “A Formal System for Information Retrieval from Files”. Communications of the ACM, 13(3):266, March 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [HH70b].

Hsiao:1970:FSI

[HH70b] David K. Hsiao and Frank D. Harary. A formal system for information retrieval from files. Communications of the ACM, 13(2): 67–73, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigenda [HH70a].

Hammer:1977:VHL

[HHKW77] Michael Hammer, W. Gerry Howe, Vincent J. Kruskal, and Irving Wladawsky. Very high level programming language for data processing applications. Communications of the ACM, 20(11):832– 840, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1970:AASa

[Hil70a] G. W. Hill. ACM Algorithm 395: Student’s t-distribution. Communications of the ACM, 13(10): 617–619, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [eL79b, Hil81a].

Hill:1970:AASb

[Hil70b] G. W. Hill. ACM Algorithm 396: Student’s t-quantiles. Communications of the ACM, 13(10): 619–620, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Hil81a, Hil81b, eL79b].

Hillstrom:1970:CSA

[Hil70c] K. E. Hillstrom. Comparison of several adaptive Newton–Cotes quadrature routines in evaluating definite integrals with peaked integrands. Communications of the ACM, 13(6):362–365, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1973:AAS

[Hil73a] G. W. Hill. ACM Algorithm 465:

Student’s t frequency [S14]. Communications of the ACM, 16(11):

690, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1973:SPM

[Hil73b] J. Carver Hill. Synchronizing processors with memory-content-generated interrupts. Communications of the ACM, 16(6):350–351, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1981:RSD

[Hil81a] G. W. Hill. Remark on “Algorithm 395: Student’s t-Distribution”. ACM Transactions on Mathematical Software, 7(2):247–249, June 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Hil70a, Hil70b, eL79b].

Hill:1981:RSQ

[Hil81b] G. W. Hill. Remark on “Algorithm 396: Student’s t-Quantiles”. ACM Transactions on Mathematical Software, 7(2):250–251, June 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Hil70b].

Hinds:1975:ALA

[Hin75] James A. Hinds. Algorithm for locating adjacent storage blocks in the buddy system. Communications of the ACM, 18(4):221–222, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschsohn:1970:AHL

[Hir70] I. Hirschsohn. AMESPLOT, a higher level data plotting software system. Communications of the ACM, 13(9):546–555, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschberg:1973:CDM

[Hir73] Daniel S. Hirschberg. A class of dynamic memory allocation algorithms. Communications of the ACM, 16(10):615–618, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschberg:1975:LSA

[Hir75] Daniel S. Hirschberg. A linear space algorithm for computing maximal common subsequences. Communications of the ACM, 18(6):341– 343, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschberg:1976:ITO

[Hir76] Daniel S. Hirschberg. An insertion technique for one-sided heightbalanced trees. Communications of the ACM, 19(8):471–473, August 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschberg:1978:FPS

[Hir78] Daniel S. Hirschberg. Fast parallel sorting algorithms. Communications of the ACM, 21(8):657–661, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hurst:1972:AAG

[HK72] Rex L. Hurst and Robert E. Knop. ACM Algorithm 425: Generation of random correlated normal variables [G5]. Communications of the ACM, 15(5):355–357, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Pag74b].

Hubner:1970:AAR

[HKLS70a] H. Hübner, H. Kremer, K. O. Linn, and W. Schwering. ACM Algorithm 388: Rademacher function. Communications of the ACM, 13 (8):510–511, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hubner:1970:AAB

[HKLS70b] H. Hübner, H. Kremer, K. O. Linn, and W. Schwering. ACM Algorithm 389: Binary ordered Walsh functions. Communications of the ACM, 13(8):511, August 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hubner:1970:AAS

[HKLS70c] H. Hübner, H. Kremer, K. O. Linn, and W. Schwering. ACM Algorithm 390: Sequency ordered Walsh functions. Communications of the ACM, 13(8):511, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harms:1980:RSM

[HKM80] U. Harms, H. Kollakowski, and G. Möller. Remark on “Algorithm 408: A Sparse Matrix Package (Part 1) [F4]”. ACM Transactions on Mathematical Software, 6 (3):456–457, September 1980. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [McN71].

Heckel:1977:TOC

[HL77] Paul G. Heckel and Butler W.

Lampson. A terminal-oriented communication system. Communications of the ACM, 20(7):486–494, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hassitt:1973:IHL

[HLL73] Anthony Hassitt, J. W. Lageschulte, and Leonard E. Lyon. Implementation of a high level language machine. Communications of the ACM, 16(4):199–212, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herman:1978:RMI

[HLL78] Gabor T. Herman, Arnold Lent, and Peter H. Lutz. Relaxation methods for image reconstruction. Communications of the ACM, 21 (2):152–158, February 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hallet:1972:CAS

[HM72] P. Hallet and E. Mund. Certification of “Algorithm 379: Squank”. Communications of the ACM, 15 (12):1073–??, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Lyn70].

Houstis:1972:AAP

[HMR72] E. N. Houstis, W. F. Mitchell, and J. R. Rice. ACM Algorithm 438: Product type twopoint Gauss–Legendre–Simpson’s integration. Communications of the ACM, 15(6):1071, June 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horowitz:1972:CSP

[HMS72] Ellis Horowitz, Howard L. Morgan, and Alan C. Shaw. Computers and society: A proposed course for computer scientists. Communications of the ACM, 15(4):257–261, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hine:1979:CRT

[HMT79] J. H. Hine, Isi Mitrani, and S. Tsur. The control of response times in multi-class systems by memory allocation. Communications of the ACM, 22(7):415–424, July 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoare:1971:PPF

[Hoa71] C. A. R. Hoare. Proof of a program: FIND. Communications of the ACM, 14(1):39–45, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoare:1974:MOS

[Hoa74] C. A. R. Hoare. Monitors: An operating system structuring concept. Communications of the ACM, 17 (10):549–557, October 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Erratum in Communications of the ACM, Vol. 18, No. 2 (February), p. 95, 1975. This paper contains one of the first solutions to the Dining Philosophers problem.

Hoare:1978:CSP

[Hoa78a] C. A. R. Hoare. Communicating sequential processes. Communications of the ACM, 21(8): 666–677, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Hoa78b].

Hoare:1978:CCS

[Hoa78b] C. A. R. Hoare. Corrigendum:

“Communicating Sequential Processes”. Communications of the ACM, 21(11):958, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Hoa78a].

Hodes:1970:PSO

[Hod70] L. Hodes. Programming system for the on-line analysis of biomedical images. Communications of the ACM, 13(5):279–283, 286, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holmgren:1973:RAM

[HOK73] B. Holmgren, D. Obradović, and Å. Kolm. Remark on “Algorithm 333: Minit algorithm for linear programming”. Communications of the ACM, 16(5):310, May 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [SS68, Obr73].

Holt:1971:CPS

[Hol71] Richard C. Holt. Comments on prevention of system deadlocks. Communications of the ACM, 14(1): 36–38, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hollander:1977:RUI

[Hol77] Clifford R. Hollander. Remark on uniform insertion in structured data structures. Communications of the ACM, 20(4):261–262, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Helbig:1972:PRC

[HOR72] Robert E. Helbig, Patrick K. Orr, and Robert R. Roediger. Political redistricting by computer. Communications of the ACM, 15(8): 735–741, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horowitz:1975:SAP

[Hor75] Steven L. Horowitz. Syntactic algorithm for peak detection in waveforms with applications to cardiography. Communications of the ACM, 18(5):281–285, May 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horning:1977:SPA

[Hor77] James J. Horning. Selected papers from the ACM conference on language design for reliable software — introduction. Communications of the ACM, 20(8):539–??, 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoskins:1973:CSS

[Hos73] W. D. Hoskins. Cubic spline solutions to fourth-order boundary value problems. Communications of the ACM, 16(6):382–385, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Howell:1971:AAE

[How71] Jo Ann Howell. ACM Algorithm 406: Exact solution of linear equations using residue arithmetic [F4]. Communications of the ACM, 14 (3):180–184, March 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [How73b].

Howard:1973:MSD

[How73a] John H. Howard, Jr. Mixed solutions for the deadlock problem. Communications of the ACM, 16 (7):427–430, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Howell:1973:RAE

[How73b] Jo Ann Howell. Remark on “Algorithm 406: Exact solution of linear equations using residue arithmetic”. Communications of the ACM, 16(5):311, May 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [How71].

Henschen:1974:TPL

[HOW74] Lawrence J. Henschen, Ross A. Overbeek, and Larry Wos. A theorem-proving language for experimentation. Communications of the ACM, 17(6):308–314, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Howard:1976:PM

[How76] John H. Howard. Proving monitors. Communications of the ACM, 19(5):273–279, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1967:AAC

[HP67] I. D. Hill and M. C. Pike. ACM Algorithm 299: Chi-squared integral. Communications of the ACM, 10(4):243–244, April 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [eL76, HP85].

Halatsis:1978:PHT

[HP78] Constantine Halatsis and George Philokyprou. Pseudochaining in hash tables. Communications of the ACM, 21(7):554–557, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1985:RCS

[HP85] I. D. Hill and M. C. Pike. Remark on “Algorithm 299: Chi-Squared Integral”. ACM Transactions on Mathematical Software, 11 (2):185, June 1985. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HP67, eL76, eL79b].

Harper:1975:S

[HPSS75] L. H. Harper, T. H. Payne, John E.

Savage, and E. Straus. Sorting X + Y. Communications of the ACM, 18(6):347–349, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herriot:1973:AAP

[HR73] John G. Herriot and Christian H.

Reinsch. ACM Algorithm 472: Procedures for natural spline interpolation [E1]. Communications of the ACM, 16(12):763–768, December 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hayes-Roth:1978:IMT

[HRM78] Frederick Hayes-Roth and John P.

McDermott. An interference matching technique for inducing abstractions. Communications of the ACM, 21(5):401–410, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harrison:1976:POS

[HRU76] Michael A. Harrison, Walter L.

Ruzzo, and Jeffrey D. Ullman. Protection in operating systems. Communications of the ACM, 19(8): 461–471, August 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hunt:1977:FAC

[HS77] James W. Hunt and Thomas G.

Szymanski. A fast algorithm for computing longest common subsequences. Communications of the ACM, 20(5):350–353, May 1977.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Held:1978:BTR

[HS78] Gerald Held and Michael Stonebraker. B-trees re-examined. Communications of the ACM, 21(2): 139–143, February 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://dblp.uni-trier.
de/db/journals/cacm/HeldS78.

html. Also published in/as: UCB, Elec. Res. Lab, No.ERL-M528, July 1975.

Hunt:1975:CLK

[HSU75] Harry B. Hunt, III, Thomas G. Szymanski, and Jeffrey D. Ullman. On the complexity of LR(k) testing. Communications of the ACM, 18 (12):707–716, December 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Papers from the Second ACM Symposium on Principles of Programming Languages (Palo Alto, Calif., 1975).

Hunt:1977:OSR

[HSU77] Harry B. Hunt III, Thomas G. Szymanski, and Jeffrey D. Ullman. Operations on sparse relations. Communications of the ACM, 20(3): 171–176, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hopcroft:1973:AAE

[HT73] John E. Hopcroft and Robert Endre Tarjan. ACM Algorithm 447: Efficient algorithms for graph manipulation [H]. Communications of the ACM, 16(6):372–378, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hu:1972:CDC

[Hu72] T. C. Hu. A comment on the double-chained tree. Communications of the ACM, 15(4):276, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huang:1973:NIO

[Hua73] J. C. Huang. A note on information organization and storage. Communications of the ACM, 16(7):406– 410, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hunt:1970:CAA

[Hun70] B. R. Hunt. A comment on axiomatic approaches to programming. Communications of the ACM, 13(7):452, July 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hurwitz:1971:PDV

[Hur71] H. Hurwitz, Jr. On the probability distribution of the values of binary trees. Communications of the ACM, 14(2):99–102, February 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herbst:1972:ELP

[HW72] N. M. Herbst and P. M. Will. An experimental laboratory for pattern recognition and signal processing. Communications of the ACM, 15 (4):231–244, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hunter:1973:AAA

[HW73] D. B. Hunter and Julia M. Williams. ACM Algorithm 455: Analysis of skew representations of the symmetric group. Communications of the ACM, 16(9):571–572, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hanson:1979:MPU

[HW79] Richard J. Hanson and John A. Wisniewski. A mathematical programming updating method using modified Givens transformations and applied to LP problems. Communications of the ACM, 22 (4):245–251, April 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holt:1977:SKS

[HWBC77] Richard C. Holt, David B. Wortman, David T. Bardnard, and James R. Cordy. SP/k: A system for teaching computer programming. Communications of the ACM, 20(5):301–309, May 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harris:1978:AUS

[HZ78] R. A. Harris and J. D. Zund. An algorithm using symbolic techniques for the Bel-Petrov classification of gravitational fields. Communications of the ACM, 21(9):715– 717, September 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ibbett:1978:DMC

[IC78] R. N. Ibbett and P. C. Capon. The development of the MU5 computer system. Communications of the ACM, 21(1):13–24, January 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Irons:1972:CES

[ID72] Edgar T. Irons and Frans M.

Djorup. A CRT editing system. Communications of the ACM, 15 (1):16–20, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Isoda:1971:EBT

[IGK71] Sadahiro Isoda, Eiichi Goto, and Izumi Kimura. An efficient bit table technique for dynamic storage allocation of 2^{n}-word blocks. Communications of the ACM, 14 (9):589–592, September 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Iguchi:1972:SMS

[Igu72] Ken Iguchi. A starting method for solving nonlinear Volterra integral equations of the second kind. Communications of the ACM, 15 (6):460–461, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ichbiah:1970:TGA

[IM70] J. D. Ichbiah and S. P. Morse. A technique for generating almost optimal Floyd-Evans productions for precedence grammars. Communications of the ACM, 13(8):501–508, August 1970. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Ibramsha:1978:DLE

[IR78] M. Ibramsha and V. Rajaraman. Detection of logical errors in decision table programs. Communications of the ACM, 21(12):1016– 1025, December 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Irons:1970:EEL

[Iro70] E. T. Irons. Experience with an extensible language. Communications of the ACM, 13(1):31–40, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ito:1976:RIT

[Ito76] M. R. Ito. Remark on “Algorithm 284: Interchange of Two Blocks of Data [K2]”. ACM Transactions on Mathematical Software, 2 (4):392–393, December 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Fle66].

Ives:1976:PEF

[Ive76] F. M. Ives. Permutation enumeration: Four new permutation algorithms. Communications of the ACM, 19(2):68–72, February 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ivie:1977:PWM

[Ivi77] Evan L. Ivie. Programmer’s workbench — a machine for software development. Communications of the ACM, 20(10):746–753, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jansen:1977:RLF

[Jan77] J. K. M. Jansen. Remark on “Algorithm 259: Legendre Functions for Arguments Larger than One”. ACM Transactions on Mathematical Software, 3(2):204–250, June 1977. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Gau65].

Jordan:1973:SCA

[JB73] B. W. Jordan, Jr. and R. C. Barrett.

Scan conversion algorithm with reduced storage requirements. Communications of the ACM, 16(11): 676–682, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jordan:1974:COR

[JB74] B. W. Jordan, Jr. and R. C. Barrett. A cell organized raster display for line drawings. Communications of the ACM, 17(2):70–77, February 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jefferys:1971:AAM

[Jef71] William H. Jefferys. Automatic algebraic manipulation in celestial mechanics. Communications of the ACM, 14(8):538–541, August 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnston:1975:MRE

[JH75] H. C. Johnston and C. A. R. Hoare.

Matrix reduction — an efficient method (school timetables). Communications of the ACM, 18(3): 141–150, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnson:1972:RAI

[JK72] Stephen C. Johnson and Brian W. Kernighan. Remark on “Algorithm 397: An integer programming problem”. Communications of the ACM, 15(6):469, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [CG70].

Jones:1978:LEE

[JL78] Anita K. Jones and Barbara H. Liskov. A language extension for expressing constraints on data access. Communications of the ACM, 21(5):358–367, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:1970:VSA

[Jon70] B. Jones. A variation on sorting by address calculation. Communications of the ACM, 13(2):105–107, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:1972:CAB

[Jon72] Peter R. Jones. Comment on average binary search length. Communications of the ACM, 15(8):774, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:1986:ECP

[Jon86] Douglas W. Jones. An empirical comparison of priority-queue and event-set implementations. Communications of the ACM, 29 (4):300–311, April 1986. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/pubs/
toc/Abstracts/0001-0782/5686.

html. See [KJ85].

Jazayeri:1975:IEC

[JOR75] Mehdi Jazayeri, William F. Ogden, and William C. Rounds. The intrinsically exponential complexity of the circularity problem for attribute grammars. Communications of the ACM, 18(12):697–706, December 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Papers from the Second ACM Symposium on Principles of Programming Languages (Palo Alto, Calif., 1975).

James:1973:ACP

[JP73] E. B. James and Derek P. Partridge. Adaptive correction of program statements. Communications of the ACM, 16(1):27–37, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jenkins:1972:AAZ

[JT72] Michael A. Jenkins and J. F. Traub.

ACM Algorithm 419: Zeros of a complex polynomial [C2]. Communications of the ACM, 15(2):97–99, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Wit74].

Kailas:1970:AMC

[Kai70] M. V. Kailas. Another method of converting from hexadecimal to decimal. Communications of the ACM, 13(3):193, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kandel:1972:CSV

[Kan72] Abraham Kandel. Computer science — A vicious circle. Communications of the ACM, 15(6):470– 471, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kandel:1973:CSS

[Kan73] Abraham Kandel. Computer science — seminars for undergraduates. Communications of the ACM, 16(7):442, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karp:1975:RGM

[Kar75] Richard Alan Karp. A reply to Gentleman and Marovich. Communications of the ACM, 18(3):174–??, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kase:1963:AAT

[Kas63] R. H. Kase. ACM Algorithm 219: Topological ordering for Pert networks. Communications of the ACM, 6(12):738–739, December 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ten77].

Kimme:1975:FCA

[KBS75] Carolyn Kimme, Dana Ballard, and Jack Sklansky. Finding circles by an array of accumulators. Communications of the ACM, 18(2):120–122, February 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kozdrowicki:1973:CIC

[KC73a] Edward W. Kozdrowicki and Dennis W. Cooper. COKO III: the Cooper-Koz chess program. Communications of the ACM, 16(7): 411–427, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuki:1973:SSA

[KC73b] H. Kuki and W. J. Cody. A statistical study of the accuracy of floating point number systems. Communications of the ACM, 16(4):223–230, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kernighan:1975:STM

[KC75] Brian W. Kernighan and Lorinda L.

Cherry. A system for typesetting mathematics. Communications of the ACM, 18(3):151–157, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kraemer:1979:OES

[KC79a] Kenneth L. Kraemer and Kent W.

Colton. Overview of the EFT symposium. Communications of the ACM, 22(12):641–643, 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kraemer:1979:PVE

[KC79b] Kenneth L. Kraemer and Kent W. Colton. Policy, values, and EFT research: anatomy of a research agenda. Communications of the ACM, 22(12):660–671, December 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kolm:1971:MAL

[KD71] Å. Kolm and T. Dahlstrand. Remark on “Algorithm 333: Minit Algorithm For Linear Programming ([H])”. Communications of the ACM, 14(1):50, 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gau64a].

Kumar:1978:PEH

[KD78] B. Kumar and Edward S. Davidson. Performance evaluation of highly concurrent computers by deterministic simulation. Communications of the ACM, 21(11):904–913, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Keller:1976:FVP

[Kel76] Robert M. Keller. Formal verification of parallel programs. Communications of the ACM, 19(7):371– 384, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kennedy:1972:NOD

[Ken72] Steve Kennedy. A note on optimal doubly-chained trees. Communications of the ACM, 15(11):

997–998, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kernighan:1973:RAM

[Ker73] Brian W. Kernighan. Remark on “Algorithm 422: Minimal Spanning Tree”. Communications of the ACM, 16(7):448, July 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Whi72].

Kessels:1977:AEQ

[Kes77a] Joep L. W. Kessels. Alternative to event queues for synchronization in monitors. Communications of the ACM, 20(7):500–503, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kessels:1977:CFN

[Kes77b] Joep L. W. Kessels. A conceptual framework for a nonprocedural programming language. Communications of the ACM, 20(12): 906–913, December 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Krolak:1971:MMA

[KFM71] Patrick Krolak, Wayne Felts, and George Marble. A man-machine approach toward solving the traveling salesman problem. Communications of the ACM, 14(5):327– 334, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karlton:1976:PHB

[KFSK76] Philip L. Karlton, Samuel H.

Fuller, R. E. Scroggs, and E. B.

Kaehler. Performance of heightbalanced trees. Communications of the ACM, 19(1):23–28, January 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Khalil:1972:EBT

[Kha72] Hatem M. Khalil. The eigenproblem of block tridiagonal matrices. Communications of the ACM, 15 (9):839, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kimble:1970:VGL

[Kim70] G. W. Kimble. A variation of the Goodman-Lance method for the solution of two-point boundary value problems. Communications of the ACM, 13(9):557–558, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kimbleton:1972:RCS

[Kim72] Stephen R. Kimbleton. The role of computer system models in performance evaluation. Communications of the ACM, 15(7):586–590, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

King:1976:SEP

[Kin76] James C. King. Symbolic execution and program testing. Communications of the ACM, 19(7):385– 394, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

King:1973:SCU

[KJ73] Peter J. H. King and Roger G.

Johnson. Some comments on the use of ambiguous decision tables and their conversion to computer programs. Communications of the ACM, 16(5):287–290, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

King:1974:CAV

[KJ74] Peter J. H. King and Roger G.

Johnson. Comments on the algorithms of Verhelst for the conversion of limited-entry decision tables to flowcharts (and author’s reply). Communications of the ACM, 17 (1):43–45, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kubale:1985:GIE

[KJ85] Marek Kubale and Boguslaw Jackowski. A generalized implicit enumeration algorithm for graph coloring. Communications of the ACM, 28(4):412–418, April 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/pubs/
toc/Abstracts/0001-0782/3350.
html.

Klemes:1974:RAR

[KK74] Jiri Klemes and Jaroslav Klemsa.

Remark on “Algorithm 450: Rosenbrock function minimization”. Communications of the ACM, 17 (10):590–591, October 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [MM73, Bul74].

Karr:1978:IUP

[KL78] Michael Karr and David B. Loveman III. Incorporation of units into programming languages. Communications of the ACM, 21(5):385– 391, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Klinger:1971:PWG

[Kli71] Allen Klinger. Pattern width at a given angle. Communications of the ACM, 14(1):15–20, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kling:1978:AWC

[Kli78a] Rob Kling. Automated welfare client-tracking and service integration: the political economy of computing. Communications of the ACM, 21(6):484–493, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kling:1978:VCS

[Kli78b] Rob Kling. Value conflicts and social choice in electronic funds transfer system developments. Communications of the ACM, 21(8): 642–657, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kling:1979:IES

[Kli79a] Rob Kling. Introduction to the EFT symposium. Communications of the ACM, 22(12):639–640, 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Klint:1979:LNM

[Kli79b] Paul Klint. Line numbers made cheap. Communications of the ACM, 22(10):557–559, October 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Korfhage:1975:MSO

[KM75a] Robert R. Korfhage and David W.

Matula. More on the Salazar and Oakford paper. Communications of the ACM, 18(5):303, May 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [SO74, KM75b, GJ75].

Korfhage:1975:SO

[KM75b] Robert R. Korfhage and David W.

Matula. On Salazar and Oakford. Communications of the ACM, 18(4):240, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [SO74, KM75a, GJ75].

Katz:1976:LAP

[KM76a] Shmuel Katz and Zohar Manna.

Logical analysis of programs. Communications of the ACM, 19(4): 188–206, April 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Koppelaar:1976:RNI

[KM76b] Henk Koppelaar and Peter Molenaar. Remark on “Algorithm 486: Numerical Inversion of Laplace Transform [D5]”. ACM Transactions on Mathematical Software, 2 (4):395–396, December 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Vei74].

Knuth:1977:FPM

[KMP77] Donald E. Knuth, J. H. Morris, and V. R. Pratt. Fast pattern matching in strings. SIAM Journal on Computing, 6(2):323–350, June 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See also [BM77] and [Sun90].

Knight:1970:ACP

[Kni70] K. R. Knight. An Algol construction for procedures as parameters of procedures. Communications of the ACM, 13(4):266, April 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knight:1975:CAC

[Kni75] William J. Knight. Certification of “Algorithm 451: Chi-Square quantiles”. Communications of the ACM, 18(2):116–??, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gol73].

Knop:1970:AAR

[Kno70] R. E. Knop. ACM Algorithm 381: Random vectors uniform in solid angle. Communications of the ACM, 13(5):326, May 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sch72a].

Knop:1973:AAR

[Kno73] Robert E. Knop. ACM Algorithm 441: Random deviates from the dipole distribution [G5]. Communications of the ACM, 16(1):51, January 1973. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Knott:1974:NSC

[Kno74] Gary D. Knott. A numbering system for combinations. Communications of the ACM, 17(1):45–46, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kleinrock:1976:SLO

[KNO76a] Leonard Kleinrock, William E.

Naylor, and Holger Opderbeck. A study of line overhead in the ARPANET. Communications of the ACM, 19(1):3–13, January 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knott:1976:NSP

[Kno76b] Gary D. Knott. A numbering system for permutations of combinations. Communications of the ACM, 19(6):355–356, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knott:1977:NSB

[Kno77] Gary D. Knott. A numbering system for binary trees. Communications of the ACM, 20(2):113–115, February 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knuth:1972:ABA

[Knu72a] Donald E. Knuth. Ancient Babylonian algorithms. Communications of the ACM, 15(7):671–677, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery. See errata [Knu76].

Knuth:1972:GFD

[Knu72b] Donald E. Knuth. George Forsythe and the development of Computer Science. Communications of the ACM, 15(8):721–726, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Knuth:1973:ACP

[Knu73] D. E. Knuth. The Art of Computer Programming, Sorting and Searching, volume 3. Addison-Wesley, Reading, MA, USA, 1973. ISBN 0-201-03803-X. xi + 723 pp. LCCN QA76.5 .K74.

Knuth:1974:CPA

[Knu74a] Donald E. Knuth. Computer programming as an art. Communications of the ACM, 17(12):667– 673, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). French translation, with three supplementary paragraphs, in L’Informatique Nouvelle, No. 64 (June 1975), 20– 27. Japanese translation by Makoto Arisawa in bit 7 (1975), 434– 444; reprinted in Kunusu Sensei no Program-Ron (see under Books), 2– 19. English version reprinted with the supplementary paragraphs in ACM Turing Award Lectures: The First Twenty Years (New York: ACM Press, 1987), 33–46; reprinted with corrections in [Knu92, pp. 1– 16]. Russian translation by V. V.

Martyniuk in Lektsii laureatov premii T’ıuringa (Moscow: Mir, 1993), 48–64.

Knuth:1974:SPG

[Knu74b] Donald E. Knuth. Structured programming with go to statements. Computing Surveys, 6(4): 261–301, December 1974. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). Reprinted with revisions in Current Trends in Programming Methodology, Raymond T. Yeh, ed., 1 (Englewood Cliffs, NJ: Prentice-Hall, 1977), 140–194; Classics in Software Engineering, Edward Nash Yourdon, ed. (New York: Yourdon Press, 1979), 259–321. Reprinted with “final” revisions in [Knu92, pp. 17–89]. This paper is a response to [Dij68].

Knuth:1976:EAB

[Knu76] Donald E. Knuth. Errata: “Ancient Babylonian algorithms”. Communications of the ACM, 19(2), February 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Knu72a].

Knuth:1992:LP

[Knu92] Donald E. Knuth. Literate Programming, volume 27 of CSLI Lecture Notes. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xv + 368 pp. LCCN QA76.6.K644. URL

http://csli-www.stanford.edu/
publications/literate.html.

Koffman:1972:IIG

[Kof72] Elliot B. Koffman. Individualizing instruction in a generative CAI tutor. Communications of the ACM, 15(6):472–473, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kolbig:1972:CAC

[Köl72] K. S. Kölbig. Certification of “Algorithm 363: Complex error function”. Communications of the ACM, 15(6):465–466, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gau69].

Koppelaar:1974:CRA

[Kop74] Henk Koppelaar. Certification and remark on “Algorithm 191: Hypergeometric”. Communications of the ACM, 17(10):589–590, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Rel63].

Korman:1970:CO

[Kor70] S. M. Korman. The graphcolouring problem. In N. Christofides, A. Mingozzi, P. Toth, and C. Sandi, editors, Combinatorial Optimization, pages 211–235. Wiley, New York, 1970. See [KJ85].

Kosaraju:1978:IDO

[Kos78] S. Rao Kosaraju. Insertions and deletions in one-sided heightbalanced trees. Communications of the ACM, 21(3):226–227, March 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kowalski:1979:ALC

[Kow79] Robert A. Kowalski. Algorithm = logic + control. Communications of the ACM, 22(7):424–436, July 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karp:1981:ERPa

[KR81] R. M. Karp and M. O. Rabin. Efficient randomized pattern-matching algorithms. Technical report TR-31-81, Harvard University, Cambridge, MA, USA, ?? 1981. ?? pp. An incremental hash function is described for application to the string search problem. See [BYG92].

Krogh:1974:EIV

[Kro74a] Fred T. Krogh. Efficient implementation of a variable projection algorithm for nonlinear least squares problems. Communications of the ACM, 17(3):167–169, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [Kro74b].

Krogh:1974:EEI

[Kro74b] Fred T. Krogh. Errata: “Efficient Implementation of a Variable Projection Algorithm for Nonlinear Least Squares”. Communications of the ACM, 17(10):591, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Kro74a].

Kotiah:1977:OCO

[KS77] Thoddi C. T. Kotiah and David I.

Steinberg. Occurrences of cycling and other phenomena arising in a class of linear programming models. Communications of the ACM, 20(2):107–112, February 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kou:1978:CEC

[KSW78] Lawrence T. Kou, Larry J. Stockmeyer, and C. K. Wong. Covering edges by cliques with regard to keyword conflicts and intersection graphs. Communications of the ACM, 21(2):135–139, February 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kubicek:1973:AAL

[Kub73] Milan Kubicek. ACM Algorithm 470: Linear systems with almost tridiagonal matrix [F4]. Communications of the ACM, 16(12):760– 761, December 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuki:1972:AAC

[Kuk72a] Hirondo Kuki. ACM Algorithm 421: Complex gamma function with error control [S14]. Communications of the ACM, 15(4):271–272, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuki:1972:CGF

[Kuk72b] Hirondo Kuki. Complex gamma function with error control [S14]. Communications of the ACM, 15 (4):262–267, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kou:1975:NSB

[KW75] Lawrence T. Kou and C. K. Wong.

A note on the set basis problem related to the compaction of character sets. Communications of the ACM, 18(11):656–657, November 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knuth:1975:ICU

[KZ75] Donald E. Knuth and Charles T.

Zahn, Jr. Ill-chosen use of “event”. Communications of the ACM, 18 (6):360, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lagerlof:1974:IRR

[Lag74] Rolf O. E. Lagerlöf. Interpolation with rounded ramp functions. Communications of the ACM, 17 (8):476–479, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Laird:1979:COE

[Lai79] Philip D. Laird. Comment on: “An optimal evaluation of Boolean expressions in an online query system” [Comm. ACM, 20, 1977, no. 5, 344–347, MR 55 #9627] by M. Z. Hanani. Communications of the ACM, 22(10):549–550, October 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

LaLonde:1977:RRP

[LaL77] Wilf R. LaLonde. Regular right part grammars and their parsers. Communications of the ACM, 20(10):

731–741, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lampson:1973:NCP

[Lam73] Butler W. Lampson. A note on the confinement problem. Communications of the ACM, 16(10):613–615, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lamport:1974:NSD

[Lam74a] Leslie Lamport. A new solution of Dijkstra’s concurrent programming problem. Communications of the ACM, 17(8):453–455, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lamport:1974:PED

[Lam74b] Leslie Lamport. The parallel execution of DO loops. Communications of the ACM, 17(2):83–93, February 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lamport:1975:MBP

[Lam75] Leslie Lamport. Multiple byte processing with full-word instructions. Communications of the ACM, 18 (8):471–475, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lamport:1977:CRW

[Lam77] Leslie Lamport. Concurrent reading and writing. Communications of the ACM, 20(11):806–811, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lamport:1978:TCO

[Lam78] Leslie Lamport. Time, clocks, and the ordering of events in a distributed system. Communications of the ACM, 21(7):558–565, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lamport:1979:PCC

[Lam79] Leslie Lamport. On the proof of correctness of a calendar program. Communications of the ACM, 22 (10):554–556, October 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lau:1973:RAB

[Lau73] John Lau. Remark on “Algorithm 445: Binary Pattern Reconstruction from Projections [Z]”. Communications of the ACM, 16(3):186, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Cha73].

Lauesen:1975:LSB

[Lau75] Søren Lauesen. A large semaphore based operating system. Communications of the ACM, 18(7):377– 389, July 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lavington:1978:MMA

[Lav78] Simon H. Lavington. The Manchester Mark I and Atlas: a historical perspective. Communications of the ACM, 21(1):4–12, January 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lawrence:1973:RAS

[Law73] E. E. Lawrence. Remark on “Algorithm 408: A sparse matrix package”. Communications of the ACM, 16(9):578, September 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [McN71, Sal73b].

Lawler:1977:CCS

[Law77] Eugene L. Lawler. Comment on computing the k shortest paths in a graph. Communications of the ACM, 20(8):603–604, August 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Laflin:1970:AAS

[LB70a] S. Laflin and M. A. Brebner. ACM Algorithm 380: In-situ transposition of a rectangular matrix. Communications of the ACM, 13(5): 324–326, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Mac72].

Lamport:1970:CBQ

[LB70b] Leslie Lamport and J. R. Bell. Comment on Bell’s quadratic quotient method for hash code searching. Communications of the ACM, 13 (9):573–574, September 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ledgard:1976:CUC

[LC76] Henry F. Ledgard and William C. Cave. Cobol under control. Communications of the ACM, 19(11): 601–608, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:1974:IPS

[LCW74] R. C. T. Lee, C. L. Chang, and Richard J. Waldinger. An improved program-synthesizing algorithm and its correctness. Communications of the ACM, 17(4): 211–217, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Litecky:1976:SEE

[LD76] Charles R. Litecky and Gordon B.

Davis. Study of errors, errorproneness, and error diagnosis in Cobol. Communications of the ACM, 19(1):33–37, January 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leathers:1979:RAS

[Lea79] Burton L. Leathers. Remark on “Algorithm 513: Analysis of In-Situ Transposition [F1]” and remark on “Algorithm 467: Matrix Transposition in Place”. ACM Transactions on Mathematical Software, 5 (4):520, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [CT77, Bre73a].

Ledgard:1972:MTC

[Led72] Henry F. Ledgard. Model for type checking — with an application to Algol 60. Communications of the ACM, 15(11):956–966, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ledgard:1974:PSC

[Led74] Henry F. Ledgard. Production Systems: or can we do better than BNF? Communications of the ACM, 17(2):94–102, February 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:1977:CFP

[Lee77] John A. N. Lee. Considerations for future programming language standards activities. Communications of the ACM, 20(11):788–794, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leitch:1969:CGP

[Lei69] I. M. Leitch. Certification of ACM Algorithm 323 [G6]: Generation of permutations in lexicographic order. Communications of the ACM, 12(9):512, September 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [OS68, Roy73c].

Lemos:1979:ISW

[Lem79] Ronald S. Lemos. An implementation of structured walk-throughs in teaching Cobol programming. Communications of the ACM, 22 (6):335–340, June 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lesk:1972:PPR

[Les72] Arthur M. Lesk. Pictorial pattern recognition and the phase problem of X-ray crystallography. Communications of the ACM, 15(1):3–6, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lesk:1979:DTD

[Les79] Arthur M. Lesk. Detection of threedimensional patterns of atoms in chemical structures. Communications of the ACM, 22(4):219–224, April 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Levialdi:1972:SBP

[Lev72] Stefano Levialdi. On shrinking binary picture patterns. Communications of the ACM, 15(1):7–10, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Levin:1976:PAD

[Lev76] Joshua Z. Levin. A parametric algorithm for drawing pictures of solid objects composed of quadric surfaces. Communications of the ACM, 19(10):555–563, October 1976. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lewart:1973:AAA

[Lew73] C. R. Lewart. ACM Algorithm 463:

Algorithms SCALE1, SCALE2, and SCALE3 for determination of scales on computer generated plots [J6]. Communications of the ACM, 16 (10):639–640, October 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lew:1978:OCE

[Lew78] Art Lew. Optimal conversion of extended-entry decision tables with general cost criteria. Communications of the ACM, 21(4):269–279, April 1978. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Lientz:1976:CEV

[Lie76] Bennet P. Lientz. A comparative evaluation of versions of BASIC. Communications of the ACM, 19 (4):175–181, April 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Linz:1970:AFP

[Lin70] Peter Linz. Accurate floating-point summation. Communications of the ACM, 13(6):361–362, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liniger:1971:SCN

[Lin71] Werner Liniger. A stopping criterion for the Newton–Raphson method in implicit multistep integration algorithms for nonlinear systems of ordinary differential equations. Communications of the ACM, 14(9):600–601, September 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Linz:1972:AAF

[Lin72] Peter Linz. ACM Algorithm 427: Fourier cosine integral [D1]. Communications of the ACM, 15(5): 358–360, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ling:1973:CGA

[Lin73] Robert F. Ling. A computer generated aid for cluster analysis. Communications of the ACM, 16 (6):355–361, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lindstrom:1974:CLS

[Lin74] Gary Lindstrom. Copying list structures using bounded workspace. Communications of the ACM, 17 (4):198–202, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liniger:1975:CBA

[Lin75] Werner Liniger. Connections between accuracy and stability properties of linear multistep formulas. Communications of the ACM, 18 (1):53–56, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder.

Lipton:1975:RMP

[Lip75] Richard J. Lipton. Reduction: A method of proving properties of parallel programs. Communications of the ACM, 18(12):717–721, December 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Papers from the Second ACM Symposium on Principles of Programming Languages (Palo Alto, Calif., 1975).

Lipis:1979:CCU

[Lip79] Allen H. Lipis. Costs of the current U.S. payments system. Communications of the ACM, 22(12): 644–647, December 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liskov:1972:DVO

[Lis72] Barbara H. Liskov. The design of the Venus operating system. Communications of the ACM, 15(3): 144–149, March 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Litsios:1991:NFH

[Lit91] James Litsios. Note on fast hashing of variable length text strings. Communications of the ACM, 34(11): 118–120, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Suggests a simple extension of Pearson’s hashing algorithm [Pea90] that supports non-character data. See also comments in [Dit91, Sav91, Pea91] and early work in [Dit76].

Levitt:1972:CAS

[LK72] Karl N. Levitt and William H. Kautz. Cellular arrays for the solution of graph problems. Communications of the ACM, 15(9):789– 801, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lawrie:1975:GPL

[LLBR75] Duncan H. Lawrie, T. Layman, D. Baer, and J. M. Randal. GLYP-NIR — A programming language for Illiac IV. Communications of the ACM, 18(3):157–164, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lam:1973:AAA

[LM73] C. Lam and J. McKay. ACM Algorithm 469: Arithmetic overa finite field [A1]. Communications of the ACM, 16(11):699, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ledgard:1975:GCS

[LM75] Henry F. Ledgard and Michael Marcotty. A genealogy of control structures. Communications of the ACM, 18(11):629–639, November 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lucas:1974:SCU

[LML74] Henry C. Lucas, Jr., David B.

Montgomery, and Jean Claude Larreche. A study of computer use in a graduate school of business. Communications of the ACM, 17 (4):205–206, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loeckx:1970:ACB

[Loe70] J. Loeckx. Algorithm for the construction of bounded-context parsers. Communications of the ACM, 13(5):297–307, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loeser:1974:SPT

[Loe74] Rudolf Loeser. Some performance tests of “quicksort” and descendants. Communications of the ACM, 17(3):143–152, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

London:1970:CAM

[Lon70] R. L. London. Certification of “Algorithm 245: Treesort 3 [M1]”:

Proof of algorithms — A new kind of certification. Communications of the ACM, 13(6):371–374, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Flo64].

Long:1979:PPE

[Lon79] Robert H. Long. Public protection and education with EFT. Communications of the ACM, 22(12): 648–654, December 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loui:1976:WDT

[Lou76] Michael C. Loui. Weighted derivation trees. Communications of the ACM, 19(9):509–513, September 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lowe:1970:ASC

[Low70] T. C. Lowe. Automatic segmentation of cyclic program structures based on connectivity and processor timing. Communications of the ACM, 13(1):3–6, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Low:1973:PQE

[Low73] David W. Low. Programming by questionnaire: an effective way to use decision tables. Communications of the ACM, 16(5):282–286, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Low:1978:ADS

[Low78] James R. Low. Automatic data structure selection: an example and overview. Communications of the ACM, 21(5):376–385, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Luccio:1978:PT

[LP78] Fabrizio Luccio and Linda Pagli.

Power trees. Communications of the ACM, 21(11):941–947, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lozano-Perez:1979:APC

[LPW79] Tomás Lozano-Pérez and Michael A.

Wesley. An algorithm for planning collision-free paths among polyhedral obstacles. Communications of the ACM, 22(10):560–570, October 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

LaFata:1970:IDA

[LR70] P. LaFata and J. B. Rosen. An interactive display for approximation by linear programming. Communications of the ACM, 13(11): 651–659, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lopez:1977:SCS

[LRT77] A. A. Lopez, Robert Raymond, and Robert Tardiff. A survey of computer science offerings in small liberal arts colleges. Communications of the ACM, 20(12):902–906, December 1977. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Levine:1970:AAS

[LS70] Michael J. Levine and Stanley M. Swanson. ACM Algorithm 377: Symbolic expansion of algebraic expressions. Communications of the ACM, 13(3):191–192, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lyness:1971:AAE

[LS71] J. N. Lyness and G. Sande. ACM Algorithm 413: ENTCAF and ENTCRE: Evaluation of normalized Taylor coefficients of an analytic function (C5). Communications of the ACM, 14(10):669–675, October 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lyche:1974:AAP

[LS74] Tom Lyche and Larry L. Schumaker. ACM Algorithm 480: Procedures for computing smoothing and interpolating natural splines [E1]. Communications of the ACM, 17 (8):463–467, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lampson:1976:ROS

[LS76] Butler W. Lampson and Howard E. Sturgis. Reflections on an operating system design. Communications of the ACM, 19(5):251–265, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lucas:1977:SHC

[LS77] Henry C. Lucas, Jr. and Jimmy A.

Sutton. The stage hypothesis and the S-curve: Some contradictory evidence. Communications of the ACM, 20(4):254–259, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liskov:1977:AMC

[LSAS77] Barbara Liskov, Alan Snyder, Russell R. Atkinson, and Craig Schaffert. Abstraction mechanisms in CLU. Communications of the ACM, 20(8):564–576, August 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also published in/as: In “Readings in Object-Oriented Database Systems” edited by S.Zdonik and D.Maier, Morgan Kaufman, 1990.

Lientz:1978:CAS

[LST78] Bennet P. Lientz, E. Burton Swanson, and G. E. Tompkins. Characteristics of application software maintenance. Communications of the ACM, 21(6):466–471, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lum:1975:COA

[LSWL75] Vincent Y. Lum, Michael E. Senko, C. P. Wang, and H. Ling. A cost oriented algorithm for data set allocation in storage hierarchies. Communications of the ACM, 18(6):318– 322, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lucas:1971:AAC

[LT71] C. W. Lucas, Jr. and C. W. Terrill. ACM Algorithm 404: Complex gamma function [S14]. Communications of the ACM, 14(1):48–49, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liu:1973:AAE

[LT73] C. N. Liu and D. T. Tang. ACM Algorithm 452: Enumerating combinations of m out of n objects [G6]. Communications of the ACM, 16(8):485, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ledgard:1977:TVD

[LT77] Henry F. Ledgard and Robert W. Taylor. Two views of data abstraction — introduction to the special section. Communications of the ACM, 20(6):382–384, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Luccio:1972:WIL

[Luc72] Fabrizio Luccio. Weighted increment linear search for scatter tables. Communications of the ACM, 15(12):1045–1047, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lucas:1974:CRG

[Luc74] Henry C. Lucas, Jr. A CRT report generating system. Communications of the ACM, 17(1):47–48, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lucas:1978:UII

[Luc78] Henry C. Lucas, Jr. Use of an interactive information storage and retrieval system in medical research. Communications of the ACM, 21 (3):197–205, March 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ludwig:1963:AAI

[Lud63] O. G. Ludwig. ACM Algorithm 179:

Incomplete beta ratio. Communications of the ACM, 6(6):314, June 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [PS76, BB74].

Lum:1970:MAR

[Lum70] Vincent Y. Lum. Multi-attribute retrieval with combined indexes. Communications of the ACM, 13 (11):660–665, November 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://dblp.uni-trier.de/
db/journals/cacm/Lum70.html.

Lum:1973:GPA

[Lum73] Vincent Y. Lum. General performance analysis of key-to-address transformation methods using an abstract file concept. Communications of the ACM, 16(10):603–612, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lunde:1977:EES

[Lun77] Åmund Lunde. Empirical evaluation of some features of instruction set processor architectures. Communications of the ACM, 20(3):

143–153, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lum:1972:ARK

[LY72] Vincent Y. Lum and P. S. T. Yuen. Additional results on key-to-address transform techniques: a fundamental performance study on large existing formatted files. Communications of the ACM, 15(11):996– 997, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lum:1971:KAT

[LYD71] Vincent Y. Lum, P. S. T. Yuen, and M. Dodd. Key-to-address transform techniques: A fundamental performance study on large existing formatted files. Communications of the ACM, 14(4):228–239, April 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Survey of several hash functions, with performance results.

Lyness:1970:AAS

[Lyn70] J. N. Lyness. ACM Algorithm 379: SQUANK (Simpson Quadrature Used Adaptively-Noise Killed). Communications of the ACM, 13 (4):260–263, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [HM72].

Lynch:1972:OSP

[Lyn72] William C. Lynch. Operating system performance. Communications of the ACM, 15(7):579–585, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lyon:1974:SDL

[Lyo74] Gordon Lyon. Syntax-directed least-errors analysis for context-free languages: a practical approach. Communications of the ACM, 17 (1):3–14, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lyon:1978:PST

[Lyo78a] Gordon Lyon. Packed scatter tables. Communications of the ACM, 21(10):857–865, October 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lyons:1978:SDE

[Lyo78b] Norman R. Lyons. Systems design education: A gaming approach. Communications of the ACM, 21 (11):889–895, November 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Macleod:1972:STR

[Mac72] I. D. G. Macleod. Certification of “Algorithm 380: In-situ transposition of a rectangular matrix [F1]”. Communications of the ACM, 15(1):49–??, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [LB70a].

MacLennan:1973:FAB

[Mac73] B. J. MacLennan. Fen — an axiomatic basis for program semantics. Communications of the ACM, 16(8):468–474, August 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Magnuski:1975:RAM

[Mag75] H. S. Magnuski. Remark on “Algorithm 479: A minimal spanning tree clustering method”. Communications of the ACM, 18(2):119–??, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pag74a].

Mairson:1977:SNU

[Mai77] Harry G. Mairson. Some new upper bounds on the generation of prime numbers. Communications of the ACM, 20(9):664–669, September 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malcolm:1971:AFP

[Mal71] Michael A. Malcolm. On accurate floating-point summation. Communications of the ACM, 14(11): 731–736, November 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malcolm:1972:ARP

[Mal72] Michael A. Malcolm. Algorithms to reveal properties of floating-point arithmetic. Communications of the ACM, 15(11):949–951, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [GM74].

Maly:1976:CT

[Mal76] Kurt Maly. Compressed tries. Communications of the ACM, 19 (7):409–415, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maly:1978:NVM

[Mal78] Kurt Maly. A note on virtual memory indexes. Communications of the ACM, 21(9):786–787, September 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mamrak:1977:DRT

[Mam77] Sandra A. Mamrak. Dynamic response time prediction for computer networks. Communications of the ACM, 20(7):461–468, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Manthey:1979:MDM

[Man79] Michael J. Manthey. Model for and discussion of multi-interpreter systems. Communications of the ACM, 22(5):299–309, May 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

March:1972:AAE

[Mar72a] David L. March. ACM Algorithm 434: Exact probabilities for R × C contingency tables [G2]. Communications of the ACM, 15(11):991– 992, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Bou74, Han75a, Bou76].

Marshall:1972:RAD

[Mar72b] D. R. T. Marshall. Remark on “Algorithm 394: Decision table translation”. Communications of the ACM, 15(12):1074–??, 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Dia70].

Martin:1972:BMM

[Mar72c] David F. Martin. A Boolean matrix method for the computation of linear precedence functions. Communications of the ACM, 15(6):448– 454, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martelli:1976:AHS

[Mar76] Alberto Martelli. Application of heuristic search methods to edge and contour detection. Communications of the ACM, 19(2):73–83, February 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mashburn:1975:RAV

[Mas75] R. G. Mashburn. Remark on “Algorithm 475: Visible surface plotting program”. Communications of the ACM, 18(5):276–277, 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wri74, Fre75, Dor75].

Matsushita:1972:HLE

[Mat72] Yutaka Matsushita. Hidden lines elimination for a rotating object. Communications of the ACM, 15 (4):245–252, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maurer:1970:NCI

[Mau70] H. A. Maurer. A note on the complement of inherently ambiguous context-free languages. Communications of the ACM, 13(3):194, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mayer:1979:PLB

[May79] Richard E. Mayer. A psychology of learning BASIC. Communications of the ACM, 22(11): 589–593, November 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mifsud:1973:AMP

[MB73] Charles J. Mifsud and Michael J.

Bohlen. Addendum to a multipleprecision division algorithm. Communications of the ACM, 16(10): 628–??, 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Mif70].

Madison:1976:CPL

[MB76a] A. Wayne Madison and Alan P.

Batson. Characteristics of program localities. Communications of the ACM, 19(5):285–294, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Metcalfe:1976:EDP

[MB76b] Robert M. Metcalfe and David R.

Boggs. Ethernet: Distributed packet switching for local computer networks. Communications of the ACM, 19(7):395–404, July 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Macleod:1973:RAH

[MC73] I. D. G. Macleod and A. M.

Collins. Remark on “Algorithm 420: Hidden-line plotting program”. Communications of the ACM, 16(7):448, July 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Wil72, Ell74b, Gai74b, Wil73c, Ell74a].

McCracken:1974:PLI

[McC74] Daniel D. McCracken. A problemlist of issues concerning computers and public policy — A report of the ACM commitee on computers and public policy. Communications of the ACM, 17(9):495–503, 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCarthy:1975:AFP

[McC75] John McCarthy. ACM Forum: Proposed criterion for a cipher to be probable-word-proof. Communications of the ACM, 18(2):131–132, February 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [FH74].

McCreight:1977:PBT

[McC77] Edward M. McCreight. Pagination of B*-trees with variablelength records. Communications of the ACM, 20(9):670–674, September 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGeachie:1973:MTU

[McG73] John S. McGeachie. Multiple terminals under user program control in a time-sharing environment. Communications of the ACM, 16(10): 587–590, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKay:1970:AAU

[McK70a] J. McKay. ACM Algorithm 391: Unitary symmetric polynomials. Communications of the ACM, 13(8):512, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sch72b].

McKay:1970:AAP

[McK70b] J. K. S. McKay. ACM Algorithm 371: Partitions in natural order. Communications of the ACM, 13(1):52, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McMorrow:1973:CMC

[McM73] Clyde H. McMorrow. Concerning music and computer composition in computational linguistics. Communications of the ACM, 16(5):313, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McMaster:1978:AAD

[McM78] Colin L. McMaster. An analysis of algorithms for the Dutch National Flag Problem. Communications of the ACM, 21(10):842–846, October 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [McM79].

McMaster:1979:CAA

[McM79] C. L. McMaster. Corrigendum:

“An Analysis of Algorithms for the Dutch National Flag Problem”. Communications of the ACM, 22(3):207, March 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [McM78].

McNamee:1971:AAS

[McN71] J. M. McNamee. ACM Algorithm 408: A sparse matrix package (part I). Communications of the ACM, 14(4):265–273, April 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [Sal73b, Law73, Sip77, Gus78, HKM80].

Mullery:1970:PAM

[MD70] A. P. Mullery and G. C. Driscoll. A processor allocation method for time-sharing. Communications of the ACM, 13(1):10–14, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCracken:1979:AAE

[MDB79] Daniel D. McCracken, CCP, Peter J. Denning, and David H. Brandin. Addendum: An ACM executive committee position on the crisis in experimental computer science. Communications of the ACM, 22 (9):503–504, September 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Merner:1962:AAC

[Mer62] J. N. Merner. ACM Algorithm 149: Complete elliptic integral. Communications of the ACM, 5(12): 605, December 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Sko78].

Merrill:1973:RCR

[Mer73] R. D. Merrill. Representation of contours and regions for efficient computer search. Communications of the ACM, 16(2):69–82, February 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Merkle:1978:SCI

[Mer78] Ralph C. Merkle. Secure communications over insecure channels. Communications of the ACM, 21 (4):294–299, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mifsud:1970:MDA

[Mif70] C. J. Mifsud. A multiple-precision division algorithm. Communications of the ACM, 13(11):666– 668, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See addendum [MB73].

Mills:1970:SDP

[Mil70] H. D. Mills. Syntax-directed documentation for PL360. Communications of the ACM, 13(4):216–222, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Millstein:1973:CSI

[Mil73] Robert E. Millstein. Control structures in Illiac IV Fortran. Communications of the ACM, 16(10): 621–627, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Miller:1974:LOP

[Mil74] Perry Lowell Miller. A locallyorganized parser for spoken input. Communications of the ACM, 17 (11):621–630, November 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mills:1975:CNM

[Mil75a] Harlan D. Mills. Corrigendum: “The New Math of Computer Programming”. Communications of the ACM, 18(5):280, May 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Mil75b].

Mills:1975:NMC

[Mil75b] Harlan D. Mills. The new math of computer programming. Communications of the ACM, 18(1):43–48, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder. See corrigendum [Mil75a].

Millen:1976:SKV

[Mil76] Jonathan K. Millen. Security kernel validation in practice. Communications of the ACM, 19(5):243– 250, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Milman:1977:AOD

[Mil77] Y. Milman. An approach to optimal design of storage parameters in databases. Communications of the ACM, 20(5):347–350, May 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Minieka:1974:CSS

[Min74] Edward Minieka. On computing sets of shortest paths in a graph. Communications of the ACM, 17 (6):351–353, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Minsky:1976:IRP

[Min76] Naftaly H. Minsky. Intentional resolution of privacy protection in database systems. Communications of the ACM, 19(3):148–159, March 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Misunas:1973:PNS

[Mis73] David Misunas. Petri nets and speed independent design. Communications of the ACM, 16(8): 474–481, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Misra:1975:RG

[Mis75] Jayadev Misra. Remark on “Algorithm 246: Graycode [Z]”. ACM Transactions on Mathematical Software, 1(3):285, September 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Boo64, Er85].

Muraoka:1973:TRS

[MK73] Yoichi Muraoka and David J. Kuck.

On the time required for a sequence of matrix products. Communications of the ACM, 16(1):22–26, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mor:1972:NAA

[ML72] Moshe Mor and T. Lamdan. A new approach to automatic scanning of contour maps. Communications of the ACM, 15(9):809–812, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morgan:1977:OPD

[ML77] Howard L. Morgan and K. Dan Levin. Optimal program and data locations in computer networks. Communications of the ACM, 20 (5):315–322, May 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mathews:1970:GPC

[MM70] M. V. Mathews and F. R. Moore. GROOVE — a program to compose, store, and edit functions of time. Communications of the ACM, 13(12):715–721, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Machura:1973:AAR

[MM73] Marek Machura and Andrzej Mulawa. ACM Algorithm 450: Rosenbrock function minimization [E4]. Communications of the ACM, 16 (8):482–483, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [Bul74, KK74, Ell74a, Dav76].

Montanelli:1976:SWM

[MM76] Richard G. Montanelli, Jr. and Sandra A. Mamrak. Status of women and minorities in academic computer science. Communications of the ACM, 19(10):578–581, October 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mamrak:1978:CSF

[MM78a] Sandra A. Mamrak and Richard G.

Montanelli, Jr. Computer science faculties: The current status of minorities and woman. Communications of the ACM, 21(2):115–119, February 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martelli:1978:ODT

[MM78b] Alberto Martelli and Ugo Montanari. Optimizing decision trees through heuristically guided search. Communications of the ACM, 21 (12):1025–1039, December 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mickunas:1978:AER

[MM78c] M. Dennis Mickunas and John A.

Modry. Automatic error recovery for LR parsers. Communications of the ACM, 21(6):459–465, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martin:1972:OBT

[MN72] William A. Martin and D. N. Ness.

Optimizing binary trees grown with a sorting algorithm. Communications of the ACM, 15(2):88–93, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McFarlan:1973:CRG

[MN73] F. Warren McFarlan and Richard L.

Nolan. Curriculum recommendations for graduate professional programs in information systems: Recommended addendum on information systems administration. Communications of the ACM, 16(7): 439–441, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Manna:1973:IMP

[MNV73] Zohar Manna, Stephen Ness, and Jean Vuillemin. Inductive methods for proving properties of programs. Communications of the ACM, 16 (8):491–502, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moler:1972:AAL

[Mol72a] Cleve B. Moler. ACM Algorithm 423: Linear equation solver [F4]. Communications of the ACM, 15(4):274, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moler:1972:MCF

[Mol72b] Cleve B. Moler. Matrix computations with Fortran and paging. Communications of the ACM, 15 (4):268–270, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Montanari:1970:NML

[Mon70] Ugo Montanari. A note on minimal length polygonal approximation to a digitized contour. Communications of the ACM, 13(1):41–47, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Montanari:1971:ODC

[Mon71] Ugo Montanari. On the optimal detection of curves in noisy pictures. Communications of the ACM, 14 (5):335–345, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moorer:1972:MCC

[Moo72] James Anderson Moorer. Music and computer composition. Communications of the ACM, 15(2):104–113, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morgan:1970:IBO

[Mor70a] H. L. Morgan. An interrupt based organization for management information systems. Communications of the ACM, 13(12):734–739, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morgan:1970:SCS

[Mor70b] H. L. Morgan. Spelling correction in systems programs. Communications of the ACM, 13(2):90–94, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1971:PSS

[Mor71a] Robert Morris. Papers from the 2nd symposium on symbolic and algebraic manipulation. Communications of the ACM, 14(8):509–510, August 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1971:ARI

[Mor71b] James H. Morris, Jr. Another recursion induction principle. Communications of the ACM, 14(5):351– 354, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1972:DPT

[Mor72a] James B. Morris. Demand paging through utilization of working sets on the Maniac II. Communications of the ACM, 15(10):867–872, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1972:BVW

[Mor72b] James H. Morris, Jr. A bonus from van Wijngaarden’s device. Communications of the ACM, 15(8):773, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1973:PPL

[Mor73] James H. Morris, Jr. Protection in programming languages. Communications of the ACM, 16(1):15–21, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morgan:1974:OSA

[Mor74] Howard Lee Morgan. Optimal space allocation on disk storage devices. Communications of the ACM, 17 (3):139–142, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1976:RDF

[Mor76] Robert Morris. Remark on “Algorithm 490: The Dilogarithm Function of a Real Argument [S22]”. ACM Transactions on Mathematical Software, 2(1):112, March 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [GZ75].

Morris:1978:TSE

[Mor78a] F. Lockwood Morris. A time- and space-efficient garbage compaction algorithm. Communications of the ACM, 21(8):662–665, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1978:CLN

[Mor78b] Robert Morris. Counting large numbers of events in small registers. Communications of the ACM, 21(10):840–842, October 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1979:CGC

[Mor79] F. L. Morris. On a comparison of garbage collection techniques. Communications of the ACM, 22 (10):571, October 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moses:1971:ASG

[Mos71a] Joel Moses. Algebraic simplification: A guide for the perplexed. Communications of the ACM, 14 (8):527–537, August 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moses:1971:SIS

[Mos71b] Joel Moses. Symbolic integration: The stormy decade. Communications of the ACM, 14(8):548–560, August 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moses:1972:TGT

[Mos72] Joel Moses. Toward a general theory of special functions. Communications of the ACM, 15(7):550– 554, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Motzkin:1979:UNM

[Mot79] Dalia Motzkin. Use of normal multiplication tables for information storage and retrieval. Communications of the ACM, 22(3):193–207, March 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malcolm:1974:FMS

[MP74] Michael A. Malcolm and John Palmer. A fast method for solving a class of tridiagonal linear systems. Communications of the ACM, 17 (1):14–17, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Montangero:1977:TLC

[MPT77] Carlo Montangero, Giuliano Pacini, and Franco Turini. Two-level control structure for nondeterministic programming. Communications of the ACM, 20(10):725–730, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mendelson:1979:OSA

[MPY79] Haim Mendelson, Joseph S. Pliskin, and Uri Yechiali. Optimal storage allocation for serial files. Communications of the ACM, 22(2):124–130, February 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Muthukrishnan:1970:CDT

[MR70] C. R. Muthukrishnan and V. Rajaraman. On the conversion of decision tables to computer programs. Communications of the ACM, 13 (6):347–351, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moore:1972:NMS

[MR72] John Moore and Prentiss Robinson. New method for the solution of the Cauchy problem for parabolic equations. Communications of the ACM, 15(12):1050–1052, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercer:1973:AGP

[MR73] Andrew Mercer and Azriel Rosenfeld. An array grammar programming system. Communications of the ACM, 16(5):299–305, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKay:1974:AAT

[MR74] John McKay and E. Regener. ACM Algorithm 482: Transitivity sets [G7]. Communications of the ACM, 17(8):470, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marcus:1975:EDT

[MR75] Marvin Marcus and Herbert Robinson. Elementary divisors of tensor products. Communications of the ACM, 18(1):36–39, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morel:1979:GOS

[MR79] É. Morel and C. Renvoise. Global optimization by suppression of partial redundancies. Communications of the ACM, 22(2):96–103, February 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Data Flow Analysis.

Moler:1970:ISR

[MS70] Cleve B. Moler and Louis P. Solomon. Integrating square roots. Communications of the ACM, 13 (9):556–557, September 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCalla:1972:MMU

[MS72] Gordon I. McCalla and Jeffrey R. Sampson. MUSE: A model to understand simple English. Communications of the ACM, 15(1):29–40, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mickunas:1973:PGS

[MS73] M. Dennis Mickunas and Victor B. Schneider. A parser-generating system for constructing compressed compilers. Communications of the ACM, 16(11):669–676, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maruyama:1976:ORD

[MS76] K. Maruyama and S. E. Smith. Optimal reorganization of distributed space disk files. Communications of the ACM, 19(11):634–642, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Manna:1977:OAR

[MS77a] Zohar Manna and Adi Shamir. The optimal approach to recursive programs. Communications of the ACM, 20(11):824–831, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maruyama:1977:ADA

[MS77b] K. Maruyama and S. E. Smith.

Analysis of design alternatives for virtual memory indexes. Communications of the ACM, 20(4):245–254, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKenney:1971:SCO

[MT71] J. L. McKenney and Fred M. Tonge.

The state of computer-oriented curricula in business schools 1970. Communications of the ACM, 14 (7):443–448, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1979:PSC

[MT79] Robert Morris and Ken Thompson. Password security — A case history. Communications of the ACM, 22 (11):594–597, November 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mullin:1971:RUS

[Mul71] James K. Mullin. Retrieval-update speed tradeoffs using combined indices. Communications of the ACM, 14(12):775–776, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mullin:1972:IIS

[Mul72] James K. Mullin. An improved indexed-sequential access method using hashed overflow. Communications of the ACM, 15(5):301– 307, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Muller:1977:AMD

[Mul77] Mervin E. Muller. An approach to multidimensional data array processing by computer. Communications of the ACM, 20(2):63–77, February 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Manna:1972:FPA

[MV72] Zohar Manna and Jean Vuillemin. Fix point approach to the theory of computation. Communications of the ACM, 15(7):528–536, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Manna:1971:TAP

[MW71a] Zohar Manna and Richard J. Waldinger. Toward automatic program synthesis. Communications of the ACM, 14(3):151–165, March 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McVitie:1971:AAT

[MW71b] D. G. McVitie and L. B. Wilson. ACM Algorithm 411: Three procedures for the stable marriage problem [H]. Communications of the ACM, 14(7):491–492, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McVitie:1971:SMP

[MW71c] D. G. McVitie and L. B. Wilson. The stable marriage problem. Communications of the ACM, 14 (7):486–490, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1977:SI

[MW77] James H. Morris, Jr. and Ben Wegbreit. Subgoal induction. Communications of the ACM, 20(4): 209–222, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Manna:1978:SSB

[MW78] Zohar Manna and Richard J. Waldinger. Is ‘sometime’ sometimes better than ‘always’? (intermittent assertions in proving program correctness). Communications of the ACM, 21(2):159–172, February 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Myers:1978:CEP

[Mye78] Glenford J. Myers. A controlled experiment in program testing and code walkthroughs/inspections. Communications of the ACM, 21 (9):760–768, September 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Naur:1975:PLN

[Nau75] Peter Naur. Programming languages, natural languages, and mathematics. Communications of the ACM, 18(12):676–683, December 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Papers from the Second ACM Symposium on Principles of Programming Languages (Palo Alto, Calif., 1975).

Nelson:1979:FRS

[Nel79] Lloyd S. Nelson. Further remark on stably updating mean and standard deviation estimates. Communications of the ACM, 22(8):483, August 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Newman:1971:DP

[New71] William M. Newman. Display procedures. Communications of the ACM, 14(10):651–660, October 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ng:1970:CAE

[Ng70] E. N. Ng. Certification of Algorithm 385, Exponential Integral Ei(x). Communications of the ACM, 13 (7):448–449, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ness:1970:CEG

[NGMM70] D. N. Ness, R. S. Green, W. A.

Martin, and G. A. Moulton. Computer education in a graduate school of management. Communications of the ACM, 13(2):110– 114, 19, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nielsen:1970:ACR

[Nie70] N. R. Nielsen. Allocation of computer resources. is pricing the answer. Communications of the ACM, 13(8):467–474, August 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nielsen:1971:AST

[Nie71a] Norman R. Nielsen. An analysis of some time-sharing techniques. Communications of the ACM, 14 (2):79–90, February 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nielsen:1971:MRC

[Nie71b] Norman R. Nielsen. The merit of regional computing networks. Communications of the ACM, 14 (5):319–326, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Niessner:1972:RAE

[Nie72] Herbert Niessner. Remark on “Algorithm 343: Eigenvalues and eigenvectors of a real general matrix”. Communications of the ACM, 15(6):466, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [GB68].

Nielsen:1977:DMA

[Nie77] Norman R. Nielsen. Dynamic memory allocation in computer simulation. Communications of the ACM, 20(11):864–873, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nikolai:1973:CAD

[Nik73] Paul J. Nikolai. Certification of “Algorithm 407: DIFSUB for solution of ordinary differential equations”. Communications of the ACM, 16(7):448, July 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gea71a].

Nilson:1970:CSU

[Nil70] E. N. Nilson. Cubic splines on uniform meshes. Communications of the ACM, 13(4):255–258, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nunamaker:1976:CAA

[NKHS76] Jay F. Nunamaker, Jr., Benn R. Konsynski, Jr., Thomas Ho, and Carl Singer. Computer-aided analysis and design of information systems. Communications of the ACM, 19(12):674–687, December 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Noll:1971:SDC

[Nol71] A. Michael Noll. Scanned-display computer graphics. Communications of the ACM, 14(3):143–150, March 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nolan:1973:MCR

[Nol73] Richard L. Nolan. Managing the computer resource: A stage hypothesis. Communications of the ACM, 16(7):399–405, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nolan:1977:ECU

[Nol77] Richard L. Nolan. Effects of chargeout on user/manager attitudes. Communications of the ACM, 20 (3):177–185, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nemeth:1971:UPM

[NR71] Alan G. Nemeth and Paul D.

Rovner. User program measurement in a time-shared environment. Communications of the ACM, 14 (10):661–666, October 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Newell:1976:CSE

[NS76] Allen Newell and Herbert A. Simon. Computer science as empirical inquiry: Symbols and search. Communications of the ACM, 19 (3):113–126, March 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Needham:1978:UEA

[NS78] Roger M. Needham and Michael D. Schroeder. Using encryption for authentication in large networks of computers. Communications of the ACM, 21(12):993–999, December 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nagy:1970:NTH

[NT70] G. Nagy and N. Tuong. Normalization techniques for handprinted numerals. Communications of the ACM, 13(8):475–481, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neufeld:1974:GCC

[NT74] G. A. Neufeld and John Tartar. Graph coloring conditions for the existence of solutions to the timetable problem. Communications of the ACM, 17(8):450–453, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nagy:1979:APM

[NW79] George Nagy and Sharad G. Wagle. Approximation of polygonal maps by cellular maps. Communications of the ACM, 22(9):518– 525, September 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Obradovic:1973:RAM

[Obr73] D. Obradović. Remark on “Algorithm 333: Minit algorithm for linear programming”. Communications of the ACM, 16(5):310, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [SS68, HOK73].

Oldehoeft:1971:EAI

[OC71a] Arthur E. Oldehoeft and Samuel D.

Conte. Experiments with an automated instructional system for numerical methods. Communications of the ACM, 14(10):643–650, October 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Olle:1971:IFA

[OC71b] T. William Olle (Chairman) and Codasyl Systems Committee. Introduction to “feature analysis of generalized data base management systems”. Communications of the ACM, 14(5):308–318, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Owicki:1976:VPP

[OG76] Susan S. Owicki and David Gries.

Verifying properties of parallel programs: An axiomatic approach. Communications of the ACM, 19 (5):279–285, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Papers from the Fifth ACM Symposium on Operating Systems Principles (Univ. Texas, Austin, Tex., 1975).

Oldehoeft:1972:MCP

[OH72] Arthur E. Oldehoeft and Maurice H. Halstead. Maximum computing power and cost factors in the centralization problem. Communications of the ACM, 15(2):94–96, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

OReagan:1972:CAC

[O’R72] Robert T. O’Reagan. Computer assigned codes from verbal responses. Communications of the ACM, 15 (6):455–459, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ord-Smith:1968:AAG

[OS68] R. J. Ord-Smith. ACM Algorithm 323: Generation of permutations in lexicographic order. Communications of the ACM, 11(2):117, February 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Lei69] and remark [Roy73c].

Oden:1972:MMC

[OS72] P. H. Oden and Gerald S. Shedler. A model of memory contention in a paging machine. Communications of the ACM, 15(8):761–771, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Ottmann:1978:RBT

[OSW78] Thomas Ottmann, Hans-Werner Six, and Derick Wood. Right brother trees. Communications of the ACM, 21(9):769–776, September 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Phillips:1972:DPA

[PA72] J. Richard Phillips and H. C.

Adams. Dynamic partitioning for array languages. Communications of the ACM, 15(12):1023–1032, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Paciorek:1970:AAE

[Pac70] K. A. Paciorek. ACM Algorithm 385: Exponential integral E_{i}(x). Communications of the ACM, 13 (7):446–447, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Fri72].

Pack:1973:EMC

[Pac73] Charles D. Pack. The effects of multiplexing on a computercommunications system. Communications of the ACM, 16(3):161–168, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pager:1970:NSP

[Pag70] D. Pager. A number system for the permutations. Communications of the ACM, 13(3):193, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pager:1972:PCB

[Pag72] David Pager. A proposal for a computer-based interactive scientific community. Communications of the ACM, 15(2):71–75, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pager:1973:PCC

[Pag73] David Pager. On the problem of communicating complex information. Communications of the ACM, 16(5):275–281, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Page:1974:AAM

[Pag74a] R. L. Page. ACM Algorithm 479: A minimal spanning tree clustering method [Z]. Communications of the ACM, 17(6):321–323, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Mag75, WGL76].

Page:1974:GRC

[Pag74b] R. L. Page. Remark on “Algorithm 425: Generation of random correlated normal variables”. Communications of the ACM, 17(6):325, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [HK72].

Pankiewicz:1970:AAD

[Pan70] W. Pankiewicz. ACM Algorithm 378: Discretized Newton-like method for solving a system of simultaneous nonlinear equations. Communications of the ACM, 13 (4):259–260, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:1972:CUD

[Par72a] David Lorge Parnas. On the criteria to be used in decomposing systems into modules. Communications of the ACM, 15(12):1053–1058, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:1972:TSM

[Par72b] David Lorge Parnas. A technique for software module specification with examples. Communications of the ACM, 15(5):330–336, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:1975:SCS

[Par75] David Lorge Parnas. On a solution to the cigarette smoker’s problem (without conditional statements). Communications of the ACM, 18 (3):181–183, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parker:1979:VEI

[Par79] Donn B. Parker. Vulnerabilities of EFTs to intentionally caused losses. Communications of the ACM, 22 (12):654–660, December 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Paton:1971:ABC

[Pat71a] Keith Paton. An algorithm for the blocks and cut nodes of a graph. Communications of the ACM, 14 (7):468–475, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Pat71b].

Paton:1971:CAB

[Pat71b] Keith Paton. Corrigendum: “An algorithm for the blocks and cutnodes of a graph”. Communications of the ACM, 14(9):592, September 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pat71a].

Patrick:1972:HPA

[Pat72] Merrell L. Patrick. A highly parallel algorithm for approximating all zeros of a polynomial with only real zeros. Communications of the ACM, 15(11):952–955, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:1973:AAA

[Pat73] T. N. L. Patterson. ACM Algorithm 468: Algorithm for automatic numerical integration over a finite interval [D1]. Communications of the ACM, 16(11):694–699, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Payne:1970:FTP

[Pay70] W. H. Payne. Fortran Tausworthe pseudorandom number generator. Communications of the ACM, 13 (1):57, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Payne:1973:GEP

[Pay73] William H. Payne. Graduate education: The ph.D. glut. Communications of the ACM, 16(3):181–182, March 1973. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). See response and rebuttal [FMP74].

Peterson:1978:SOT

[PBH78] James L. Peterson, James R. Bitner, and John H. Howard. The selection of optimal tab settings. Communications of the ACM, 21 (12):1004–1007, December 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pearson:1990:FHV

[Pea90] Peter K. Pearson. Fast hashing of variable-length text strings. Communications of the ACM, 33 (6):677–680, June 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Dit76, Dit91, Lit91, Pea91, Sav91].

Pearson:1991:NFH

[Pea91] Peter Pearson. Note on fast hashing of variable length text strings. Communications of the ACM, 34 (11):120, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Responds to several comments [Dit91, Lit91, Sav91] on the author’s earlier paper [Pea90], and early work [Dit76].

Peemoller:1983:CBM

[Pee83] Jürgen Peemöller. A correction to Brélaz’s modification of Brown’s coloring algorithm. Communications of the ACM, 26(8):595–597, August 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peto:1970:RAM

[Pet70] R. Peto. Remark on “Algorithm 347: An efficient algorithm for sorting with minimal storage [M1]”. Communications of the ACM, 13 (10):624, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [GR70, Sin69a].

Pratt:1971:LEG

[PF71] Terrence W. Pratt and Daniel P. Friedman. A language extension for graph processing, and its formal semantics. Communications of the ACM, 14(7):460–467, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Prieve:1976:VOV

[PF76] Barton G. Prieve and Robert S. Fabry. VMIN — an optimal variable-space page replacement algorithm. Communications of the ACM, 19(5):295–297, May 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Popek:1978:MVD

[PF78] Gerald J. Popek and David A. Farber. Model for verification of data security in operating systems. Communications of the ACM, 21(9): 737–749, September 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pfefferkorn:1975:HPS

[Pfe75] Charles E. Pfefferkorn. Heuristic problem solving design system for equipment or furniture layouts. Communications of the ACM, 18 (5):286–297, May 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Popek:1974:FRV

[PG74] Gerald J. Popek and Robert P.

Goldberg. Formal requirements for virtualizable third generation architectures. Communications of the ACM, 17(7):412–421, July 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pike:1965:AAP

[PH65] M. C. Pike and I. D. Hill. ACM Algorithm 266: Pseudo-random numbers. Communications of the ACM, 8(10):605–606, October 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Sul72].

Preparata:1977:CHF

[PH77] Franco P. Preparata and S. J. Hong.

Convex hulls of finite sets of points in two and three dimensions. Communications of the ACM, 20(2): 87–93, February 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:1972:CDP

[PHH72] David Lorge Parnas, A. Nico Habermann, and R. C. Holt. Comment on deadlock prevention method. Communications of the ACM, 15(9): 840–841, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Phillips:1971:NBO

[Phi71] David L. Phillips. A note on best one-sided approximations. Communications of the ACM, 14(9):598– 600, September 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Phong:1975:ICG

[Pho75] Bui Tuong Phong. Illumination for computer generated pictures. Communications of the ACM, 18 (6):311–317, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Perl:1978:ISL

[PIA78] Yehoshua Perl, Alon Itai, and Haim Avni. Interpolation search — A log log N search. Communications of the ACM, 21(7):550–553, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Piessens:1973:AAG

[Pie73] Robert Piessens. ACM Algorithm 453: Gaussian quadrature formulas for Bromwich’s integral [D1]. Communications of the ACM, 16 (8):486–487, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Piessens:1974:AAC

[Pie74a] Robert Piessens. ACM Algorithm 473: Computation of Legendre series coefficients [C6]. Communications of the ACM, 17(1):25, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Piessens:1974:RAM

[Pie74b] Robert Piessens. Remark on “Algorithm 400: Modified Haåvie integration”. Communications of the ACM, 17(6):324, June 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wal70].

Piessens:1974:RAC

[Pie74c] Robert Piessens. Remark on “Algorithm 418: Calculation of Fourier integrals”. Communications of the ACM, 17(6):324, June 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ein72a, Ein72b].

Piepmeier:1975:OBR

[Pie75] William F. Piepmeier. Optimal balancing of I/O requests to disks. Communications of the ACM, 18 (9):524–527, September 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pinter:1970:DBF

[Pin70] C. C. Pinter. On decomposing a Boolean function and simplifying the components separately. Communications of the ACM, 13(12): 758, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peterson:1973:CWR

[PKT73] W. W. Peterson, Tadao Kasami, and Nobuki Tokura. On the capabilities of while, repeat, and exit statements. Communications of the ACM, 16(8):503–512, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Purdom:1972:AAI

[PM72] Paul Walton Purdom, Jr. and Edward F. Moore. ACM Algorithm 430: Immediate predominators in a directed graph [H]. Communications of the ACM, 15(8):777–778, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Piessens:1975:RCA

[PM75] Robert Piessens and Irene Mertens. Remark and certification on “Algorithm 446: Ten subroutines for the manipulation of Chebyshev series”. Communications of the ACM, 18 (5):276, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Bro73].

Payne:1978:OEN

[PM78] W. H. Payne and K. L. McMillen. Orderly enumeration of nonsingular binary matrices applied to text encryption. Communications of the ACM, 21(4):259–263, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peterson:1977:BS

[PN77] James L. Peterson and Theodore A. Norman. Buddy systems. Communications of the ACM, 20(6):421– 431, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pohl:1972:SPC

[Poh72] Ira Pohl. A sorting problem and its complexity. Communications of the ACM, 15(6):462–464, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pollack:1971:CCD

[Pol71] S. L. Pollack. Comment on the conversion of decision tables to computer programs. Communications of the ACM, 14(1):52, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pomeranz:1970:NAP

[Pom70] J. E. Pomeranz. Note on an anomaly in paging. Communications of the ACM, 13(7):451, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pomeranz:1974:AAE

[Pom74] John Pomeranz. ACM Algorithm 487: Exact cumulative distribution of the Kolmogorov–Smirnov statistic for small samples. Communications of the ACM, 17(12): 703–704, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Pom76].

Pomeranz:1976:REC

[Pom76] J. Pomeranz. Remark on “Algorithm 487: Exact Cumulative Distribution of the Kolmogorov– Smirnov Statistic for Small Samples [S14]”. ACM Transactions on Mathematical Software, 2(1): 111, March 1976. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Pom74].

Pooch:1976:PSI

[Poo76] Udo W. Pooch. Papers from SIG-GRAPH 76 — introduction. Communications of the ACM, 19(10):

541–??, 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Probst:1972:ISI

[POW^{+}72] G. G. Probst, Robert S. Oelman, Sam Wyly, C. Lester Hogan, F. T. Cary, Gordon Bell, and Donald P. Kircher. As the industry sees it. Communications of the ACM, 15 (7):506–517, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peleg:1979:BSC

[PR79] Shmuel Peleg and Azriel Rosenfeld. Breaking substitution ciphers using a relaxation algorithm. Communications of the ACM, 22(11): 598–605, November 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Press:1974:AMC

[Pre74] Laurance I. Press. Arguments for a moratorium on the construction of a community information utility. Communications of the ACM, 17 (12):674–678, December 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Preparata:1979:ORT

[Pre79] Franco P. Preparata. An optimal real-time algorithm for planar convex hulls. Communications of the ACM, 22(7):402–405, July 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Prieve:1973:UPR

[Pri73] Barton G. Prieve. Using page residency to select the working set parameter. Communications of the ACM, 16(10):619–620, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Proll:1971:FEA

[Pro71] L. G. Proll. Further evidence for the analysis of algorithms for the zero-one programming problem. Communications of the ACM, 14 (1):46–47, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Proll:1972:RAA

[Pro72] L. G. Proll. Remark on “ACM Algorithm 370: General random number generator [G5]”. Communications of the ACM, 15(6):467–468, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [But70].

Proll:1974:RAC

[Pro74] L. G. Proll. Remark on “Algorithm 431: A computer routine for quadratic and linear programming problems”. Communications of the ACM, 17(10):590, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Rav72, Rav74].

Parnas:1975:UCT

[PS75] David Lorge Parnas and Daniel P.

Siewiorek. Use of the concept of transparency in the design of hierarchically structured systems. Communications of the ACM, 18 (7):401–408, July 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pike:1976:RIB

[PS76] Malcolm C. Pike and Jennie SooHoo. Remark on “Algorithm 179: Incomplete Beta Ratio [S14]”. ACM Transactions on Mathematical Software, 2(2):207–208, June 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Lud63].

Purdy:1974:HSL

[Pur74] George B. Purdy. A high security log-in procedure. Communications of the ACM, 17(8):442–445, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pomeranz:1970:CMP

[PW70] J. E. Pomeranz and R. L. Weil, Jr. The cyclical majority problem. Communications of the ACM, 13 (4):251–254, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peters:1975:SGJ

[PW75] G. Peters and James Hardy Wilkinson. On the stability of Gauss– Jordan elimination with pivoting. Communications of the ACM, 18 (1):20–24, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder.

Rabin:1977:CC

[Rab77] Michael O. Rabin. Complexity of computations. Communications of the ACM, 20(9):625–633, September 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Rab78].

Rabin:1978:CCC

[Rab78] Michael O. Rabin. Corrigendum: “Complexity of computations”. Communications of the ACM, 21(3), March 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Rab77].

Radke:1970:UQR

[Rad70a] C. E. Radke. The use of quadratic residue research. Communications of the ACM, 13(2):103–105, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Raduchel:1970:EHB

[Rad70b] W. J. Raduchel. Efficient handling of binary data. Communications of the ACM, 13(12):758–759, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Raduchel:1971:RAS

[Rad71] William J. Raduchel. Remark on “Algorithm 316: Solution of simultaneous nonlinear equations [C5]”. Communications of the ACM, 14(7):493–??, 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Bro67].

Rader:1974:MCS

[Rad74] Gary M. Rader. Method for composing simple traditional music by computer. Communications of the ACM, 17(11):631–638, November 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ralston:1972:NTY

[Ral72] Anthony Ralston. The next two years (ACM President’s letter). Communications of the ACM, 15(7):499–500, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ravindran:1972:AAC

[Rav72] Arunachalam Ravindran. ACM Algorithm 431: A computer routine for quadratic and linear programming problems [H]. Communications of the ACM, 15(9):818– 820, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Rav74, Pro74].

Ravindran:1974:RAC

[Rav74] Arunachalam Ravindran. Remark on “Algorithm 431: A computer routine for quadratic and linear programming problems”. Communications of the ACM, 17(3):157–??, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Rav72, Pro74].

Raymond:1976:LLA

[Ray76] J. Raymond. LG: A language for analytic geometry. Communications of the ACM, 19(4):182–187, April 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robbins:1970:ICS

[RB70] M. F. Robbins and J. D. Beyer. An interactive computer system using graphical flowchart input. Communications of the ACM, 13(2): 115–119, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Redish:1970:RAE

[Red70] K. A. Redish. Remark on Algorithm 385, Exponential Integral Ei(x). Communications of the ACM, 13 (12):750, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pac70, Fri72].

Redish:1971:ACL

[Red71] K. A. Redish. Algorithms: Comment on London’s certification of Algorithm 245. Communications of the ACM, 14(1):50–51, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reingold:1973:NLM

[Rei73a] Edward M. Reingold. A nonrecursive list moving algorithm. Communications of the ACM, 16(5):305– 307, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reinsch:1973:AAE

[Rei73b] Christian H. Reinsch. ACM Algorithm 464: Eigenvalues of a real, symmetric, tridiagonal matrix [F2]. Communications of the ACM, 16 (11):689, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Relph:1963:AAH

[Rel63] A. P. Relph. ACM Algorithm 191: Hypergeometric. Communications of the ACM, 6(7):388–389, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Kop74].

Revens:1972:FTF

[Rev72] Lee Revens. The first twenty-five years: ACM 1947–1962. Communications of the ACM, 15(7):485– 490, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Reynolds:1970:GST

[Rey70] John C. Reynolds. GEDANKEN — A simple typeless language based on the principle of completeness and the reference concept. Communications of the ACM, 13(5):308– 319, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reynolds:1979:RAA

[Rey79] John C. Reynolds. Reasoning about arrays. Communications of the ACM, 22(5):290–299, May 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rubinstein:1972:CBT

[RF72] Richard Rubinstein and Julian Feldman. A controller for a Braille terminal. Communications of the ACM, 15(9):841–842, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ruschitzka:1977:UAS

[RF77] Manfred Ruschitzka and Robert S.

Fabry. A unifying approach to scheduling. Communications of the ACM, 20(7):469–477, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosin:1972:ERM

[RFE72] Robert F. Rosin, Gideon Frieder, and Richard H. Eckhouse, Jr. Environment for research in microprogramming and emulation. Communications of the ACM, 15(8): 748–760, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramamoorthy:1971:SOE

[RG71] C. V. Ramamoorthy and M. J. Gonzalez. Subexpression ordering in the execution of arithmetic expressions. Communications of the ACM, 14(7):479–485, July 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ragland:1973:CAG

[RG73] Larry C. Ragland and Donald I.

Good. Certification of “Algorithm 386: Greatest common divisor of n integers and multipliers”. Communications of the ACM, 16(4):257, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Bra70a].

Read:1972:NGR

[RH72] Ronald C. Read and K. Harada.

A note on the generation of rosary permutations. with a response by K. Harada. Communications of the ACM, 15(8):775, August 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles in honor of George E. Forsythe (ACM Student Competition Award Papers).

Richards:1971:HKA

[Ric71] Donald L. Richards. How to keep the addresses short. Communications of the ACM, 14(5):346–349, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rice:1972:PFS

[Ric72a] John R. Rice. On the present and future of scientific computation. Communications of the ACM, 15 (7):637–639, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Richman:1972:AEA

[Ric72b] Paul L. Richman. Automatic error analysis for determining precision. Communications of the ACM, 15 (9):813–817, September 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Richman:1973:VPE

[Ric73] Paul L. Richman. Variableprecision exponentiation. Communications of the ACM, 16(1):38–40, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ripley:1978:SRO

[Rip78] G. David Ripley. A simple recoveryonly procedure for simple precedence parsers. Communications of the ACM, 21(11):928–930, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rivest:1976:SOS

[Riv76] Ronald L. Rivest. On selforganizing sequential search heuristics. Communications of the ACM, 19(2):63–67, February 1976. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Richardson:1973:AAC

[RK73] Joel A. Richardson and J. L.

Kuester. ACM Algorithm 454: The complex method for constrained optimization [E4]. Communications of the ACM, 16(8):487–489, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [She74].

Reiser:1975:HRE

[RK75] Martin Reiser and H. Kobayashi.

Horner’s rule for the evaluation of general closed queueing networks. Communications of the ACM, 18 (10):592–593, October 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reed:1979:SES

[RK79] David P. Reed and Rajendra K.

Kanodia. Synchronization with eventcounts and sequencers. Communications of the ACM, 22(2): 115–123, February 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rokne:1971:CIA

[RL71] J. Rokne and P. Lancaster. Complex interval arithmetic. Communications of the ACM, 14(2):111–112, February 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rothnie:1974:ABF

[RL74] James B. Rothnie, Jr. and Tomas Lozano. Attribute based file organization in a paged memory environment. Communications of the ACM, 17(2):63–69, February 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robinson:1977:PTH

[RL77] Lawrence Robinson and Karl N. Levitt. Proof techniques for hierarchically structured programs. Communications of the ACM, 20(4): 271–283, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robertson:1972:RAT

[Rob72] J. Douglas Robertson. Remark on “Algorithm 398: Tableless date conversion”. Communications of the ACM, 15(10):918–??, 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Sto70].

Robson:1977:BSA

[Rob77] J. M. Robson. A bounded storage algorithm for copying cyclic structures. Communications of the ACM, 20(6):431–433, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rochfeld:1971:NLT

[Roc71] Arnold Rochfeld. New LISP techniques for a paging environment. Communications of the ACM, 14 (12):791–795, December 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rokne:1970:CNP

[Rok70] J. Rokne. Condition numbers of Pei matrices. Communications of the ACM, 13(11):699, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rokne:1973:AES

[Rok73] J. Rokne. Automatic errorbounds for simple zeros of analytic functions. Communications of the ACM, 16(2):101–104, February 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosen:1972:PSL

[Ros72] Saul Rosen. Programming systems and languages (1965–1975). Communications of the ACM, 15 (7):591–600, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosin:1973:TAP

[Ros73] Robert F. Rosin. Teaching “about programming”. Communications of the ACM, 16(7):435–438, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ross:1976:HHS

[Ros76] Douglas T. Ross. Homilies for humble standards (computer applications). Communications of the ACM, 19(11):595–600, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosen:1977:HLD

[Ros77] Barry K. Rosen. High-level data flow analysis. Communications of the ACM, 20(10):712–724, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roy:1973:RPR

[Roy73a] M. K. Roy. Reflection-free permutations, rosary permutations and adjacent transposition algorithms. Communications of the ACM, 16 (5):312, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roy:1973:RFP

[Roy73b] Mohit Kumar Roy. Reflectionfree permutations, rosary permutations, and adjacent transposition algorithms. Communications of the ACM, 16(5):312–313, 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roy:1973:RAG

[Roy73c] Mohit Kumar Roy. Remark on “Algorithm 323: Generation of permutations in lexicographic order [G6]”. Communications of the ACM, 16 (9):577–578, September 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [OS68, Lei69].

Roy:1977:NRP

[Roy77a] M. K. Roy. A note on reflection-free permutation enumeration. Communications of the ACM, 20(11): 823, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roy:1977:NRF

[Roy77b] Mohit Kumar Roy. A note on refection-free permutation enumeration. Communications of the ACM, 20(11):823–??, 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robers:1973:AAD

[RR73a] P. D. Robers and S. S. Robers.

ACM Algorithm 458: Discrete linear L_{1} approximation by interval linear programming [E2]. Communications of the ACM, 16(10): 629–631, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rodriguez-Rosell:1973:EWS

[RR73b] Juan Rodriguez-Rosell. Empirical working set behavior. Communications of the ACM, 16(9):556– 560, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rodriguez-Rosell:1973:DIE

[RRD73] Juan Rodriguez-Rosell and Jean-Pierre Dupuy. The design, implementation, and evaluation of a working set dispatcher. Communications of the ACM, 16(4):247–253, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramberg:1972:AMG

[RS72] John S. Ramberg and Bruce W. Schmeiser. Approximate method for generating symmetric random variables. Communications of the ACM, 15(11):987–990, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramberg:1974:AMG

[RS74] John S. Ramberg and Bruce W. Schmeiser. An approximate method for generating asymmetric random variables. Communications of the ACM, 17(2):78–82, February 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rivest:1978:MOD

[RSA78] Ronald L. Rivest, Adi Shamir, and Leonard M. Adleman. A method for obtaining digital signatures and public-key cryptosystems. Communications of the ACM, 21(2): 120–126, February 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). The basics of trap-door functions and the famous RSA public key cryptosystem are presented in this paper.

Ritchie:1974:UTS

[RT74] Dennis M. Ritchie and Ken Thompson. The UNIX time-sharing system. Communications of the ACM, 17(7):365–375, July 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rubin:1976:ETF

[Rub76] Frank Rubin. Experiments in text file compression. Communications of the ACM, 19(11):617– 623, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Russell:1978:CCS

[Rus78] Richard M. Russell. The Cray-1 computer system. Communications of the ACM, 21(1):63–72, January 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Raiha:1979:OIA

[RZ79] Kari-Jouko Räihä and Stuart H.

Zweben. An optimal insertion algorithm for one-sided height-balanced binary search trees. Communications of the ACM, 22(9):508– 512, September 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stockman:1977:EHC

[SA77] G. C. Stockman and Ashok K.

Agrawala. Equivalence of Hough curve detection to template matching. Communications of the ACM, 20(11):820–822, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sale:1972:RAS

[Sal72a] Arthur H. J. Sale. Remark on “Algorithm 393: Special series summation with arbitrary precision”. Communications of the ACM, 15 (6):468–469, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Abd70].

Salton:1972:DDP

[Sal72b] Gerard Salton. Dynamic document processing. Communications of the ACM, 15(7):658–668, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Salzer:1972:OFF

[Sal72c] Herbert E. Salzer. Ordering +- f(+-f(+-f(… +-f(x)..))) when f(x) is positive monotonic. Communications of the ACM, 15(1):45–46, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Salasin:1973:HSI

[Sal73a] John Salasin. Hierarchical storage in information retrieval. Communications of the ACM, 16(5):291– 295, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sale:1973:SMP

[Sal73b] Arthur H. J. Sale. Remark on “Algorithm 408: A sparse matrix package”. Communications of the ACM, 16(5):311, May 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [McN71, Law73].

Salzer:1973:GCS

[Sal73c] Herbert E. Salzer. Gray code and the ± sequence when ±f(±f(±f(⋯± f(x) ⋯))) is ordered. Communications of the ACM, 16(3):180, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Salzer:1973:RSC

[Sal73d] Herbert E. Salzer. A recurrence scheme for converting from one orthogonal expansion into another. Communications of the ACM, 16 (11):705–707, November 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Saltzer:1974:PCI

[Sal74a] Jerome H. Saltzer. Protection and the control of information sharing in Multics. Communications of the ACM, 17(7):388–402, July 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Saltzer:1974:SLM

[Sal74b] Jerome H. Saltzer. A simple linear model of demand paging performance. Communications of the ACM, 17(4):181–186, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sammet:1972:PLH

[Sam72] Jean E. Sammet. Programming languages: History and future. Communications of the ACM, 15 (7):601–610, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sammet:1976:RPL

[Sam76] Jean E. Sammet. Roster of programming languages for 1974–75. Communications of the ACM, 19 (12):655–669, December 1976. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samet:1978:PCH

[Sam78] Hanan Samet. Proving the correctness of heuristically optimized code. Communications of the ACM, 21 (7):570–582, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Savoy:1991:NFH

[Sav91] Jacques Savoy. Note on fast hashing of variable length text strings. Communications of the ACM, 34(11): 118–119, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Suggests an improvement to Pearson’s hashing algorithm [Pea90] that avoids secondary clustering. Exhibits a key set for which Pearson’s algorithm produces alarming clustering. See also comments in [Dit91, Lit91, Pea91] and early work [Dit76].

Stillman:1977:CHS

[SB77] Neil J. Stillman and P. Bruce Berra. A comparison of hardware and software associative memories in the context of computer graphics. Communications of the ACM, 20(5):331–339, May 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sherman:1972:TDM

[SBB72] Stephen Sherman, Forest Baskett III, and James C. Browne. Tracedriven modeling and analysis of CPU scheduling in a multiprogramming system. Communications of the ACM, 15(12):1063–1069, December 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sasser:1970:ASS

[SBGN70] W. E. Sasser, D. S. Burdick, D. A.

Graham, and T. H. Naylor. Application of sequential sampling to simulation. an example inventory model. Communications of the ACM, 13(5):287–296, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schwarcz:1970:DQA

[SBS70] Robert M. Schwarcz, John F.

Burger, and Robert F. Simmons. A deductive question answerer for natural language inference. Communications of the ACM, 13(3):167–183, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1975:OPR

[SC75] John Miles Smith and Philip Yen-Tang Chang. Optimizing the performance of a relational algebra data base interface. Communications of the ACM, 18(10): 568–579, October 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://dblp.uni-trier.de/db/
journals/cacm/SmithC75.html.

Scheuermann:1978:SAP

[SC78] Peter Scheuermann and C. Robert Carlson. Self-assessment procedure V: a self-assessment procedure dealing with database systems. Communications of the ACM, 21(8): 687–693, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sauer:1979:IDD

[SC79] Charles H. Sauer and K. Mani Chandy. The impact of distributions and disciplines on multiple processor systems. Communications of the ACM, 22(1):25–34, January 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schaffer:1970:AAG

[Sch70a] H. E. Schaffer. ACM Algorithm 369: Generator of random numbers satisfying the Poisson distribution. Communications of the ACM, 13(1):49, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schweikert:1970:CEI

[Sch70b] D. G. Schweikert. A comparison of error improvement estimates for adaptive trapezoid integration. Communications of the ACM, 13 (3):163–166, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schmitt:1971:AAD

[Sch71] H. Schmitt. ACM Algorithm 409: Discrete Chebychev curve fit [E2]. Communications of the ACM, 14 (5):355–356, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schrack:1972:RAR

[Sch72a] Günther F. Schrack. Remark on “Algorithm 381: Random vectors uniform in solid angle”. Communications of the ACM, 15(6):468, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Kno70].

Schrack:1972:RAU

[Sch72b] Günther F. Schrack. Remark on “Algorithm 391: Unitary symmetric polynomials [Z]”. Communications of the ACM, 15(1):49–??, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [McK70a].

Schuegraf:1972:RAL

[Sch72c] Ernst Schuegraf. Remark on “Algorithm 176: Least squares surface fit”. Communications of the ACM, 15(12):1073–??, 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Art63].

Schuegraf:1972:RAB

[Sch72d] Ernst Schuegraf. Remark on “Algorithm 195: BANDSOLVE”. Communications of the ACM, 15(12): 1074–??, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Thu63].

Schkolnick:1974:ERT

[Sch74] Mario Schkolnick. The equivalence of reducing transition languages and deterministic languages. Communications of the ACM, 17(9):517– 519, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schwartz:1975:ADS

[Sch75] Jacob T. Schwartz. Automatic data structure choice in a language of very high level. Communications of the ACM, 18(12):722–728, December 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scheifler:1977:AIS

[Sch77] Robert W. Scheifler. An analysis of inline substitution for a structured programming language. Communications of the ACM, 20(9):647– 654, September 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schoene:1978:RMI

[Sch78a] Andrew Y. Schoene. Remark on “Algorithm 435: Modified Incomplete Gamma Function [S14]”. ACM Transactions on Mathematical Software, 4(3):296–304, September 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Ful72].

Schwetman:1978:HSM

[Sch78b] Herbert D. Schwetman. Hybrid simulation models of computer systems. Communications of the ACM, 21(9):718–723, September 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scott:1977:LPL

[Sco77] Dana S. Scott. Logic and programming languages. Communications of the ACM, 20(9):634– 641, September 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Slagle:1970:EMT

[SD70] J. R. Slagle and J. K. Dixon.

Experiments with the M & N tree-searching program. Communications of the ACM, 13(3):147– 154, 159, March 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Severance:1976:PGA

[SD76] Dennis G. Severance and Ricardo A.

Duhne. A Practitioner’s guide to addressing algorithms. Communications of the ACM, 19(6):314–326, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sedelow:1972:LAH

[Sed72] Sally Yeates Sedelow. Language analysis in the humanities. Communications of the ACM, 15(7):644– 647, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sedgewick:1978:IQP

[Sed78] Robert Sedgewick. Implementing Quicksort programs. Communications of the ACM, 21(10): 847–857, October 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Sed79].

Sedgewick:1979:CIQ

[Sed79] Robert Sedgewick. Corrigendum: “Implementing Quicksort Programs”. Communications of the ACM, 22(6):368, June 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Sed78].

Seppanen:1970:AAS

[Sep70] J. J. Seppänen. ACM Algorithm 399: Spanning tree. Communications of the ACM, 13(10):621–622, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Slagle:1971:EAL

[SF71] James R. Slagle and Carl D. Farrell. Experiments in automatic learning for a multipurpose heuristic program. Communications of the ACM, 14(2):91–99, February 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stone:1973:NOS

[SF73] Harold S. Stone and Samuel H. Fuller. On the near-optimality of the shortest-latency-time-first drum scheduling discipline. Communications of the ACM, 16(6):352–353, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also published in/as: Technical Note No.12, DSL.

Shapira:1979:COP

[SF79] Ruth Shapira and Herbert Freeman. The cyclic order property of vertices as an aid in scene analysis. Communications of the ACM, 22(6):368– 375, June 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Saltzer:1970:IM

[SG70] J. H. Saltzer and J. W. Gintell. The instrumentation of Multics. Communications of the ACM, 13(8): 495–500, August 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sager:1975:RLC

[SG75] Naomi Sager and Ralph Grishman. Restriction language for computer grammars of natural language. Communications of the ACM, 18(7):390–400, July 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sutherland:1974:RPC

[SH74] Ivan E. Sutherland and Gary W.

Hodgman. Reentrant polygon clipping. Communications of the ACM, 17(1):32–42, January 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shapiro:1970:LSG

[Sha70] S. C. Shapiro. The list set generator: a construct for evaluating set expressions. Communications of the ACM, 13(12):741–744, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shaw:1974:RCC

[Sha74] Mary Shaw. Reduction of compilation costs through language contraction. Communications of the ACM, 17(5):245–250, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shampine:1975:DLS

[Sha75] L. F. Shampine. Discrete least squares polynomial fits. Communications of the ACM, 18(3):179–180, March 1975. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Shapiro:1977:CRP

[Sha77] Marvin B. Shapiro. Choice of reference points in best-match file searching. Communications of the ACM, 20(5):339–343, May 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shamir:1979:HSS

[Sha79] Adi Shamir. How to share a secret. Communications of the ACM, 22 (11):612–613, November 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shell:1971:OPS

[She71] Donald L. Shell. Optimizing the polyphase sort. Communications of the ACM, 14(11):713–719, November 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [She72].

Shell:1972:COP

[She72] Donald L. Shell. Corrigendum: “Optimizing the Polyphase Sort”. Communications of the ACM, 15 (1):28, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [She71].

Shedler:1973:QMM

[She73] Gerald S. Shedler. A queuing model of a multiprogrammed computer with a two-level storage system. Communications of the ACM, 16(1):3–10, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shere:1974:RAC

[She74] Kenneth D. Shere. Remark on “Algorithm 454: The complex method for constrained optimization”. Communications of the ACM, 17(8):471–??, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [RK73].

Sheil:1978:MST

[She78] B. A. Sheil. Median split trees: A fast lookup technique for frequently occurring keys. Communications of the ACM, 21(11):947–958, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Early work on minimal perfect hash functions.

Shu:1975:CHL

[SHL75] Nan C. Shu, Barron C. Housel, and Vincent Y. Lum. CONVERT: A high level translation definition language for data conversion. Communications of the ACM, 18(10): 557–567, October 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also published in/as: 19 ACM SIGMOD Conf. on the Management of Data, King(ed), May.1975.

Shneiderman:1973:ODB

[Shn73] Ben Shneiderman. Optimum data base reorganization points. Communications of the ACM, 16(6): 362–365, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://dblp.uni-trier.de/db/
journals/cacm/Shneiderman73.
html.

Shneiderman:1978:JSF

[Shn78] Ben Shneiderman. Jump searching: A fast sequential search technique. Communications of the ACM, 21 (10):831–834, October 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shore:1975:ESF

[Sho75] John E. Shore. On the external storage fragmentation produced by first-fit and best-fit allocation strategies. Communications of the ACM, 18(8):433–440, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shore:1977:ABF

[Sho77] John E. Shore. Anomalous behavior of the fifty-percent rule in dynamic memory allocation. Communications of the ACM, 20(11): 812–820, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shostak:1978:ARA

[Sho78] Robert E. Shostak. An algorithm for reasoning about equality. Communications of the ACM, 21 (7):583–585, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shrager:1972:QPN

[Shr72] Richard I. Shrager. Quadratic programming for nonlinear regression. Communications of the ACM, 15 (1):41–45, January 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shwayder:1971:CLE

[Shw71] Keith Shwayder. Conversion of limited-entry decision tables to computer programs. a proposed modification to Pollack’s algorithm. Communications of the ACM, 14 (2):69–73, February 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shwayder:1974:EIT

[Shw74] Keith Shwayder. Extending the information theory approach to converting limited-entry decision tables to computer programs. Communications of the ACM, 17(9):532– 537, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shwayder:1975:CDR

[Shw75] Keith Shwayder. Combining decision rules in a decision table. Communications of the ACM, 18(8): 476–480, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Siler:1976:SEM

[Sil76] Kenneth F. Siler. A stochastic evaluation model for database organization in data retrieval systems. Communications of the ACM, 19 (2):84–95, February 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simmons:1970:NLQ

[Sim70] R. F. Simmons. Natural language question-answering systems: 1969. Communications of the ACM, 13 (1):15–30, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simmons:1974:AMP

[Sim74] Grover C. Simmons. Addendum to M. L. Patrick paper. Communications of the ACM, 17(4):209–??, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singleton:1969:AAEa

[Sin69a] R. C. Singleton. ACM Algorithm 347: An efficient algorithm for sorting with minimal storage. Communications of the ACM, 12(3): 185–187, March 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [GR70, Pet70].

Singleton:1969:AAEb

[Sin69b] R. C. Singleton. ACM Algorithm 357: An efficient prime number generator. Communications of the ACM, 12(10):563–564, October 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [De 73].

Sipala:1977:RAS

[Sip77] Paolo Sipala. Remark on “Algorithm 408: A Sparse Matrix Package (Part I) [F4]”. ACM Transactions on Mathematical Software, 3(3):303, September 1977. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [McN71].

Sitton:1970:OGA

[Sit70] G. A. Sitton. Operations on generalized arrays with the Genie compiler. Communications of the ACM, 13(5):284–286, May 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sites:1978:OSS

[Sit78] Richard L. Sites. Optimal shift strategy for a block-transfer CCD memory. Communications of the ACM, 21(5):423–425, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sreenivasan:1974:CRS

[SK74] K. Sreenivasan and A. J. Kleinman.

On the construction of a representative synthetic workload. Communications of the ACM, 17(3):127–133, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shapiro:1975:ICN

[SK75] Stuart C. Shapiro and Stanley C.

Kwasny. Interactive consulting via natural language. Communications of the ACM, 18(8):459–462, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Skinner:1974:HAI

[Ski74] C. William Skinner. Heuristic approach to inductive inference in fact retrieval systems. Communications of the ACM, 17(12):707–712, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stockman:1976:SPR

[SKK76] George Stockman, Laveen N.

Kanal, and M. C. Kyle. Structural pattern recognition of carotid pulse waves using a general waveform parsing system. Communications of the ACM, 19(12):688–695, December 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Skovgaard:1975:RBF

[Sko75a] Ove Skovgaard. Remark on “Algorithm 236: Bessel Functions of the First Kind [S17]”. ACM Transactions on Mathematical Software, 1 (3):282–284, September 1975. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Gau64b].

Skovgaard:1975:RAJ

[Sko75b] Ove Skovgaard. Remark on “Algorithm 332: Jacobi polynomials”. Communications of the ACM, 18(2):116–117, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wit68].

Skovgaard:1978:RCE

[Sko78] Ove Skovgaard. Remark on “Algorithm 149: Complete Elliptic Integral [S21]”. ACM Transactions on Mathematical Software, 4(1):95, March 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Mer62].

Slagle:1971:AGT

[SL71] James R. Slagle and Richard C. T. Lee. Application of game tree searching techniques to sequential pattern recognition. Communications of the ACM, 14(2):103–110, February 1971. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Slagle:1979:ESA

[Sla79] James R. Slagle. Experiments with some algorithms that find central solutions for pattern classification. Communications of the ACM, 22 (3):178–183, March 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spitzen:1978:EHD

[SLR78] Jay M. Spitzen, Karl N. Levitt, and Lawrence Robinson. An example of hierarchical design and proof. Communications of the ACM, 21(12): 1064–1075, December 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sterling:1975:CHL

[SLSH75] T. Sterling, V. Lum, N. Shu, and B. Housel. CONVERT : A high level translation definition language for data conversion. Communications of the ACM, 18(10):557–567, October 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Slysz:1974:ESS

[Sly74] William D. Slysz. An evaluation of statistical software in the social sciences. Communications of the ACM, 17(6):326–332, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sattley:1970:CPL

[SM70a] K. Sattley and R. Millstein. Comments on a paper by Lowe. Communications of the ACM, 13(7):450– 451, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1970:AAS

[SM70b] Robert R. Smith and Dennis Mc-Call. ACM Algorithm 392: Systems of hyperbolic P.D.E. Communications of the ACM, 13(9):567– 570, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Fri72].

Stein:1971:CD

[SM71] Marvin L. Stein and William D. Munro. On complement division. Communications of the ACM, 14 (4):260–264, April 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1970:CMU

[Smi70a] J. L. Smith. Comment on multiprogramming under a page on demand strategy. Communications of the ACM, 13(4):265, April 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1970:UIG

[Smi70b] L. B. Smith. The use of interactive graphics to solve numerical problems. Communications of the ACM, 13(10):625–634, October 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1973:LPW

[Smi73] Michael H. Smith. A learning program which plays partnership dominoes. Communications of the ACM, 16(8):462–467, August 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1975:CPT

[Smi75a] A. J. Smith. Comments on a paper by T. C. Chen and I. T. Ho. Communications of the ACM, 18(8):463, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1975:CPC

[Smi75b] Alan Jay Smith. Comments on a paper by T. C. Chen and I. T. Ho. Communications of the ACM, 18(8):463, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [CH75].

Smith:1975:MOL

[Smi75c] Graham Smith. On maintenance of the opportunity list for class-teacher timetable problems. Communications of the ACM, 18(4):203–208, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1977:MMO

[Smi77] Alan Jay Smith. Multiprocessor memory organization and memory interference. Communications of the ACM, 20(10):754–761, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shneiderman:1977:EIU

[SMMH77] Ben Shneiderman, Richard E.

Mayer, Don McKay, and Peter Heller. Experimental investigations of the utility of detailed flowcharts in programming. Communications of the ACM, 20(6):373–381, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smoliar:1972:CMM

[Smo72] Stephen W. Smoliar. Comments on Moorer’s Music and computer composition. Communications of the ACM, 15(11):1000–1001, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smoliar:1974:ECC

[Smo74] Stephen W. Smoliar. Emotional content considered dangerous. Communications of the ACM, 17(3):164–165, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smoot:1976:DIS

[Smo76] Oliver R. Smoot. Development of an international system for legal protection of computer programs. Communications of the ACM, 19 (4):171–174, April 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Slagle:1973:EAT

[SN73] James R. Slagle and Lewis M. Norton. Experiments with an automatic theorem-prover having partial ordering inference rules. Communications of the ACM, 16(11): 682–688, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Snyder:1978:BRE

[Sny78a] Lawrence Snyder. B-Trees reexamined. Communications of the ACM, 21(7):594, July 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Snyder:1978:BTR

[Sny78b] Lawrence Snyder. On B-trees reexamined. Communications of the ACM, 21(7):594–??, 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Salazar:1974:GFS

[SO74] A. Salazar and R. V. Oakford.

A graph formulation of a school scheduling algorithm. Communications of the ACM, 17(12): 696–698, December 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [KM75b, KM75a, GJ75].

Soule:1975:AAB

[Sou75] Stephen Soule. Addition in an arbitrary base without radix conversion. Communications of the ACM, 18(6):344–346, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shen:1974:WBM

[SP74a] Kenneth K. Shen and James L.

Peterson. A weighted buddy method for dynamic storage allocation. Communications of the ACM, 17(10):558–562, October 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [SP75].

Sterling:1974:ITM

[SP74b] Theodor D. Sterling and Seymour V. Pollack. Ideal teaching machines — a solution to the pedagogic language problem. Communications of the ACM, 17(4):207–208, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shen:1975:CWB

[SP75] K. K. Shen and J. L. Peterson. Corrigendum: “A Weighted Buddy Method for Dynamic Storage Allocation”. Communications of the ACM, 18(4):202, April 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [SP74a].

Spath:1970:AAF

[Spä70a] H. Späth. ACM Algorithm 375: Fitting data to one exponential. Communications of the ACM, 13 (2):120–121, February 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spath:1970:AAL

[Spä70b] H. Späth. ACM Algorithm 376: Least squares fit by f(x) = A cos(Bx + C). Communications of the ACM, 13(2):121–122, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spacek:1972:PEP

[Spa72] Thomas R. Spacek. A proposal to establish a pseudo virtual memory via writable overlays. Communications of the ACM, 15(6):421– 426, June 1972. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Sprague:1972:WVC

[Spr72] Richard E. Sprague. A Western view of computer history. Communications of the ACM, 15(7):686– 692, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Sprugnoli:1977:PHF

[Spr77] Renzo Sprugnoli. Perfect hashing functions: A single probe retrieving method for static sets. Communications of the ACM, 20(11): 841–850, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This is the first discussion on perfect hashing; describes heuristics for constructing perfect hash functions. See comments in [AA79a].

Squire:1972:AAL

[Squ72] W. Squire. ACM Algorithm 429:

Localization of the roots of a polynomial [C2]. Communications of the ACM, 15(8):776–777, August 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wil73a, DH73].

Strong:1973:RCT

[SR73] James P. Strong, III and Azriel Rosenfeld. A region coloring technique for scene analysis. Communications of the ACM, 16(4):237–246, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schachter:1978:SNM

[SR78] Bruce J. Schachter and Azriel Rosenfeld. Some new methods of detecting step edges in digital pictures. Communications of the ACM, 21(2):172–176, February 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Salazar:1968:AAM

[SS68] R. C. Salazar and S. K. Sen. ACM Algorithm 333: Minit algorithm for linear programming. Communications of the ACM, 11(6):437–440, June 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [HOK73, Obr73].

Schroeder:1972:HAI

[SS72a] Michael D. Schroeder and Jerome H. Saltzer. A hardware architecture for implementing protection rings. Communications of the ACM, 15 (3):157–170, March 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simmons:1972:GED

[SS72b] Robert F. Simmons and Jonathan Slocum. Generating English discourse from semantic networks. Communications of the ACM, 15 (10):891–905, October 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Saylor:1973:AAC

[SS73] Paul E. Saylor and James D. Sebastian. ACM Algorithm 460: Calculation of optimum parameters for alternating direction implicit procedures [D3]. Communications of the ACM, 16(10):633–635, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shneiderman:1974:SDS

[SS74a] Ben Shneiderman and Peter Scheuermann. Structured data structures. Communications of the ACM, 17(10):566–574, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1974:LCE

[SS74b] Graham Smith and Ian M. Sefton.

On Lions’ counter example for Gotlieb’s method for the construction of school timetables. Communications of the ACM, 17(4): 196–197, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schumacher:1976:SAD

[SS76] Helmut Schumacher and Kenneth C. Sevcik. The synthetic approach to decision table conversion. Communications of the ACM, 19 (6):343–351, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1977:DAA

[SS77] John Miles Smith and Diane C. P.

Smith. Database abstractions: Aggregation. Communications of the ACM, 20(6):405–413, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://dblp.uni-trier.
                                        de/db/journals/cacm/SmithS77.
                                        html.

Sibley:1973:DDM

[ST73] Edgar H. Sibley and Robert W. Taylor. A data definition and mapping language. Communications of the ACM, 16(12):750– 759, December 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://dblp.uni-trier.de/db/
                                        journals/cacm/SibleyT73.html.

Slutz:1974:NCA

[ST74] Donald R. Slutz and Irving L. Traiger. A note on the calculation of average working set size. Communications of the ACM, 17(10): 563–565, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stehfest:1970:AAN

[Ste70a] Harald Stehfest. ACM Algorithm 368: Numerical inversion of Laplace transforms. Communications of the ACM, 13(1):47–49, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stewart:1970:AAE

[Ste70b] G. W. Stewart. ACM Algorithm 384: Eigenvalues and eigenvectors of a real symmetric matrix [F2]. Communications of the ACM, 13 (6):369–371, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [Ste70d].

Stewart:1970:IOS

[Ste70c] G. W. Stewart. Incorporating origin shifts into the QR algorithm for symmetric tridiagonal matrices. Communications of the ACM, 13 (6):365–367, June 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stewart:1970:RAF

[Ste70d] G. W. Stewart. Remark on “Algorithm 384: Eigenvalues and eigenvectors of a real symmetric matrix [F2]”. Communications of the ACM, 13(12):750, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ste70b].

Stenger:1973:AAS

[Ste73] Frank Stenger. An algorithm for the approximate solution of Wiener– Hopf integral equations. Communications of the ACM, 16(11):708– 710, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sterling:1974:GHC

[Ste74] Theodor D. Sterling. Guidelines for humanizing computerized information systems, A report from Stanley House. Communications of the ACM, 17(11):609–613, November 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Steele:1975:MCG

[Ste75] Guy L. Steele, Jr. Multiprocessing compactifying garbage collection. Communications of the ACM, 18 (9):495–508, September 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Ste76].

Steele:1976:CMC

[Ste76] Guy L. Steele, Jr. Corrigendum: “Multiprocessing Compactifying Garbage Collection”. Communications of the ACM, 19(6):354, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ste75].

Stewart:1978:CNT

[Ste78] William J. Stewart. A comparison of numerical techniques in Markov modeling. Communications of the ACM, 21(2):144–152, February 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sterling:1979:CDC

[Ste79] Theodor D. Sterling. Consumer difficulties with computerized transactions: an empirical investigation. Communications of the ACM, 22 (5):283–289, May 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stone:1970:AAT

[Sto70] R. A. Stone. ACM Algorithm 398: Tableless date conversion. Communications of the ACM, 13(10):621, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Rob72].

Stockhausen:1973:AOC

[Sto73] Peter F. Stockhausen. Adapting optimal code generation for arithmetic expressions to the instruction sets available on present-day computers. Communications of the ACM, 16 (6):353–354, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stone:1974:NCP

[Sto74] Harold S. Stone. A note on a combinatorial problem of Burnett and Coffman. Communications of the ACM, 17(3):165–166, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stonebraker:1988:RDS

[Sto88] Michael Stonebraker, editor. Readings in Database Systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1988. ISBN 0-934613-65-6. xii + 644 pp. LCCN QA76.9.D3 R4 1988. US$29.95.

Strunz:1973:DDT

[Str73] Horst Strunz. The development of decision tables via parsing of complex decision situations. Communications of the ACM, 16(6):366– 369, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Strauss:1974:AMH

[Str74a] Jon C. Strauss. An analytic model of the Hasp execution task monitor. Communications of the ACM, 17 (12):679–685, December 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stroud:1974:GHI

[Str74b] A. H. Stroud. Gauss harmonic interpolation formulas. Communications of the ACM, 17(8):471–475, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sturman:1972:CUS

[Stu72] Gerald M. Sturman. Computers and urban society. Communications of the ACM, 15(7):652–657, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sullins:1972:CAP

[Sul72] Walter L. Sullins. Certification of “Algorithm 266: Pseudo-random numbers [G5]”. Communications of the ACM, 15(12):1072–1073, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [PH65].

Sunday:1990:VFS

[Sun90] Daniel M. Sunday. A very fast substring search algorithm. Communications of the ACM, 33(8):132–142, August 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [BM77] and [KMP77].

Shanno:1972:MSV

[SW72] David F. Shanno and Roman L. Weil. Management science: a view from nonlinear programming. Communications of the ACM, 15 (7):542–549, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Schneider:1977:CSI

[SW77a] Ben Ross Schneider, Jr. and Reid M. Watts. Corrigendum: SITAR: An interactive text processing system for small computers. Communications of the ACM, 20(10):761–??, 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schneider:1977:SIT

[SW77b] Ben Ross Schneider, Jr. and Reid M. Watts. SITAR: an interactive text processing system for small computers. Communications of the ACM, 20(7):495–499, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shaw:1977:AVA

[SWL77] Mary Shaw, William A. Wulf, and Ralph L. London. Abstraction and verification in Alphard: Defining and specifying iteration and generators. Communications of the ACM, 20(8):553–564, August 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Salton:1975:VSM

[SWY75] Gerard Salton, A. Wong, and C. S.

Yang. A vector space model for automatic indexing. Communications of the ACM, 18(11):613– 620, November 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Syslo:1973:AAE

[Sys73] Maciej M. Syslo. ACM Algorithm 459: The elementary circuits of a graph [H]. Communications of the ACM, 16(10):632–633, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sys75].

Syslo:1975:RAE

[Sys75] Maciej M. Syslo. Remark on “Algorithm 459: The elementary circuits of a graph”. Communications of the ACM, 18(2):119–??, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Sys73].

Szymanski:1978:ACM

[Szy78] Thomas G. Szymanski. Assembling code for machines with spandependent instructions. Communications of the ACM, 21(4):300–308, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tadikamalla:1978:CGGa

[Tad78a] Pandu R. Tadikamalla. Computer generation of gamma random variables. Communications of the ACM, 21(5):419–422, May 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tadikamalla:1978:CGGb

[Tad78b] Pandu R. Tadikamalla. Computer generation of gamma random variables — II. Communications of the ACM, 21(11):925–928, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tajibnapis:1977:CPT

[Taj77] William D. Tajibnapis. A correctness proof of a topology information maintenance protocol for a distributed computer network. Communications of the ACM, 20 (7):477–485, July 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tan:1972:FIS

[Tan72] K. C. Tan. On Foster’s information storage and retrieval using AVL trees. Communications of the ACM, 15(9):843, September 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See C. C. Foster, Proc. ACM 20th National Conference, pp. 192–205 (1965).

Tanenbaum:1978:ISP

[Tan78] Andrew S. Tanenbaum. Implications of structured programming for machine architecture. Communications of the ACM, 21(3):237–246, March 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Taulbee:1976:PEP

[TC76] Orrin E. Taulbee and Samuel D.

Conte. Production and employment of Ph.D.’s in computer science. Communications of the ACM, 19(6):311–313, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Taulbee:1977:CPE

[TC77a] Orrin E. Taulbee and Samuel D.

Conte. Corrigendum: Production and employment of Ph.D.’s in computer science — 1976. Communications of the ACM, 20(8):602–??, 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Taulbee:1977:PEP

[TC77b] Orrin E. Taulbee and Samuel D. Conte. Production and employment of Ph.D.’s in computer science — 1976. Communications of the ACM, 20(6):370–372, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Taulbee:1979:PEP

[TC79] Orrin E. Taulbee and Samuel D. Conte. Production and employment of Ph.D.s in computer science — 1977 and 1978. Communications of the ACM, 22(2):75–76, February 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Towsley:1978:MPP

[TCB78] Donald F. Towsley, K. Mani Chandy, and James C. Browne. Models for parallel processing within programs: Application to CPU:I/O and I/O:I/O overlap. Communications of the ACM, 21 (10):821–831, October 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Talbot:1971:ALT

[TCCH71] Peggy Anne Talbot, John W. Carr III, Richard R. Coulter, Jr., and Rosa C. Hwang. Animator: An on-line two-dimensional film animation system. Communications of the ACM, 14(4):251–259, April 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teichroew:1971:ERU

[Tei71] Daniel Teichroew. Education related to the use of computers in organizations. Communications of the ACM, 14(9):573–588, September 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tennent:1976:DSP

[Ten76] Robert D. Tennent. The denotational semantics of programming languages. Communications of the ACM, 19(8):437–453, August 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tenney:1977:RAT

[Ten77] Dennis Tenney. Remark on “Algorithm 219: Topological Ordering for PERT Networks”. ACM Transactions on Mathematical Software, 3 (3):303, September 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Kas63].

Tenenbaum:1978:SDS

[Ten78] Aaron M. Tenenbaum. Simulations of dynamic sequential search algorithms. Communications of the ACM, 21(9):790–791, September 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teorey:1978:GEI

[Teo78] Toby J. Teorey. General equations for idealized CPU-I/O overlap configurations. Communications of the ACM, 21(6):500–507, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tesch:1974:RAR

[Tes74] Gerhard Tesch. Remark on “Algorithm 456: Routing problem”. Communications of the ACM, 17(12):706–??, 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Fen73].

Tucker:1971:DMP

[TF71] Allen B. Tucker and Michael J. Flynn. Dynamic microprogramming. processor organization and programming. Communications of the ACM, 14(4):240–250, April 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thompson:1970:IRR

[Tho70] R. J. Thompson. Improving roundoff in Runge–Kutta computations with Gill’s method. Communications of the ACM, 13(12):739– 740, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thurnau:1963:AAB

[Thu63] D. H. Thurnau. ACM Algorithm 195: BANDSOLVE. Communications of the ACM, 6(8):441, August 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sch72d].

Tiernan:1970:ESA

[Tie70] James C. Tiernan. An efficient search algorithm to find the elementary circuits of a graph. Communications of the ACM, 13(12): 722–726, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thompson:1977:SMC

[TK77] C. D. Thompson and H. T. Kung.

Sorting on a mesh-connected parallel computer. Communications of the ACM, 20(4):263–271, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Todd:1975:CLC

[Tod75a] John Todd. Corrigendum: “The Lemniscate Constants”. Communications of the ACM, 18(8):462, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Tod75b].

Todd:1975:LC

[Tod75b] John Todd. The lemniscate constants. Communications of the ACM, 18(1):14–19, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder. See corrigendum [Tod75a].

Teorey:1972:CAD

[TP72] Toby J. Teorey and Tad B. Pinkerton. A comparative analysis of disk scheduling policies. Communications of the ACM, 15(3): 177–184, March 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://dblp.uni-trier.de/db/
journals/cacm/TeoreyP72.html.

Tanimoto:1977:EPS

[TP77] Steven L. Tanimoto and Theodosios Pavlidis. Editing of picture segmentations using local analysis of graphs. Communications of the ACM, 20(4):223–229, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Traub:1972:NMC

[Tra72] J. F. Traub. Numerical mathematics and computer science. Communications of the ACM, 15(7):537– 541, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Turner:1977:ULS

[TS77] Rollins Turner and William D. Strecker. Use of the LRU stack depth distribution for simulation of paging behavior. Communications of the ACM, 20(11):795– 798, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tsao:1974:DSD

[Tsa74] Nai Kuan Tsao. On the distributions of significant digits and roundoff errors. Communications of the ACM, 17(5):269–271, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tharp:1982:PTS

[TT82] A. L. Tharp and K.-C. Tai. The practicality of text signatures for accelerating string searching. Software — Practice and Experience, 12 (1):35–44, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Expands Harrison’s work [Har71b].

Tarjan:1979:SST

[TY79] Robert Endre Tarjan and Andrew Chi Chih Yao. Storing a sparse table. Communications of the ACM, 22(11):606–611, November 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ulrich:1978:EMD

[Ulr78] Ernst G. Ulrich. Event manipulation for discrete simulations requiring large numbers of events. Communications of the ACM, 21 (9):777–785, September 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanEmden:1970:AAI

[van70a] M. H. van Emden. ACM Algorithm 402: Increasing the efficiency of Quicksort. Communications of the ACM, 13(11):693–694, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Whe73].

vanEmden:1970:IEQ

[van70b] M. H. van Emden. Increasing the efficiency of quicksort. Communications of the ACM, 13(9):563– 567, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanderPoel:1972:CCS

[van72] W. L. van der Poel. Comment on the composition of semantics in Algol 68. Communications of the ACM, 15(8):772, August 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

VanRijsbergen:1974:BMP

[Van74] C. J. Van Rijsbergen. The bestmatch problem in document retrieval. Communications of the ACM, 17(11):648–649, November 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanderSluis:1975:CPE

[van75a] A. van der Sluis. Corrigendum: “Perturbations of eigenvalues of non-normal matrices”. Communications of the ACM, 18(3):180, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [van75b].

vanderSluis:1975:PEN

[van75b] A. van der Sluis. Perturbations of eigenvalues of non-normal matrices. Communications of the ACM, 18 (1):30–36, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder. See corrigendum [van75a].

vanGelder:1977:SPC

[van77] Allen van Gelder. Structured programming in COBOL: an approach for application programmers. Communications of the ACM, 20(1): 2–12, January 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vonBochmann:1973:MEL

[vB73] Gregor von Bochmann. Multiple exits from a loop without the GOTO. Communications of the ACM, 16 (7):443–444, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vonBochmann:1976:SEL

[vB76] Gregor von Bochmann. Semantic evaluation from left to right. Communications of the ACM, 19 (2):55–62, February 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vaucher:1975:CCS

[VD75a] J. G. Vaucher and P. Duval. Corrigendum: “A comparison of simulation event list algorithms”. Communications of the ACM, 18(8):462, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [VD75b].

Vaucher:1975:CSE

[VD75b] Jean G. Vaucher and Pierre Duval.

A comparison of simulation event list algorithms. Communications of the ACM, 18(4):223–230, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [VD75a].

VanderBrug:1975:SSP

[vdBM75] Gordon J. van der Brug and Jack Minker. State-space, problemreduction, and theorem proving — some relationships. Communications of the ACM, 18(2):107–115, February 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nat:1979:IS

[vdN79] M. van der Nat. On interpolation search. Communications of the ACM, 22(12):681–??, 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Veillon:1974:AAN

[Vei74] Francoise Veillon. ACM Algorithm 486: Numerical inversion of Laplace transform. Communications of the ACM, 17(10):587–589, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [KM76b, Vei77].

Veillon:1977:RNI

[Vei77] Françoise Veillon. Remark on “Algorithm 486: Numerical Inversion of Laplace Transform”. ACM Transactions on Mathematical Software, 3(1):111, March 1977. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Vei74].

Vere:1970:TEE

[Ver70a] S. Vere. Translation equations (errata). Communications of the ACM, 13(5):286, May 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ver70b].

Vere:1970:TE

[Ver70b] Steven Vere. Translation equations. Communications of the ACM, 13 (2):83–89, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See errata [Ver70a].

Verhelst:1972:CLE

[Ver72] M. Verhelst. The conversion of limited-entry decision tables to optimal and near-optimal flowcharts: Two new algorithms. Communications of the ACM, 15(11):974– 980, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vos:1973:RAC

[Vos73] H. Vos. Remark on “Algorithm 300:

Coulomb wave functions”. Communications of the ACM, 16(5):308– 309, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gun67].

vonGlaserfeld:1970:MPH

[vP70] E. von Glaserfeld and P. P. Pisani.

The multistore parser for hierarchical syntactic structures. Communications of the ACM, 13(2):74–82, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Swieten:1979:RVS

[vSdH79] A. C. M. van Swieten and J. Th. M.

de Hosson. Remark on “Algorithm 475: Visible Surface Plotting Program”. ACM Transactions on Mathematical Software, 5 (4):521–523, December 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Wri74].

Vuillemin:1978:DSM

[Vui78] Jean Vuillemin. A data structure for manipulating priority queues. Communications of the ACM, 21 (4):309–315, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wadler:1976:AAR

[Wad76] Philip L. Wadler. Analysis of an algorithm for real time garbage collection. Communications of the ACM, 19(9):491–500, September 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See erratum [Gri77b].

Wagner:1970:FAI

[Wag70] Robert A. Wagner. Finiteness assumptions and intellectual isolation of computer scientists. Communications of the ACM, 13(12):759– 760, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagner:1973:AAA

[Wag73a] Robert A. Wagner. ACM Algorithm 444: An algorithm for extracting phrases in a space-optimal fashion [Z]. Communications of the ACM, 16(3):183–185, March 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagner:1973:CPM

[Wag73b] Robert A. Wagner. Common phrases and minimum-space text storage. Communications of the ACM, 16(3):148–152, March 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagner:1974:OCR

[Wag74] Robert A. Wagner. Order-n correction for regular languages. Communications of the ACM, 17(5):265– 268, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wu:1977:ICG

[WAG77] Sheng-Chuan C. Wu, John F. Abel, and Donald P. Greenberg. An interactive computer graphics approach to surface representation. Communications of the ACM, 20(10): 703–712, October 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Waite:1970:MPS

[Wai70] W. M. Waite. The mobile programming system: STAGE 2. Communications of the ACM, 13(7):415– 421, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wallick:1970:AAM

[Wal70] G. C. Wallick. ACM Algorithm 400:

Modified Håvie integration. Communications of the ACM, 13(10): 622–624, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Pie74b].

Walker:1971:BS

[Wal71] R. J. Walker. Binary summation. Communications of the ACM, 14(6):417, June 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Walden:1972:NCN

[Wal72a] David C. Walden. A note on Cheney’s nonrecursive list-compacting algorithm. Communications of the ACM, 15(4):275, April 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Walden:1972:SIC

[Wal72b] David C. Walden. A system for interprocess communication in a resource sharing computer network. Communications of the ACM, 15 (4):221–230, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reproduced in “Advances in Computer Communications”, W. W. Chu, (Ed.), (1974), 340–349.

Wallace:1974:CGG

[Wal74] N. D. Wallace. Computer generation of gamma random variates with non-integral shape parameters. Communications of the ACM, 17 (12):691–695, December 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Walter:1976:RAC

[Wal76] Kenneth G. Walter. Recursion analysis for compiler optimization. Communications of the ACM, 19 (9):514–516, September 1976. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Waltz:1978:ELQ

[Wal78] David L. Waltz. An English language query answering system for a large relational data base. Communications of the ACM, 21(7):526– 539, July 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Warren:1975:MWA

[War75] Henry S. Warren, Jr. A modification of Warshall’s algorithm for the transitive closure of binary relations. Communications of the ACM, 18(4):218–220, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://dblp.uni-trier.
de/db/journals/cacm/Warren75.
html.

Warren:1977:FRW

[War77] Henry S. Warren, Jr. Functions realizable with word-parallel logical and two’s-complement addition instructions. Communications of the ACM, 20(6):439–441, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ward:1978:RTP

[War78] Stephen A. Ward. Real time plotting of approximate contour maps. Communications of the ACM, 21 (9):788–790, September 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wasserman:1975:PLP

[Was75] Anthony I. Wasserman. A problemlist of public policy issues concerning computers and health care. Communications of the ACM, 18 (5):279–280, May 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Watson:1970:ANM

[Wat70] G. A. Watson. On an algorithm for nonlinear minimax approximation. Communications of the ACM, 13(3):160–162, March 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Watkins:1973:RAG

[Wat73] Richard P. Watkins. Remark on “Algorithm 412: Graph plotter”. Communications of the ACM, 16 (8):489–490, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Čer71].

Watkins:1974:AAM

[Wat74] Steven L. Watkins. ACM Algorithm 483: Masked three-dimensional plot program with rotations [J6]. Communications of the ACM, 17(9): 520–523, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Fei75].

Wong:1971:CSA

[WC71] Eugene Wong and T. C. Chiang. Canonical structure in attribute based file organization. Communications of the ACM, 14(9):593– 597, September 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also published in/as: PhD Th., UCB, EECS, 1973.

Wulf:1974:HKM

[WCC^{+}74] William A. Wulf, Ellis S. Cohen, William M. Corwin, Anita K. Jones, Roy Levin, C. Pierson, and Fred J. Pollack. HYDRA: The kernel of a multiprocessor operating system. Communications of the ACM, 17 (6):337–345, June 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wetherbe:1978:SPM

[WD78] James C. Wetherbe and V. Thomas Dock. Strategic planning methodology for the computing effort in higher education: an empirical evaluation. Communications of the ACM, 21(12):1008–1015, December 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilcox:1976:DIT

[WDT76] Thomas R. Wilcox, Alan M. Davis, and Michael H. Tindall. Design and implementation of a table driven, interactive diagnostic programming system. Communications of the ACM, 19(11):609–616, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegner:1973:TSP

[Weg73] Eberhard Wegner. Tree-structured programs. Communications of the ACM, 16(11):704–705, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegbreit:1974:SLP

[Weg74a] Ben Wegbreit. The synthesis of loop predicates. Communications of the ACM, 17(2):102–112, February 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegbreit:1974:TDT

[Weg74b] Ben Wegbreit. The treatment of data types in EL1. Communications of the ACM, 17(5):251–264, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegbreit:1975:MPA

[Weg75] Ben Wegbreit. Mechanical program analysis. Communications of the ACM, 18(9):528–539, September 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegbreit:1976:FRC

[Weg76] Ben Wegbreit. Faster retrieval from context trees. Communications of the ACM, 19(9):526–529, September 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weiss:1972:PCI

[Wei72] Eric A. Weiss. Publications in computing: An informal review. Communications of the ACM, 15 (7):491–497, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Wendorf:1978:SEM

[Wen78] J. W. Wendorf. A simply extended and modified batch environment graphical system (sembegs). Communications of the ACM, 21 (11):897–904, November 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

West:1979:UMV

[Wes79] D. H. D. West. Updating mean and variance estimates: an improved method. Communications of the ACM, 22(9):532–535, September 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wexelblat:1973:ACC

[Wex73] Richard L. Wexelblat. Another comment on computer music. Communications of the ACM, 16(5): 313–314, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weber:1975:DEW

[WG75] Richard E. Weber and Bruce Gilchrist. Discrimination in the employment of women in the computer industry. Communications of the ACM, 18(7):416–418, July 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wong:1979:SAP

[WG79] J. W. Wong and G. Scott Graham.

Self-assessment procedure VI: a selfassessment procedure dealing with queueing network models of computer systems. Communications of the ACM, 22(8):449–454, August 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

White:1976:RMS

[WGL76] G. M. White, S. Goudreau, and J. L. Legros. Remark on “Algorithm 479: A Minimal Spanning Tree Clustering Method [Z]”. ACM Transactions on Mathematical Software, 2(1):110–111, March 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Pag74a].

Wheeler:1973:IEQ

[Whe73] Robert E. Wheeler. Remark on “Algorithm 402: Increasing the efficiency of Quicksort”. Communications of the ACM, 16(5):311, May 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [van70a].

White:1970:AAN

[Whi70a] J. S. White. ACM Algorithm 373: Number of doubly restricted partitions. Communications of the ACM, 13(2):120, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

White:1970:AAR

[Whi70b] J. S. White. ACM Algorithm 374: Restricted partition generator. Communications of the ACM, 13(2):120, February 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whitney:1972:AAM

[Whi72] V. Kevin M. Whitney. ACM Algorithm 422: Minimal spanning tree [H]. Communications of the ACM, 15(4):273–274, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Ker73].

Wilkes:1970:CAM

[Wil70a] M. A. Wilkes. Conversational access to a 2048-word machine. Communications of the ACM, 13(7):407– 414, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilson:1970:OSA

[Wil70b] M. Wayne Wilson. Optimal starting approximations for generating square root for slow or no divide. Communications of the ACM, 13 (9):559–560, September 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilson:1970:GMG

[Wil70c] W. J. Wilson. A generalized method for generating argument/ function values. Communications of the ACM, 13(11):696–697, November 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Williams:1971:SUM

[Wil71] John G. Williams. Storage utilization in a memory hierarchy when storage assignment is performed by a hashing algorithm. Communications of the ACM, 14(3):172–175, March 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Williamson:1972:AAH

[Wil72] Hugh Williamson. ACM Algorithm 420: Hidden-line plotting program [J6]. Communications of the ACM, 15(2):100–103, February 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [MC73, Wil73c, Ell74a, Gai74b, Ell74b].

Williams:1973:RAL

[Wil73a] Edward J. Williams. Remark on “Algorithm 429: Localization of the roots of a polynomial”. Communications of the ACM, 16(8):490, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Squ72, DH73].

Williams:1973:AMH

[Wil73b] John G. Williams. Asymmetric memory hierarchies. Communications of the ACM, 16(4):213–222, April 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Williamson:1973:RAH

[Wil73c] Hugh Williamson. Remark on “Algorithm 420: Hidden-line plotting program”. Communications of the ACM, 16(9):578–579, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wil72, MC73, Ell74a, Ell74b, Gai74b].

Willers:1974:NIA

[Wil74] I. M. Willers. A new integration algorithm for ordinary differential equations based on continued fraction approximations. Communications of the ACM, 17(9):504– 508, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilks:1975:IAU

[Wil75a] Yorick Wilks. An intelligent analyzer and understander of English. Communications of the ACM, 18 (5):264–274, May 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Willoughby:1975:SIH

[Wil75b] Ralph A. Willoughby. Special issue honoring Alston S. Householder — Editor’s foreword. Communications of the ACM, 18(1):3–??, 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilhelm:1976:ADS

[Wil76] Neil C. Wilhelm. An anomaly in disk scheduling: A comparison of FCFS and SSTF seek scheduling using an empirical model for disk accesses. Communications of the ACM, 19(1):13–17, January 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winters:1971:SAC

[Win71] William K. Winters. Scheduling algorithm for a computer assisted registration system. Communications of the ACM, 14(3):166–171, March 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winograd:1979:BPL

[Win79] Terry Winograd. Beyond programming languages. Communications of the ACM, 22(7):391–401, July 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1969:MMC

[Wir69] N. Wirth. On multiprogramming, machine coding, and computer organization. Communications of the ACM, 12(9):489–498, September 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Wir70].

Wirth:1970:CMM

[Wir70] N. Wirth. Corrigenda: “On Multiprogramming, Machine Coding, and Computer Organization”. Communications of the ACM, 13 (4):266, April 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wir69].

Wirth:1971:PDS

[Wir71] Niklaus Wirth. Program development by stepwise refinement. Communications of the ACM, 14(4): 221–227, April 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1977:TDR

[Wir77a] Niklaus Wirth. Toward a discipline of real-time programming. Communications of the ACM, 20(8): 577–583, August 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:1977:WCW

[Wir77b] Niklaus Wirth. What can we do about the unnecessary diversity of notation for syntactic definitions? Communications of the ACM, 20 (11):822–823, November 1977. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wise:1976:RLE

[Wis76] David S. Wise. Referencing lists by an edge. Communications of the ACM, 19(6):338–342, June 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Witte:1968:AAJ

[Wit68] B. F. W. Witte. ACM Algorithm 332: Jacobi polynomials. Communications of the ACM, 11(6):436– 437, June 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sko75b].

Withers:1974:RAZ

[Wit74] David H. Withers. Remark on “Algorithm 419: Zeros of a complex polynomial”. Communications of the ACM, 17(3):157–??, 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [JT72].

Wu:1992:FTS

[WM92] Sun Wu and Udi Manber. Fast text searching allowing errors. Communications of the ACM, 35 (10):83–91, October 1992. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0001-
0782/135244.html. This algorithm

in this paper is implemented in the agrep program, publicly available via ANONYMOUS FTP to cs.

arizona.edu in the agrep subdi-

rectory. See also [BYG92].

Woods:1970:CSP

[Woo70a] W. A. Woods. Context-sensitive parsing. Communications of the ACM, 13(7):437–445, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Woods:1970:TNG

[Woo70b] W. A. Woods. Transition network grammars for natural language analysis. Communications of the ACM, 13(10):591–606, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wulf:1971:BLS

[WRH71] W. A. (William A.) Wulf, D. B. Russell, and A. N. Habermann. BLISS: A language for systems programming. Communications of the ACM, 14(12):780–790, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://doi.acm.
org/10.1145/362919.362936.

Wright:1974:AAV

[Wri74] Thomas Wright. ACM Algorithm 475: Visible surface plotting program [J6]. Communications of the ACM, 17(3):152–155, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [Fre75, Mas75, Dor75, Dut76, vSdH79].

Wunderlich:1974:DNT

[WS74] Marvin C. Wunderlich and J. L. Selfridge. Design for a number theory package with an optimized trial division routine. Communications of the ACM, 17(5):272–276, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whitt:1975:ASA

[WS75] Jerry D. Whitt and A. Gale Sullenberger. The algorithm sequential access method: An alternative to index sequential. Communications of the ACM, 18(3):174–176, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wyman:1975:IES

[Wym75] F. Paul Wyman. Improved eventscanning mechanisms for discrete event simulation. Communications of the ACM, 18(6):350–353, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yao:1977:ANA

[Yao77] S. Bing Yao. Approximating the number of accesses in database organizations. Communications of the ACM, 20(4):260–261, April 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://dblp.uni-trier.de/
db/journals/cacm/Yao77.html.

Yap:1976:NUB

[Yap76] Chee K. Yap. New upper bounds for selection. Communications of the ACM, 19(9):501–508, September 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cheung:1974:IGD

[yC74] To yat. Cheung. Interactive graphic display for region partitioning by linear programming. Communications of the ACM, 17(9):513– 516, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yohe:1972:AAH

[Yoh72] J. M. Yohe. ACM Algorithm 428: Hu-Tucker minimum redundancy alphabetic coding method [Z]. Communications of the ACM, 15(5):360–362, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Byr73b].

Young:1974:FOA

[You74] John W. Young. A first order approximation to the optimal checkpoint interval. Communications of the ACM, 17(9):530–531, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yu:1977:EIR

[YS77] Clement T. Yu and Gerard Salton. Effective information retrieval using term accuracy. Communications of the ACM, 20(3):135–142, March 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zobrist:1977:DCO

[ZC77] Albert L. Zobrist and Frederick R. Carlson, Jr. Detection of combined occurrences. Communications of the ACM, 20(1):31–35, January 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zloof:1977:SBA

[Zd77] Moshé M. Zloof and S. Peter de Jong. The system for Business Automation (SBA): Programming language. Communications of the ACM, 20(6):385–396, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zelkowitz:1971:IDP

[Zel71] Marvin V. Zelkowitz. Interrupt driven programming. Communications of the ACM, 14(6):417–418, June 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zelkowitz:1973:RE

[Zel73] Marvin V. Zelkowitz. Reversible execution. Communications of the ACM, 16(9):566, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zinn:1972:CIP

[Zin72] Karl L. Zinn. Computers in the instructional process: directions for research and development. Communications of the ACM, 15(7):648– 651, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zweben:1978:OMD

[ZM78] Stuart H. Zweben and M. A.

McDonald. An optimal method for deletion in one-sided heightbalanced trees. Communications of the ACM, 21(6):441–445, June 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).